A Complete Bibliography of the Journal Family Printing History

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: http://www.math.utah.edu/~beebe/

04 April 2024 Version 1.08

Title word cross-reference

0f [Gob98b, Sto98].

1 [Car92, Rum98]. 1473-1500 [Ame79]. 1478 [Tra17]. 1480-1599 [Tra84]. 1516 [Dan02]. 1599 [Tra84]. 1690-1832 [Ber16]. 16th [Szm86, Ann85]. 18 [La 16a]. 1800 [Isa86, Sil82a, Met81]. 1812 [Bum82]. 1820 [Ste80a]. 1825 [Tur81]. 1830 [Gre81]. 1840 [Kod10c]. 1860 [WP09d]. 1865 [Sul83]. 1876 [Pan85]. 1880 [Kod10b]. 1880s [Dun88a, Sta85]. 1890 [Joh09c, Wal95]. 1890s [Gob98b]. 1896-1912 [Kel08c]. 1899 [Kra79b]. 1900 [Thé83]. 1912 [WP09c]. 1914 [Spa83]. 1920s [Smi05a]. 1928 [Bid92]. 1931 [Win92]. 1935 [La 16a]. 1940 [Kod09a]. 1940s [Ste85, Sul86]. 1970s [Smi05a]. 1972 [Ano72]. 1975 [All83, Ste79a]. 1977 [Mat82]. 1978 [Teb79]. 1980 [Tan99]. 1981 [New84]. 1982 [CH83, Gil82, Dav84]. 1992 [Dre93]. 1998 [BS99]. 19th [Pan88, Har88].

2 [GK10]. 2000 [Ano12a]. 2004 [WP09a]. 20th [Pan88]. 23 [Sha18b]. 2d [Lar92, Pet85].

31/32 [BBRD99].

98 [Han00].

A. [Sta81]. Abraham [Wei02]. Account [Mat82, Win92, DG93, MS80]. Active [Bur07, Kel09a]. Adam [Mif16b].

Adaptation [Arg92, Lan05, WP07g]. Adapted [Bue82]. additions [Hen95]. Adobe [Arg92]. Adventure [Sul83, WCB82]. Advertising [Can00]. Advocating [Dun79, Dun80]. Aesthetic [Eri17, Maz15, Mif17a]. After [Bel95, Sto98]. Age [Sha07, Tho80a, BBB^{+}02, Teb78]. Agent [Tra82, Eis94]. Aikens [Can00]. Alan [Ano11a, DeC87, Joh09a, Sul84a]. Albany [Cyr93]. Albion [WP07m, Sto05]. Album [Gar91]. Aldus [Bar85, Bar92, Cyr93, Fle86, Fle93, Fle09]. Alexander [Ber87, Joh07c, Kel91, Pom05, Sre86]. Alice [Sil82b, WP08]. Allegations [BBRD99]. Allied [Thé83, Hul82]. Almanacs [Sha07, Win11]. Alphabet [HJB03, Sme08]. Alpheus [La 16a]. Altar [Amo90]. am [Sch87]. Amelia [Ash10]. Amelung [de 82]. America [Gre86, Hen95, Joh09c, Kod09a, Kod10b, Kod10c, Kos10, Sil82a, Ste79a, Tur81, Ann75, Bar76, BF09, Blu77a, Blu77b, Bri05, Caz84, Day66, Far19, FB^{+}09, Fli07, GK06, GK10, Han93, Han05, Hud08, KR09, Kod09b, Lan83, Lan16, Lar84, Mor91, Rum92a, Ste80c, Teb87, TM96, Wei89, Wei05, Win81, Win11, Wol80, Bid82, Gre87, Kra79a, WP07b]. American [Bro88a, Bro79, Dal18, Har81, Joh10, Kas82, Ste80a, Sul83, Sul87, Swo18, Tho79, Tur81, WP09i, Ano97, Ano80, BH80, Ber16, Bid80, Bid12, Bou12, Bum82, Dic86, Hud78, Joh79, Las05, Mak19, Old06, Phe85, Rei75, Rum92b, SD91, Sax91, SD92, Shi10, Sil79b, Sto87, Tho80b, WCB82, Wil84, Wol80]. American-Made [Phe85, Sax91]. Amert [Geh15, Dan02]. Amici [Mar01]. Among [Ber90]. Amory [Coa03]. Analytical [Tan81]. Anathema [Dro83, Nes84]. Anatomy [Law90a, Kel91]. Anderson [Joh07c, Pom05]. Andrew [You10a]. Ann [Fra89, For90]. Anne [Szm86]. Annenberg [Hen95]. Anninger [Szm86]. Anniversary [Pan85, Dwi80]. Annotated [Kos10, La 15, Bid84, GK06, Lan04, Pom05, SB82]. Announcement [Kri10]. anthology [MS73, Ash80]. Anthony [Wee85]. Antiquarian [Ste85, Sul86]. Any [Pan99]. APHA [Tho79]. Apocalypse [Dan02]. appreciation [You10a]. April [Bec07, Gil82]. Arab [Dal18]. Arabic [Gre09]. Arabic-Script [Gre09]. Archer [Sha18a]. Archer-Parré [Sha18a]. Architect [DeC87, Cra85, Cra05]. Architecture [WP07r, Smi05a, Smi05b]. Archive [Ano19a]. Ari [WP07a]. Aries [Sen17, Keg16]. Art [Ano72, Ano11b, Ano12a, Bri05, Fli07, Gra90, Mif12b, Sax86, Sha18a, Smi88, Spa83, Tho80b, WP07b, WP07e, Win95, HS85, Hop11, Joh79, Kel10, Rai13, Ray82, Roy86, Tho88, Tic05, Mif15, Tho79]. Article [BBRD99]. Artist [Kel91, Law90c, Swa13, LRS07, WP09g]. artistic [BW84]. Arts [Ano27, Ano82, Cav84, DeC84, Dun88a, Gar91, Har85, Smi88, Sta85, Whi05, MP82, Roy86, Smi05a, Smi05b, Ann85, Ano88a, HF06, Szm86, WP07r]. Artzt [Bre90]. Ashall [Tan82]. Ashbee [DeC87, Cra85, Cra05]. Ashendene [Fra88, Fra86a]. Asia [Dav11]. Aspects [Pre81, GL80]. Assaults [Sil79b]. ATF [Cos94, Panxx]. Atlantic [Sha07]. Attributing [Sin80]. August [La 16a, WP09a]. Augusta [Rum98]. Auji [Dal18]. Ausstellung [Ame79, de 82]. author [Ano19a]. Autobiography [Sil79a]. Autographed [La 16a]. Axel [Van84]. Axel-Nilsson [Van84].

B [Bid82, Fra89, Hil80, Kas82, Sil82a, Sul86, WP07n]. Back [Tho83, Win95]. Background [Sin80]. Bailyn [Kas82]. Bakhareva [Gle86]. Ballads [Kod10a, Kel08a]. Ballou [Gen84]. Baltimore [WP09a, FIbibBLL09]. Balzac [La 14, You12]. Band [de 82]. Banham [WP09h]. Barker [Fle86, Lar92, Pet85, Teb79]. Barlow [Ste79a, Tra79, Hod78]. Barnard [Auj14]. Baron [Han93]. Bartram [Joh09a]. Basil [Rai81, Rai82]. Baskerville [Sha18a]. Bay [Amo90]. be [Lan16]. Beautiful [Hel11, Kri10]. Beck [Sti85]. Beckwith [WP08]. Bed [Swo18]. Before [Har88]. Beginning [Wei05]. Beginnings [Har81, Ano80]. Behold [Hel11]. being [HS78]. Beirut [Dal18]. Beletsky [Ano18c]. Bell [Joy85]. Ben [Bid79]. Benjamin [GS06, Win07]. Benson [Ant10]. Benton [Cos94]. Beorum [Mar07]. bequest [Ano88a, Joh91]. Berger [Ano83]. Berkeley [Hut93]. Bermuda [Len88]. Bernard [Ano89, Kas82]. Berry [Haw17]. Bertha [La 16b]. best [Gre94]. Betty [WP07b]. Between [Teb87, Teb78, Gre87]. Bewick [Ugl06, WP07f]. Beyond [Mif12a, Sch18, Pea08]. Bible [Mak19]. Bibles [Far19]. Bibliographer [Wel80, BY79]. Bibliographical [Fra89, Wee88, Ann85, BW84, TB88, Mor91]. Bibliography [Kri82, May87, Rat92, Sul83, Tan81, Tra84, Eas83, GK06, HTBR05, HH08, Pet84, Pom05, WCB82, Dre93, Far83b, Joh07a, Kos10, WP09k]. Bibliophiles [WP09a, Ano03, FIbibBLL09, WP07q, BR03]. biblioteki [BL^{+}83, Gle86]. Bibliothek [Hut89, SIvA87]. Bidwell [Ber16]. Bill [DeV81]. Bimonthly [DeC84, Ano82]. Biographical [Kod10a, Kel08a]. Biography [Gob90, Lar92, Sul84a, Dye82, Sch89]. Birth [Kri10, Dwi80, Pan85]. Black [New84, Joy83]. Black-Owned [New84, Joy83]. Blackletter [BS^{+}98, BS99]. Blades [Ber93]. Blake [Bla72, Bla00, Gro01]. Blind [Har86]. Blockbook [Dan02]. Blood [Pan88]. Blumenthal [All83, Cor86, Kra79a, Sti90]. Board [Eri17]. Bodoni [Mat82, Ano91, MS80]. Bolton [Tra17].

Book

[All83, Ano79, Ano11a, Ano11b, Ano12a, Ano12b, Ano18b, Ano18a, Ano18c, Ano18d, Anoxx, Ant08, Ant10, Ash80, Auj14, Ber93, Ber16, Ber87, Bia87, Bid79, Bid82, Bid84, Bid92, Blu80, Bod93, Bon83, Bri05, Bro88a, Bro88b, Bro79, Bue82, Car94, Cav84, Cor86, Dal18, Dav11, Dav84, Day66, DeC84, DeC87, Dre87, Dre93, Due84, Due87, Due91, Dun79, Dun80, Dun88a, EK83, Far83a, Far83b, Far19, Fea86, Fle86, Fra86b, Fra88, Fra89, Geh98, Geh09, Geh15, Gle86, Gob90, Gra88, Gra90, Gre81, Gre86, Gre87, Har80, Har81, Hen95, Her09, Hey85, Hil80, Hut89, Isa86, Joh91, Joh07a, Joh08, Joh07b, Joh07d, Joh07c, Joh09a, Joh09b, Joh09c, Joh09d, Joh10, Joh11a, Joh11b, Joy85, KR09, Kas82, Kel91, Kel98, Kel03, Kel08b, Kel09a, Kod84]. Book [Kod09a, Kod09b, Kod10a, Kod10b, Kod10c, Kod11, Kor83, Kos10, Kra79a, Kri10, La 14, Lan83, Lar92, Law90b, Leu94, Lie79, Lis80, Mat82, May87, McT84, McT85, Met81, Mif12a, Mif12b, Mif13, Mif15, Mif16a, Mif16b, Mif16c, Mif16d, Mif17a, Mif17b, MW05, Nes84, New84, Pal19a, Pal19b, Pan85, Pan07, PS14, PM84, Pet85, Pre81, Rae88, Raf05, Rai81, Rai82, Ruf17, Ruf19, Rum92b, Sax80a, Sax86, Sax07, SIvA87, Sch17, Sen17, Sha18a, Sil82a, Sil84, Smi88, Spa83, Sta81, Ste79a, Ste80a, Sti85, Sto87, SW10, Sul83, Sul84a, Sul86, Sul87, Swo18, Szm86, Szm89, Tan82, Tan99, Tay90, Teb79, Teb86, Thé83, Tho80b, Tra79, Tra82, Tra84, Tra17, Tri84, Tur81, Van79, Van84, WP07a, WP07b, WP07c, WP07d, WP07e, WP07g, WP07f, WP07h, WP07k, WP07j]. Book [WP07l, WP07i, WP07o, WP07m, WP07n, WP07p, WP07q, WP07r, WP08, WP09a, WP09b, WP09k, WP09d, WP09c, WP09e, WP09f, WP09g, WP09j, WP09i, WP09h, Wal95, Wat84, Wee85, Wee88, Wel80, Whi05, Win92, Win95, Win07, de 82, AB12, BF86, Blu77a, Blu77b, Cas07, Caz84, DSC^{+}05, Dic86, Dre94, Dro83, Eve08, Gla79a, Gla79b, GK10, Hod78, HH08, Hun80b, Joy83, Kel10, Maz15, Mey83, MP82, Ray82, Hin82, Roy86, Teb78, Teb87, Win81, WB10, You12, McT85, Mif16b, Tra81, WP07a, WP07k, Ano18a, Auj14, Cav84, Kod11, Amo90, Anoxx, Hut90, Szm89]. Bookbinding [ER82, McT81, Tan82, Gla79a, Gla79b, Kod84]. Bookcloth [TM96, Tan99]. Booklet [Tho11]. bookmaker [Ano97, Tho79]. bookman [Bru86, Gra88]. Bookplate [Kri10]. bookplates [Hop11, Ano11b, Mif12b]. Books [Ano11a, Bid80, Dre93, Hel00, Hel15, Kod84, Kod10a, Kor83, Mif12a, Pea08, Spu82, Sto98, Szm86, Tra84, Uza79, WP08, WP09d, WP09g, Win11, Ann85, BL^{+}83, Bec07, BB05, Chr82, ER82, Kel08a, LRS07, LQS10, Pet07b, Rat92, Wil08, Mif12a, Wei83, WP09e]. Bookseller [Sin92, Yea12]. Booksellers [Thé83, Hul82]. Bookselling [Sul86, Ste85]. Boston [Ant08, Kra79b, Kos06, O’B03, Yea12]. Bowden [Fra89]. Box [Hop09]. Boyd [Cos94]. BR [Gen84]. Bradbury [Gra90]. Bradley [Dan80]. Bride [RB08]. Bridson [Wee88]. Brief [McK79, Sax81]. Bright [WP07b]. Brigid [McT85]. Britain [Joh09c, WP09d, BW84, BB05, Gre09, Hud08, TM96]. British [Rai82, Swo18, WP09k, HH08, Hun80a]. Broadside [Kri10]. Broadsides [Bum82]. Brotherhood [Mif13, Lea10]. Brothers [Kod10b, FB^{+}09, BF09]. Brown [WP09d]. Bruccoli [Gra88]. Bruce [Ano18c, WP09c, WP09g, Ano19a, Gen84, Kel08c, KBKHF16, LRS07, Sti85, Tho80a, Tho83]. Bruckner [Law90b]. Bruntjen [Wel80]. Brussel [Ano89]. Brussel-Smith [Ano89]. Buchanan [WP09d]. Buchanan-Brown [WP09d]. Buchgraphik [SIvA87]. Buchgrapik [Hut89]. Buchillustration [Hut89, SIvA87]. Buechler [Pan85]. builders [Har89]. Bullen [Leexxa]. Burgess [BBRD99, BBRD99, Par94].

Burke [Kel09a]. Business

[Fli07, Kri10, Met81, aD79]. Buxton

[BC83, Lar92, Pet85].

C [Bro88a, BY79]. C.

[Ben93, Cra85, Cra05, DeC87, Har03]. ca [Lan04]. California [Wal95, Gra07, Joh11c, LA94]. calligrapher [Ano72, HJ07, Kel08b]. Calligraphy [Hid07]. Came [Lan16]. Camp [Sul83]. Canadian [Pre81, GL80]. Cape [Ash10]. Capital [Sme08]. Carbutt [Han05]. Card [Kri10]. Career [BB88, McK79, Mif15, Mid85, Rai13]. Carl [Pan85, Gob90, Kod09a, Wal82]. Caroline [Sha18a]. Carroll [Bro79]. Carter [Lar92, Pet85, BC83, Lan02]. Case [Ber11, Blu80, Ryd79]. Caslon [Due84, Ber11, Bid80]. Casper [Kod09b]. Catalog [BS99, Har81]. catalogs [Ann75, Hen95]. Catalogue [CH83, Dav84, Joh91, WP08, Ann85, Ano80, Ano88a, AN83, Bec07, Bid84, FIbibBLL09, SB82, Pal19a, Van84]. Catalogues [Sil82a, Win81]. Cataneo [Wat84, HC81]. Cave [Tri84]. Caxton [Nee86]. Cazden [Gre86]. Celebration [Dun88b]. census [PP11, Ano12b]. Centaur [Ano18c, KBKHF16]. Centenary [Joh93, Ano72, Gil82]. Centennial [Dun88b]. Cento [Mar01]. Centuries [Ano18a, SIvA87, WP09e, Wil08, Hel00]. Century [All83, Ano12a, Ant08, Arg92, Bar93, Bid82, Dal18, Dav81, Dav84, Hel15, Joh09d, Joh10, Joh11c, Kor83, Lan83, Lar92, Mif17a, Pet85, Szm86, Tra17, Van79, WP07o, Win83, BL^{+}83, BC83, Bar85, Bar92, BB82, Can00, CP34, CP71, CP92, Clo08, CH83, Dys84, God16, Gra76, HS78, Kel10, Kos06, Maz15, OW80, PM84, Sou05, Ste80c, Wee85, Ann85, Bol16, Far19, Har88, Las05, Mak19, Mif17b, O’B03, Pan88, Pet07a, Rum92a, Spu82, Tho19, Win11, Fle86, Joh09b, McT85, Sta81].

Certain [Lar92, Pet85, BC83, CP34, CP71, CP92]. Chalmers [Tho80b]. Chambers [Wei83]. Champ [Sme08]. Chancery [Sta81, OW80]. Change [Fal02, Leu94, Tra82, Tra84, Chr82, Eis94, Kos06, Ant08]. Changes [Pom94]. Chants [Wil84, Sul87]. Chap [Dre93, Rat92]. Charing [Nee86]. Charles [Sul83, Swo18, WP07j, Ano18e, Bue83, Dun80, Pan85, Wat01, Wat04, Win11]. Charter [Rum14]. Chaucer [Ano12b, PP11]. Checklist [Sil82a, DM79, Gil82, Kel08c, Win81, You10a]. Checklists [Kod10a, Kel08a]. Chemists [Pet81]. Chicago [Han00, Eck87]. Children [McT84, Mey83]. Chinese [jK86]. Chinook [Vor19]. Chip [AAG87]. Chiswick [Bue82]. Chrisman [Tra84]. Christian [Van84]. Christopher [Gre86, Kel09a, La 15, WP07c]. Chronicles [Com79, Sax80a]. Circus [Fli07]. City [Sul87, Sax80b, Wil84]. Civil [Bou12, Gre86, Caz84]. Claiming [All87]. Claire [Tra17]. Clare [Ash10]. Clarifying [Kru86]. class [Wil84, Sul87]. Classification [Eli15]. Claude [Lan05]. Clinton [Lie79]. Clouse [Joh09b]. Club [Bec07]. Co [Wei83]. Co. [Sax81]. Cockerel [Cav82, DG93, Win92]. Cod [Ash10]. Codex [CCS18]. Cody [La 16a]. Colin [Fra88]. Colines [Geh15, Ame92, AB12]. collaboration [DG93]. Collected [DeV81]. Collecting [Her09, WP07l, HH06]. Collection [Ano88a, Van84, Ann85, AN83, BL^{+}83, Ano83, Har88, Her09, Shi10]. Collections [Mos08]. collectors [Dic86, Bro88a]. College [Ano72, Mar07]. Collins [Lar92, Pet85]. Collotype [RK91]. Colonial [Bro79, Ste79a, Sul84a, Yea12, Dye82, Rei75, Bar76]. Color [Fli07, HB13, Joh10, Las05, Wei05]. Colour [Tho19, Bla72, Gro01]. colours [Bla00]. Columbia [Har81, Ano80]. come [LQS10, Ano11a]. Comic [Win11]. Coming [Geh98]. commentary [Wol80]. Commercial [Han93, Har88]. Common [Bid92, Hut90, Lie79, Sax90, Sax91, HS78]. Communication [RB08]. Communications [Tra82, Eis94]. Communist [Zla13]. comp [Ste79a]. Comp. [Pre81]. Companion [Kod11, SW10]. Company [Sax80a, Sch87, Ano80, Bak92, Com79, Mar87, Rom14, Mif16a]. Comparative [Bod93, SZ91]. Comparato [Sax80a]. Comparison [Ame03]. complete [Bla00]. Composition [Due87, Rum92a, Hus85]. Compositor [Kel91, Law90c, Gob98a]. Computing [Lab17]. conference [Tra81]. Congress [Mil89a, Mil89b]. conjunction [Gil82]. Connecticut [Kod10b, BF09, FB^{+}09]. Connection [McG94]. Conquest [CCS18]. Consequences [Fal02]. Conservation [Kod84, ER82]. considered [DSC^{+}05, WP07c]. Considering [Shi10]. conspectus [Ver10, Joh11a]. Constantly [Joh92]. Construction [Car92]. Contemporary [WP07g, Lan05, PS14, Mif16b]. Context [Sme08]. Continental [Geh98, McT81]. contributions [Ano89]. Convivial [Sil82b]. Cookbooks [Tho19]. Copeland [Kra79b, Lis80]. Copperplate [Lar84]. copybook [HC81, Wat84]. Cornhill [Kra79b, Tho11]. correspondence [Mac79, Hil80]. costs [TM96]. Countries [Lan04]. Country [How80]. Coverly [Kel08a, Kod10a]. Covers [Gre87, Teb87]. Cowl [Gre94]. Craft [Pre84, Sax86, WP07e, HS85, Tic05]. Crafts [Ano82, DeC84, Sta85, Dun88a]. Craftsman [Kel91, Law90c]. Crawford [DeC87]. Creating [Rai13, Mif15]. Cree [McK79]. Critical [Sul83, WCB82]. Crosby [Lox08a]. Cross [Nee86]. Crump [Dav84]. Cubery [Har82]. Cultural [Tra82, Eis94]. Culture [Dal18, Fal02, Geh15, Kod10c, Mif13, New84, Pre84, Tra84, AB12, BBB^{+}02, Chr82, GK10, Joy83, Lea10]. Curious [Sin80]. Curse [Dan94]. curses [Dro83, Nes84]. Cutter [Pal19a].

D [Hil80, Joh09d, Joh11a, Kod09b, Pal19a, Rae88, Sch87, Sil82b]. D. [Law90b, Sch87]. Dale [Smi88]. Damned [Pet81]. Dana [Ste79a]. Daniel [All90, Hut93]. Danish [Ben93]. Dard [Wri92]. Darnton [EK83]. Dated [Auj14, La 16a, WB10]. David [Hil80, McK79, Mif12a]. Davis [Pan85]. Day [Kra79b, WP09b, Eve08, Lis80]. Deberny [HF06]. Defense [All90]. del [Mar01]. delineation [HS78]. democracy [Com79, Sax80a]. Democratic [Sul87, Wil84]. Denslow [Ano83]. Department [Szm86, Ann85, Ano88a, RB08]. Depictions [Per15]. description [Ann85, HS78]. descriptions [Lan04]. Descriptive [Kod84, Sil82a, ER82, Win81]. Design [Dun79, Dun80, Hid07, HF06, Joh08, Joh09c, Nea79, WP07o, WP07r, BBB^{+}02, Hud08, Jan04, Maz15, Smi05a, Smi05b, Sou05, Tho88, Ano19a, Gra90, Mif17a]. Designer [BBRD99, DeC87, Par94, WP07k, WP07n, WP09g, Cra85, Cra05, Dou06, LRS07, MW05]. Designers [Ano12a, Kel10, Mac06, WP09f]. Designs [Gro01, Tra94, Bla72, TM96]. Destruction [Mif16b, PS14]. deutschen [Ame79, de 82]. Developed [O’B03]. Development [Fle86, Gre09, Pan88, Bar85, Bar92, Hus85]. Devices [Hel00, Sch17]. Devil [Mif16d, Lox08b]. DeVinne [WP07e]. Devoted [DeC84, Ano82]. Diamond [Auj14, WB10]. Diane [Tho80b]. Dick [Sha18a]. Dickinson [Bro88a]. Dictionary [Bro79, Dic86, Fea86, Kod84, Rai82, Tay90, Thé83, ER82, Hul82, Hun80a, Rei75, Bro88a]. Digital [Ame03, Ano18d, Mif16c, Sha17, Sha18b, WP09j, Dod06, Hop12]. Dimensions [Mor81]. Dining [Sil82b].

Directions [Sul84b]. Directory

[Ber16, Bid12]. Disbound

[DSC^{+}05, WP07c]. Disciple [Cyr93]. Discovering [WP09e, Wil08]. Discovery [Sav16]. Discussion [CCS18]. Dispersed [WP07c, DSC^{+}05]. Distribution [Ber16, Bid12]. Diversity [Ber90]. Division [Har85]. Do [Sto98]. Doctoral [Sul84b, Web85]. Document [Sin80]. Documentation [Har88, HS78]. Documents [Mil89a, Mil89b]. Dodd [WP09j]. Doggett [Gre81]. Dokumentation [Due91]. Dolmetsch [Gre86]. Dolphin [Bro88b, Som86]. dominance [TM96]. Don [Kod84, WP07d]. Donald [Bro88a, New84, WP09a]. Doubleday [WP07n]. Doug [Joh09b]. Douglas [Ano18e, BY79, Swo18, Wel80]. Doves [Kel03, Tid02]. Dr. [BJS87]. Dragon [Pan88]. Drawings [WP08, Bec07]. Dreyfus [Far83a, Kel98, Win92, DG93]. Drogin [Nes84]. Dungeons [Pan88]. Duplicator [Zla13]. During [Bou12]. Dust [Mif17b, God16]. dust-jackets [God16]. Dwiggins [Ano19a, Pan85, Dwi80, Rum14]. Dyer [Sul84a]. Dyson [Wee85].

Eagle [Hel15]. earlier [Tho79, Wol80]. Earliest [Auj14, WP09j, Dod06, WB10]. Early [Bak92, BB05, CC85, Eis81, Har88, Hel17, Hud78, Lan04, Lee89, O’B03, Pan88, Per15, Hin82, Sil84, Tho79, Tra82, Wol80, Ame79, Eis94, HS78, TM96, Joh07d, Lar84, Mak19, Ste80a, Tur81, WP09d]. Early-Modern [Tra82, Eis94]. earnest [Wat01, Wat04]. Easton [Hey85]. Ed [Mif16b, Ant08, Bid82, Gob90, Hil80, Joy85, Kod10b, Lar92, Pet85, Pre81, Sil84, Sti85]. Ed. [Wei83]. Eden [Sen17, Keg16]. Edited [Ano91]. Edition [Win92, DG93, Cha90, Gob90]. Editions [Dan02, Fra89, Kel08c, TB88]. Editor [Leexxa, Panxx]. Editors [Cav88, Len88]. eds [Dav84, Kas82, Kod09a, Kod09b, Kod10c, Kod11, Sha18a, WP09a, WP09h]. Education [AP17, Hid07]. Educator [BJS87]. Edward [Ano72, HJ07, Kel08b, Tra79]. Edwin [Bia87, Gli84]. Egloffstein [Han93]. Egyptian [Tra94]. eight [Mid85]. Eighteenth [Dav84, CH83, HS78, Hel00]. eighteenth-century [HS78]. Eighty [Fra87]. Eisenstein [Tra82]. Elbert [Ano83]. Electric [Lab17]. Electrification [Dev85]. Elephants [Her09]. Eleven [Ano12a, Kel10]. Elizabeth [Bro79, Lie79, Thé83, Tra82, WP09b]. Ellen [Mif17a]. Ellms [Win11]. Emery [Dun88b]. Emma [Ano18b, Ano18a]. Encyclopedie [Met81, aD79]. End [Uza79]. England [Hun80b, Spu82, Kor83, Rai81]. English [Fra89, Tra79, Fal02, Hod78, McT81, TB88]. Engraved [Her09, Hun80b, Rai81]. Engraver [Tho79, WP07f, Joh07c, Pom05, Ugl06]. Engravers [O’B03, O’B14, Hun80a, Rai82]. Engraving [Lar84, O’B03, Tur81, Wee85, Dys84, Wol80]. Engravings [Har88, Lar84, Pom94]. Enlightenment [Met81, aD79, Sha18a]. Enquiry [Lar92, Pet85, BC83, CP34, CP71, CP92]. Enterprise [Bou08]. Entertainment [Bid82, Ste80c]. Ephemera [Bax00, Joh09c, Pan82, Sil82b, Hud08]. epilogue [CP92]. Era [Rom03, Gor14]. Eragny [WP08, Bec07]. Erasmus [Sch19]. Eric [DG93, Gil82]. Ernest [WP07j]. Esoteric [Sch17]. Essay [Geh98, La 15, FIbibBLL09]. Essays [Cav84, Sil84, Jan04, Mid85, MP82, Hin82]. Estiennes [Bid84, SB82]. Etherington [Kod84]. Europe [Day66, Eis94, Fra86c, Tra82]. European [Roy86, Smi88]. Europeans [Geh98]. Evenden [WP09b]. Everson [Pei98]. Every [Ste80b]. Ex [Ano11b, Hop11, Mif12b]. examined [BC83, Lar92, Pet85]. Excerpts [Tho79, LA94]. Exchange [You10b]. exhibit [Gil82]. Exhibition [BS99, Joh91, WP08, Ame79, Ann85, Ano72, Ano88a, Bec07, Leexxb]. Expansion [KR09, Kod09a]. Expedition [Bou08]. Experience [Mar07]. Experiments [Ste80b]. Exploration [Sul83, WCB82]. Exploring [Bou08]. Explosion [Joh10, Las05]. Extensive [Kod10c, GK10]. Externals [Wei83]. Eyes [Tho19].

F [FIbibBLL09, Kod09a, Kod11, WP07l]. Fable [Bri12]. Face [Pan99]. Facing [Pan07]. Fairchild [La 13]. Fall [Panxx, You10b]. Family [Ber90]. Fantasies [Rum15, Ruf17]. far [Sme03, Joh07b]. Fared [Smi86]. Farren [WP09a]. Fashioned [Sch81, TS68]. Feast [Tho19]. Feather [Tay90]. Feb [Ano72]. February [Bec07]. Fenn [Mif15, Rai13]. Fiction [Kor83, Spu82]. Field [Bou12, WP09i, Old06, Joh11b, You10a]. Fifteenth [Bar93, Fle86, Tra17, Bar85, Bar92, Bol16]. Fifteenth-Century [Bar93, Tra17, Bol16]. Fifty [WP09a, FIbibBLL09]. Fine [Auc01, Har98, Pei98]. Finlay [Kod10b]. Fiona [WP09h]. Firma [Rae88, Sch87]. First [Ano18b, Ano18a, Bon83, Han93, Joh11c, Phe85, Smi15, Tra79, WP09c, BL^{+}83, Hod78, Kel08c]. Five [WP09e, Tho79, Wil08, SIvA87]. Fleuron [Bro88b, Ash80, MS73, Som86]. Fleury [Sme08]. Flowering [WP09c, Kel08c]. Folio [Ano18a, Smi15, Ano18b, O’B14]. Folly [Sax80a, Com79]. Font [Dan02, Sch87]. fonts [Dod06, WP09j]. Forgeries [Lar92, Pet85, BC83]. Forgers [Lar92]. Formal [La 16b]. Forman [Lar92, Pet85, BC83]. Format [Nea79]. Forms [Smi05a, Smi05b, Tho06, WP07r]. Fortunes [Gra88, Bru86]. Founders [Ano80, Har81, Sax80b, Wil64]. Foundries [Hen95, Ann75]. Foundry [HF06, Eck87].

Four [Ano18a, Win92, DG93]. Fowle [All90]. Fragments [Sav16]. Frames [Hel11]. francaise [Cha90]. France [WP09d, BB05, Lan04]. Frances [Auj14]. Francis [Ash80, Hod78, Tra79]. Francisco [Har98]. Frank [Mif16a, Sax80a]. Frankfurt [Sch87]. Franklin [Fra88, New84, Sax90, Win07, Can00, For90, GS06, Sha07]. Frans [Joh08]. Fraser [Ano89]. Fraternal [Lan02]. Fred [Bid84, Joh07b]. Frederic [Ano79, Mif16d, Ano78, Bru90, La 13, La 16a, La 16b, Law90b, Lox08a, Lox08b]. Frederick [Han93]. Freedom [Joy85, RB84]. French [Joh09d, Spa83, Ame92, Joh11a, Ray82, Ver10]. Frisket [Sav16]. Fritz [Swa13]. Frost [Cor86, Blu85]. Frühdruck [Ame79, de 82]. Fruits [Bou08]. Fuller [Cos94]. Fulton [BP85]. Fünf [SIvA87, Hut89]. Future [Pan07, Wri07].

G [Lan83]. G. [Gle86]. Gabriel [Ruf17]. Galleries [Ano72]. Garamond [Arg92, Arg92, Lan05, WP07g]. Gary [Isa86]. Gatekeepers [Joy83, New84]. Gavin [Wee88]. Gay [Pan85]. Gender [Rum98]. Genial [Eck85]. Genius [Sax80a, Com79]. Geoffrey [Wee88, Tan82]. Geofroy [Sme08]. George [Bod93, Tra94, WP07h, Wal04]. Georgian [Gre09]. German [Gre86, Ame79, Caz84, Leu94, SIvA87, Sch87]. Germany [WP09d, BB05]. Giage [Lox08a]. Giancarlo [Geh09]. Giants [Rum92a]. Gibbings [DG93]. Gibson [Pre81]. Gill [DG93, Gil82, Mif16b, Pan85]. Giovanni [Mat82]. Glaister [Tan82, Gla79a, Gla79b, Tan82]. Glance [Joh11c]. Glick [Bia87]. Glorious [Bou08]. Glossary [Tan82, Gla79a, Gla79b]. Godburn [Mif17b]. Gods [Amo90]. Going [HH13]. Golden [Teb78, Win92, Cav82, DG93]. Golding [Sax81]. Goldman [Kos10]. Good [Dun79, Dun80, Far19, WP08, Bec07].

Goodman [Sax91]. Gordon [Spa83]. Gospel [CC85, Win92]. Gospels [DG93]. Goudy [Ano78, Bru90, HH13, La 13, La 16a, La 16b, Ano79, Law90b]. Government [CB18, Mil89a, Mil89b, Wal88]. Grabhorn [Bax00]. Graham [Lar92, Pet85, BC83, Joh09c]. GramLee [Har88]. Grand [Ben93]. Grandfather [Swa13]. Grandville [La 14, You12]. Granjon [Pal19a, Lan05]. Graphic [Ann85, Ano88a, Gar91, Gra90, Har85, HF06, RB08, Smi88, Szm86, Woo86, Roy86, SIvA87, Tho88, Tho06]. graphiques [Ano27, HF06]. Gravure [Wal88]. Gray [Van79, Bla72, Bla00, Gro01]. Great [Fli07, McT84, Mey83]. Greek [Fle86, Bar85, Bar92]. Green [Win07, Ber90]. Gregorio [Far83b, Kri82]. Gregynog [Pet87]. Grolier [Bec07]. Gross [Kod10c]. Groves [Kod09b]. Guard [Joh92]. Guide [WP09i, BW84, Old06, Wee88]. Gutenberg [Dod06, Smi88, WP09j, Mar07, Roy86, SFRM83, Sin80]. Gwasg [Pet87].

H

[Har03, Kod11, Lar92, Pet85, WP08, WP09k]. H. [BC83]. Hague [Pan85, Gil82]. Hala [Dal18]. Half [All83, HJB03, BL^{+}83, BB82]. Halftone [Han93]. Hall [Teb86]. Hallock [Coa03]. Hamel [WP07c]. Hammer [McG94]. Hampshire [Can00]. Hand [Lan83, Mat82, Rum92b, Sax86, Sta81, WP09i, HS85, MS80, Old06, OW80, SD91, SD92]. Handbook [Sta81, OW80]. Handmade [Mos07, Sax07]. handpress [HS78]. Hard [Ruf17, Rum15]. Harmsen [Joh07a]. Harris [Dav84, Lie79, WP07l]. Harry [Lar92, Mif15, Rai13]. Hartford [Kod10b, BF09, FB^{+}09]. Hartz [WP07i, LL06]. Harvard [Wat84, Ano88b, Hal86, Teb86]. Hauser [WP07d]. Hebrew [GK06, Hel00, Hel11, Hel15, Hel17, Kos10, Whi92]. Heiba [Ano18f]. Helen [Ano83]. Hench [Kas82].

Hendrik [Joh09d, Joh11a, Pal19a]. Henry [Har03, Ano89, Dun79, Sul83]. Her [All87]. Herbert [Ash80]. Hermann [Kel09b]. Highlights [WP07q, Ano03, BR03, Bid84, SB82]. Hill [Lox08a, WP09k]. Hindman [Sil84]. Histoire [Cha90]. Historian [Wel80, BY79]. Historical [Bum82, Fra86c]. Histories [Kor83, Spu82]. History [Ano18c, Bak92, Ber93, Bri12, CCS18, Cre06, Dre85, Eli15, Eri17, Far83a, Fea86, Fra86b, Geh98, Gre86, Har80, Har81, Hop09, Joh08, Joh92, KR09, Kel98, Kod09a, Kod09b, Kod10c, Kos10, Lan83, Met81, Mif12a, Nes84, Rom14, Ruf19, Sax81, Sul84b, Sul86, Tan81, Tan87, Tur81, Van84, WP07p, WP09j, Web85, Win95, Ano80, AN83, Bar78, Bar07, Bla83, Bla92, Caz84, aD79, Dod06, Dre81, Dre94, Dro83, Eas83, GL80, GK06, GK10, Hal86, HS78, HH06, Hus85, Jan04, KBKHF16, Lox04, Pea08, Ste85, Sut78, Teb78, TB88, Wol80, BBRD99, Mif16a, Teb79, Fra89, Joh09a, Pre81, Sha18b, Tay90, Teb86, WP07l, Ano94]. Hitherto [Bre90]. Hitler [Lan16]. Hodnett [Tra79]. Hoe [Com79, Sax80a]. Hofer [Ano88a, Joh91, Ano88a]. Holliday [Kel08b]. Hollywood [Sch18]. Holton [Ano12b]. Homan [Coa03]. Honor [Sil84, Hin82]. Hopkins [Mif16c]. Hopkinson [Ano11b, Mif12b]. Hornby [Har03]. Horse [Lan83]. Hospital [Nee86]. Houghton [Ano88a]. Hours [Zla13]. Housekeeper [Pan82]. Howard [WP09k]. Howard-Hill [WP09k]. Hubbard [Ano83]. Hudak [Ste80a]. Hudson [Joh09c]. Hugh [Coa03]. Hulse [Thé83]. Humane [LRS07, WP09g]. Humanist [Fle93, Hut93, Eli15]. Humanitarian [BJS87]. Hundred [Ant08, Zla13, Kos06]. Hundredth [Pan85, Dwi80]. Hunnisett [Rai81, Rai82]. Hunter [Sti85, Wri92]. Huntington [Dan02]. Huss [Due87]. Hutner [Bid92].

Ice [Ano19b]. Iconic [Ano18a]. Ideal [Cav84, Fle93, Pet81, MP82]. identification [TM96]. identity [BS^{+}98, BS99]. II [Auc01, Dun80, Eis94, Mar07, Mil89b]. III [Teb78, Har80]. Illuminated [Tan82, Gla79a, Gla79b]. Illustrated [Nea16, Sil84, Spa83, WP09d, WP09j, Bar78, BB05, Dod06, Ray82, Hin82]. Illustrating [Bec07, Can00, WP08]. Illustration [Rai81, SIvA87, Hod78, Hun80b, You12, La 14, Tra79]. illustrations [Rei75, Bro79]. Illustrator [Swa13, Pom05, Joh07c]. Illustrators [McT85, Mey83, PM84, McT84]. im [Ame79, de 82]. image [Hel00]. Imholtz [WP09a]. Importance [Mif12a, Pea08]. Imprints [Tra84, DM79]. Inclusion [Rum98]. Incunabula [Hel11]. Indeed [Far19]. Index [Ano94, Som86, Bro88b]. Indies [Cav88]. indulgence [Nee86]. Industrial [Fli07, BW84, Cas07]. Industry [Dev85, Sha18a]. Infant [Lee89]. Influential [Ano12a, Kel10]. Informal [Teb79, Sut78]. Ingenious [McC95, Swo18]. Initials [Bue82]. Ink [jK86, Pet81]. Inkmaking [jK86]. Inland [Tho79, Ann77]. Innocent [Bri05, WP07b]. Innovation [Ber90, Mif17a, Maz15]. Inspired [Ano18d, Sha17, Sha18b]. Instantaneous [HF18]. Institution [Har85, HS78]. Intarsia [Sch19]. Intellectual [Isa86, Mar85]. Interest [Tho79]. International [Fra89, You10b, BBB^{+}02, TB88]. Interview [Ano18e, Ano18f, Ano19a, Ano19b, Ano17]. Introduction [Lan83, Lar84, Law71, Leexxb, Panxx, Sil82b, Tra81, Whi92, Hen95, Due84, Kel91]. Inventing [Har86]. Invention [Can00]. Inventor [Han93, HF18, Sch89, Sil79a]. Invisible [CB18]. Involvement [Dre85]. Irene [WP07e]. Irish [Fal02, McG94]. Iron [Rum92b, SD91, SD92]. Isaac [Ste79b]. Island [Gre81, DM79, For90]. Issues [Ano94]. Italian [Pet07b]. Italic [Ame92, Wat84, HC81, Lan05]. Itinerant [Vor19]. IV [Dan02]. IV/V [Dan02].

J [Ano91, BC83, Bia87, Gra88, Lar92, Law90b, Pet85, Hin82, Sch17, Sil84, Tur81, WP07l]. J. [Bid79]. jackets [God16, Mif17b]. Jacobs [Rum14]. Jahrhunderte [Hut89, SIvA87]. Jam [Her09]. James [Lar92, Pan85, Sul84a, Ano89, Due84, Dye82, Joy85, Ruf19, Sax07, Win07]. Jan [Kel09a, Bur07, Dou06, WP07n]. Jane [Hey85, All87, Joh07c]. Janice [Kod09a]. Janson [Kri17]. Janssen [Joh08]. Jar [Her09]. Jargon [Vor19]. Jay [AP17, Joh10]. Jeanne [Bro88b]. Jeffers [Har87]. Jefferson [Sto98, Whi05]. Jeffrey [Kod09b]. Jenny [WP07f]. Jerry [Ano18c, Ano12a, Arg92, Sha18b, WP09c]. Jewish [Sin80]. Joe [Lis80]. Joel [WP07c, Cyr93]. Johann [Tra17, Bol16, Sin80]. John [Har03, Lar92, Pet85, Pre81, WP09b, BC83, Ber16, Blu80, Cav88, Dre93, DG93, Eve08, Far83a, Gre87, Han05, Har80, Har03, Joh07d, Joy85, Kel98, Lar92, McC95, Pan85, Pet85, RB84, Sha18a, Tay90, WP07g, WP09d, Win92]. Johnson [Tho80b]. Johnston [Kel08b, Ano72, HJ07]. Jonas [Bar83]. Jones [WP07h, Tra94, Wal04]. Jordan [Joh09b, Clo08]. Joseph [Kod10a, All83, Cor86, Kel08a, Kra79a, Sti90]. Josiah [Ste80b]. Journals [Tho79]. Journey [All83, Ann77, BB82]. Journeyman [Ano18e]. Journeymen [Dav81]. Joyce [New84]. Jr [Kod10a, WP09a]. Jr. [Kel08a]. Judah [Sin92]. Juliana [LL06, WP07i]. July [CH83, Dav84, Gil82].

Kaestle [Kod09a]. Kampf [Lan16]. Kate [Kod10a]. Kay [Dan02, Geh15]. Keeping [Sha07]. Kegan [Dun80]. Kegler [Sen17]. Keller [Kod10a]. Kelley [Kod10c]. Kellogg [Kod10b, BF09, FB^{+}09]. Kelly [Ano12a, Ano18c, Arg92, WP09c, Sha18b, Shi10]. Kelmscott [Ano12b, May87, Pet81, Pet84, PP11]. Kennerley [Gra88, Bru86]. Kennet [Ano19a]. Keri [La 14]. Kiesel [Sch17]. Kneeland [Yea12]. knigi [BL^{+}83, Gle86]. Knocks [Ruf17, Rum15]. Know [Sto98]. knows [Gre94]. Kono [Fra87]. Kosofsky [Ant08]. Kraus [Lis80]. Kredel [Swa13]. Krivatsy [Far83b]. Kseniya [Ano17].

L [Gre86, Joh09d, Joh11a, Mif16c, Pal19a, Sul83, Tra82, Wel80]. Labels [Her09]. Lamara [Ano18f]. Landesbibliothek [Ame79, de 82]. Landis [Sax93]. Landon [WP09g]. Lane [Joh07d, WP07g]. Lanes [Hau06, WP07d]. Language [AAG87, BBB^{+}02]. Lanston [Mif16c, Hop12]. Lapides [FIbibBLL09]. Larchmont [Ano79]. Large [Mak19]. Large-Scale [Mak19]. Last [Joh10]. Late [Joh09b, Mif17a, Pan88, Clo08, Maz15]. Late-Nineteenth-Century [Mif17a, Maz15]. later [Wol80]. Latest [WP09j, Dod06]. Latin [RB08, WP09h]. laureate [Wal04, WP07h]. Laurie [Pre81]. Lawrence [Tra94, WP07h]. Lawson [Kel91]. Lay [Chr82, Tra84]. Leadenhall [Joh11b, You10a]. Leader [Dev85, Tic89]. Leaf [Dav11, WP07c, DSC^{+}05]. Learned [Tra84, Chr82]. Learning [Teb79, Bar78]. Leary [Mif13]. Leaves [Gar91]. Lectures [Cav84, MP82, Tho06]. Legends [Sta10]. Legibility [Smi86, Ryd79, Blu80]. Leighton [Ash10]. Lent [Cav88]. Leo [WP07l]. Leona [Ste80a]. Leonard [AP17]. Leslie [BJS87]. Lessing [Sil84, Hin82]. Leti [Kri82, Far83b]. Letter [Ano91, La 15, La 16a, Pal19a, Gen84, Wil64]. Letter-Cutter [Pal19a]. Letterform [Ano19a, Car92]. Letterforms [Ame03, WP09j, Dod06]. Letterpress [Dun88b, Fli07, Hid07]. Letters [Har03, Nes80, WP09g, LRS07, Lox04, Mid85, Sti85, WP07p, Leexxa]. Levenson [Wal95]. Lever [Bre90]. Lewis [Pre81, Rum98]. libraries [BL^{+}83, Ano80, Har81]. Library [Ame79, Gil82, Har81, Nea79, Nea16, RB08, Ano80, Ano88a, SIvA87]. libri [Geh09, Pet07b]. Libris [Ano11b, Mif12b, Hop11]. Libro [Mar01]. Lieberman [Bid79]. Life [Ano19a, Isa86, Mif16d, Nes80, Ruf17, Rum14, Sam14, WP07e, WP07f, WP08, Bec07, Lox08b, Mar85, Mid85, Rum15, Tic05, Ugl06]. Lillian [HTBR05, Joh07a]. Lincoln [BJS87]. Linda [FIbibBLL09]. Linn [Cos94]. Linotype [Mif16a, WP07i, Gob98b, LL06, Rom14, Rum14, Sch89, Tra94]. Lion [Hel17]. Lions [Car01]. List [Mos08]. Literacy [Stu18]. Literary [EK83, Dar82]. Literature [Kel09a, Bur07]. Litho [Her09]. Lithographic [Wei89]. Lithography [Las05, Joh10]. Little [Sto98]. Lommen [WP07i]. London [Lea10, RB08, Wil64, Mif13]. Loney [Ano11a]. Long [DM79, Gre81]. Longer [Bri05, WP07b]. Longevity [Smi86]. Looks [Tho83]. Lord [BP85]. Lost [Cav82]. Love [Hel11]. Low [Tic05, WP07e, Lan04]. Loxley [Mif16d, WP07p]. Loya [Sch19]. Lucy [McT85]. Luppov [Gle86]. Lyon [Dav81].

M [Dav84, Har82, La 16b, Ste79b, Ste80a, Tra17]. M. [Dav84]. Machine [Can00, Ste79b, Tra94]. Machinery [Mak19]. Machines [Swo18]. MacKellar [Joh09b, Clo08]. Macmillan [WP09f]. Made [BBRD99, CB18, Phe85, Pom94, Sax91]. Madeleine [Bid82, Sul86]. Magazine [HF06, O’B14, How80]. Main [Sch87]. Making [Ano18b, Bid92, Smi15, WP07i, Hut90, LL06]. Malcolm [Sha18a]. Man [Bar83, Ben93, Ste80b, Sti85, HH13, Mid85, Eis81].

Manifesto [Joh07b, Sme03]. Manifestos [Pal19b]. Manley [Mar93]. manual [MW05, WP07k]. manufacture [TM96]. manufacturers [Old06, WP09i]. Manufactures [Lee89]. Manufacturing [Ber16, Bid12, WP07o, Sou05]. Manuscript [Wat84, HC81]. Manuscripts [Tan82, Gla79a, Gla79b]. Manutius [Fle86, Bar85, Bar92, Lan02]. Manutius/Miller [Lan02]. Maps [Woo86]. Marbling [Eas83, Hey85]. Marc [Nes84]. March [BS99]. Marching [Sil82b]. Mardersteig [Mat82]. Margaret [Kor83, WP09e]. Marguerite [Gre81]. Marianne [Kel03]. Marjorie [Ste79a]. Mark [Auj14, Gob98a, Mif17b, Sch81]. Marker [Isa86]. Marks [Joh07a, HTBR05]. Martin [Ano11b, Ant08, Bid92, Mif12b]. Mary [Kod10c, Nee86]. Mass [Bid82, Ste80c]. Master [Kel08b, Tra79, Ano72, Hod78, HJ07]. Masterpiece [Win92, DG93]. Masterpieces [SIvA87]. Masters [Tan99]. material [Lar92]. Materials [Mos08, Mos07, Sax07]. Mathematics [Pri13]. Mathieu [WP07i]. Matrices [Kri17, Van84, AN83]. Matrix [WP07q, Hus85, Due87, Ano03, BR03]. Matt [Kod84]. Matter [Bax00]. Matters [La 15, WP09e, Wil08]. Matthew [Gra88, Joh11b, Lan02]. Maureen [WP07j]. Maurice [Hen95]. Max [Teb86]. Maxim [Bod93]. May [BS99, Dun88b]. Mazur [Mif17a]. McKitterick [Hil80]. McLennan [Joh11b]. McMurtrie [Wel80, BY79]. Me [Fle09]. Mechanisms [Mak19]. Medallic [Ber93, Bla83, Bla92]. Medals [Wro91]. Medicine [Bar83]. Medieval [Mif16b, Nes84, Dro83, PS14]. Mein [Lan16]. Meisterwerke [Hut89, SIvA87]. Melbourne [Hau06, WP07d]. Melissa [Wel80]. Memento [Mar93]. Memoirs [Har82]. Memoriam [Ano88b, Ano89, BJS87]. Men [Pet07b].

Mergenthaler [Gob90, Gob98b, Sch89]. Meriden [Wal88]. Merrymount [Joh93, Kri17]. messenger [Ano65, Tho79]. Messerschmidt [Bly15]. Messrs. [Kra79b, Lis80]. Methods [Ber16, Bid12, Sou05, TM96, WP07o]. Métiers [HF06, Ano27]. Meyer [McT84]. Meynell [Ash80]. Michael [Kod09b, Kod11, WP07k]. Michalis [Pal19b]. Micklethwait [McT85]. Mid [Mif13, Lea10]. Mid-Victorian [Mif13, Lea10]. Middleton [Sti85, Mid85]. Mike [BBRD99]. Miller [Lan02]. Mills [Ber16, Bid12]. Mimeography [Vor19]. Minerva [Dav81]. Minnesota [HH06, WP07l]. Miriam [Tra84]. Mirror [Bri03, Hel00]. Mirror-image [Hel00]. Misha [Ano18c]. Misprint [Dav81]. Missing [Bax00]. Mitchell [Gra88, WP07k, Bru86]. modeled [Lan05]. Modern [Ben93, Nea79, Nea16, Tra82, Eis94, Gor14]. Modernism [WP07r, Smi05a, Smi05b]. Modernity [Dal18]. Moira [WP07l]. Monograms [Hel00]. Monotype [Mif16c, Hop12]. Monticello [Cre06]. Morely [La 15]. Moretus [Joh07d, Lan04]. Mori [Mar93]. Morison [Hil80]. Morphology [Sto87]. Morris [Cav84, Dun88a, Cos94, Dun88b, Pet81, Sta85]. Mosley [Sax07, Due84]. mostly [Lan04]. Motif [Hel15, Hel17]. Motion [KR09, Kod09a]. movement [Ano82, DeC84]. Mr. [Bre90, Sto98]. Multicolor [HF18]. Multilingual [Bod93, SZ91]. Multitudinous [HF18]. Mummy [Dan94, Sta10]. Munsell [Cyr93]. Museum [Har85, Sax93, Lan04, Joh07d]. Museums [Wri07]. Music [Tur81, Woo86, Wol80]. Musical [Kru86]. My [Hel11, Pan99, Ruf17, Rum15]. Mycall [McC95]. Mysteries [Joh92]. Mystery [Sav16]. Myth [Bas18].

N [Spa83, Win07]. Name [Pan99]. Nancy [Kod10b]. Naphtali [Sin92]. Nash [Har03]. Nathaniel [Kod10a, Kel08a]. Nati [Far83b]. Nation [Kod10c, Mil89a, Mil89b, GK10]. National [BS99, BS^{+}98]. Nature [Lar92, Pet85, Ugl06, BC83, CP34, CP71, CP92, WP07f]. nauchnykh [BL^{+}83, Gle86]. Needham [Dre87]. Needs [Amo90]. Neil [WP09f]. nel [Geh09, Pet07b]. Nemesis [Gob98a]. Newburyport [McC95]. Newport [For90]. Newspapers [Kod10a, Kel08a]. Nicholas [Lar92, Teb79]. Nicolas [Pet85, Fle86]. Nicolette [Van79]. Nilsson [Van84]. Nineteenth [Bid82, Can00, Dal18, God16, Gra76, Joh09b, Joh10, Lar92, Mak19, Mif17a, Pet85, Wee85, BC83, CP34, CP71, CP92, Clo08, Dys84, Maz15, Ste80c, Far19, Lan83, Las05, Mif17b, O’B03, Pet07a, Rum92a, Tho19, Van79, Win83, Win11]. Nineteenth-Century [Dal18, Can00, God16, Wee85, Dys84, Far19, Las05, Mif17b, O’B03, Pet07a, Rum92a, Tho19, Win11]. Ninety [Kel09b]. Nissenbaum [Kod09b]. No [BBRD99, Bri05, Rum98, Tic05, WP07b, WP07e]. Noah [Rus90]. Noblest [Ano18c, KBKHF16]. Non [RB08, WP09h]. Non-Latin [RB08, WP09h]. Nonesuch [Far83a, Dre81]. Nordlunde [Ben93]. Norstedt [Van84, AN83]. North [WP09i, Old06]. Northern [Wal95, LA94]. Northwest [Vor19]. Note [Dan02, Tho80a]. Notes [Bar76, Ber16, Bid12, Kod10a, Tan82, Due84, Gla79a, Gla79b, Kel08a, Lan04, Ste79a]. nouveau [Joh79, Tho80b]. Numismata [Bla83, Bla92, Ber93]. NY [Ano79].

O [Hen95, Rum92b, Wei02]. Observation [Fra86c]. observing [Gil82]. Oeuvre [Pal19a]. Office [Van84, AN83, CB18]. offices [Lan04]. Officina [Mat82, MS80]. Offspring [Lan02]. Oh [Sto98]. Old [Ame92, Ben93, Bue82, EK83, Sax94, Sch81, Sul84b, TS68, Dar82]. Old-Fashioned [Sch81]. Old-Style [Ame92]. Oldest [Sax91]. Oldham [WP09i]. One [Ant08, Fal02, Tho19, Zla13, Kos06]. Opentype [WP09j, Dod06]. Operate [Szm89, BF86]. Origin [Mif16c, Hop12, Hus85]. Origins [Ame92, Fli07, Har98, Isa86, Lab17, Sul86, Mar85, Ste85]. Ornamental [Hel11]. Ornamented [Van79, Gra76]. Ornaments [Bid80, Bro79, Rei75]. Orville [BB88]. Oscar [Dun88b]. Osley [Sta81]. Other [BB88, Pan99]. Ottmar [Gob90, Sch89]. Otto [Hut89, SIvA87]. Ouija [Eri17]. our [Amo90, HH06, WP07l, Lee89]. Own [Ste80b]. Owned [New84, Joy83]. Oxford [Kod11, Teb79, Bar78, SW10, Sut78].

P [Gle86, Joy85, Pan85, Sil79a, Wal82]. Pacific [Vor19]. Packages [Win11]. Page [Kru86, Mor81]. Pages [Hel00, Hel17]. Paige [Gob98a]. Paleo [Joh09d]. Paleo-Typography [Joh09d]. Palm [Dav11]. Pamphlets [Pet85, CP34, CP71, CP92, Lar92, Pet85, BC83]. Paper [Ber16, Bid12, Dan94, Mif15, Sta10, WP07l, HH06, Rai13]. Papermaking [Bar93, Tan82, Gla79a, Gla79b]. Papers [Dav84, Joh09d, BL^{+}83, CH83, Tra81]. pardoner [Nee86, Dre87]. Paris [Dan02, AB12, Geh15]. Parker [Sul84a, BBRD99, Dye82]. Parré [Sha18a]. Parson [Gre94]. Part [Car92, Mil89a, Mil89b]. Partial [Sil79a]. Partington [Mif16b]. Past [Bum82, Mil89a, Mil89b, Sha17, Ano18d, Sha18b]. Patents [Eve08, Gob98b, WP09b]. Patricia [Ber87]. Patrick [Mif13]. Patronage [WP09b, Eve08]. Paul [Ano18d, Bly15, Dre87, Dun80, Mar87, Sha18b]. Pearson [Mif12a]. Peasant [Leu94]. Peculiar [All90]. Peignot [HF06]. Penguin [WP07n, Dou06]. Peppin [McT85]. Perceptive [Dun80]. Permanence [Sil82b]. Perspective [Sto87]. pervoî [Gle86]. pervoi [BL^{+}83]. Peter [Joy85, RB84, Win07, Dun88a, Kel08b, Teb79, de 82]. Peterson [Ano12b, May87]. Petrella [Geh09]. Philadelphia [Har89, HH13, Lee89, O’B14]. Philip [Joh91, Ano88a, Ano91]. Philippic [Eck85]. Phoebe [Hey85]. Phonetic [HJB03]. Photography [Gra07, HB13]. Phototypesetting [Rom03]. Picasso [Roy86, Smi88]. Pichler [Pal19b]. Picture [Wee88, Wro91, Ben08, BW84, Ant10]. Pictures [Dys84, WP09b, Eve08, Wee85]. Picturing [BF09, FB^{+}09, Kod10b]. Pioneer [AP17, Len88]. Piracy [Pan99]. Pirates [Win11]. Pissarro [WP08]. Pissarros [Bec07]. Pitz [Ano89]. Place [All87]. Plains [Sul83, WCB82]. Plantin [Joh07a, Joh07d, HTBR05, Lan04]. Platemaking [Pan88]. Platen [Swo18]. Plates [Win83]. Pleasant [Kor83, Spu82]. plus [Joh07b, Sme03]. Pocket [Bon83]. Poems [Gro01, Bla72, Bla00]. Poland [Zla13]. Politics [Rum98]. Pollard [Lar92, Pet85, BC83]. Pollishings [Amo90]. polovine [BL^{+}83, Gle86]. Polyglot [Bod93, SZ91]. Pomeroy [Joh07c]. Popular [Kor83, Spu82]. Port [O’B14]. Portable [Bou12]. Portfolio [Wro91]. Portrait [BP85, Sch87]. Portraits [Ash10]. Porträt [Rae88, Sch87]. Portuguese [Szm86, Ann85]. Posters [Fli07]. Power [Lan83]. Practical [Reh93, Car94]. Practices [Tra17, Bol16, Wol80]. Prayer [Bid92, Hut90]. preliminary [Lan04]. Premier [WP07g, Lan05]. prepared [Gil82]. Present [Mil89a, Mil89b, Tay95]. Presented [Dav84, CH83]. Preservation [Sax07, Mos07]. Preservative [Win95]. Press [Ano79, BH80, Bou12, Bre90, Cav83, Dal18, Gre86, Har89, Joh07a, Kas82, Lan83, Mat82, Per15, Pet87, Sax80a, Sax91, Sax94, Sen17, Teb79, Teb86, Tra82, WP07j, WP08, WP09c, Win92, Bec07, Com79, Eis94, HS78, MS80, RB84, Ano19b, Bar78, Bax00, Bec07, Bue82, Car01, Cav82, Dre81, DG93, Fra86a, Gro01, Hal86, HTBR05, Joh93, Keg16, Kel08c, Lox08a, Mar07, Pet81, Pet84, Sax89, Sax90, Sax91, Sax93, Smi86, Sto05, Sut78, Tid02, Wat01, Wat04, Wei02, You10a, Far83a, Fra88, Joh11b, Joy85, Kel03, Lie79, May87, Tri84, WP07m, WP08]. Press-builders [Har89]. Presses [Har93, WP09i, Gla79a, Gla79b, Old06, Pet07b, SD91, SD92, SB82, Rum92b, Tan82, Bid84]. Pressmarks [Hel17]. Prince [RK91]. Print [Anoxx, KR09, Kel98, Kod09a, Kod10c, Mif13, Swo18, Wee85, Dre94, Dys84, GK10, Lea10]. Printed [Ant10, Auj14, Gil82, Hel00, Kra79a, Sch19, Sil82a, WP07l, Ben08, Blu77a, Blu77b, HH06, Nee86, Win81, WB10, Pan85]. Printer [All83, Ano18e, Bro79, Dav81, Dre87, Due87, Eis81, Fle93, Hel00, How80, Hut93, Lox08b, Mar93, Pan85, Pei98, Sil82b, Sou05, Ste80b, WP07h, Win07, Woo86, Yea12, Ann77, BB82, Bue83, Dye82, GS06, Hus85, Nee86, Rei75, Rum15, Tic05, TS68, Wal04, Mif16d, WP07o, You10b, Ruf17, Sul84a, WP07e, Sch81, Tho79]. Printers [Ant08, Ber90, BB88, Cav88, Har87, Hau06, Kel08a, Pre84, Ste79a, Ste80a, Thé83, WP07q, Ano03, Bar76, BR03, Blu85, Hud78, Hul82, Kod10a, Kos06, Sch17, WP07d, Cor86]. Printing [Ann85, Ano88a, AP17, Auc01, AN83, Ben93, Bou12, Bre90, CB18, CC85, CCS18, Dev85, Dre85, Due84, Dun88b, Fli07, Fra86b, Fra86c, Gar91, Gre81, HB13, Har98, Har81, Har85, Har86, Her09, Hop09, HF18, Isa86, Joh92, Joh09c, Joh11a, Joh11c, Kel98, Kos10, Lar84, Law71, Mak19, Mar07, McK79, Mil89a, Mil89b, Mor91, Per15, Pre81, Pri13, RK91, Sax80a, Sax86, Sax94, Sch18, Sil79b, Sul84b, Swo18, Szm86, Tan81, Tan82, Tan87, Tho19, Tra82, Tra17, Tur81, Van84, Wal88, Wal95, Web85, Wee88, Wei05, Win83, Win95, Wri07, Wro91, Zla13, Ame79, Ano80, Bla83, Bla92, Bol16, BW84, BY79, Com79, DM79, Dre94, Eis94, GL80, Gla79a, Gla79b, GK06, HS85, Hud08, Jan04, Lan04, LA94, Mar85, Tho79, Ver10]. printing [Wil64, Wol80, BBRD99, Dal18, Kel91, Ber93, Joh08, Wel80, Ano94, Sha18b]. Printings [Gor14]. Printmaking [BW84, Wee88]. Prints [BF09, Kod10b, WP08, Bec07, FB^{+}09]. Priority [Gob98b]. Prisoner [CB18]. Private [Cav83, Tan82, Tri84, Gla79a, Gla79b]. Pro [WP07g, Lan05]. Process [Han93, Pom94]. Processes [Pan88, RK91]. Producing [Mil89a, Mil89b, DG93]. production [Dre94, TM96, Kel98]. Products [Ber16, Bid12]. Profile [Fra87]. Project [Mar07]. Proof [Sax89]. Prospectus [Car01]. Psalm [Amo90]. Publication [Phe85]. Published [Lan16, Lar84, Ano19a, Sha18b]. Publisher [Dun80, Haw17, Mar87, WP07j, Yea12, Sre86, Wat01, Wat04, Ber87]. Publishers [Han00, Ste80a, Ste80c, Thé83, Hud78, Hul82, Bid82]. Publishing [Bou08, Gre87, Har80, KR09, Kod09a, La 15, Mar85, Met81, Mor91, New84, Tan82, Tur81, aD79, Gla79a, Gla79b, Joy83, Teb78, Teb87, Wol80, Isa86, Pal19b, Sch17]. Punch [Mif13, Sax86, HS85, Lea10]. Punchcutter [Coa03]. Punchcutting [Sax86]. punching [HS85]. Pup [Rus90]. Pursuit [HF18]. Push [Gob98b].

qualities [TM96]. Quality [Swo18]. Quest [Dav11, Swo18].

R [Cra85, Cra05, DeC87, Kod11, Law90b, Mif17b, Sax80a, Sul83, WP08, Wei83]. R. [Com79]. Rabbit [AB12, Geh15]. Radway [Kod09a]. Raecke [Fra86b]. Rafaeli [WP07a]. Rail [Har93]. Rainey [Mif15]. Ramage [Sax93, Sax89]. Rampant [Car01]. Rare [Her09]. Rathé [Dre93]. Raven [Ruf19]. Ray [Spa83]. re [BC83, Lar92, Pet85]. re-examined [BC83, Lar92, Pet85]. Readability [Ber11]. Readers [Gre94]. Readership [Kor83, Spu82]. Reading [KR09, Kod09a, Wil08, RB08, WP09e]. Rebuttal [BBRD99, Cav88]. Reconstruction [Sme08]. record [HS78]. Redesigning [Sta85, Dun88a]. Reflections [Joh93]. Regime [Dar82, EK83]. Regression [Swo18]. Rehak [Car94]. Reilly [Bro79]. Reincarnation [Pet87]. Related [WP08, Bec07, Tho79]. Religion [Yea12]. Reminiscence [Dan80]. Renaissance [Dan02, Geh15, Joh09d, Joh11a, AB12, Eis81, Gor14, Mor81, Pet07b, Sme08, Tan81, Tra81, Ver10]. Renate [Fra86b]. Representation [Leu94]. reprint [How80, Wei83]. reprinted [Uza79]. Reprinting [Har88]. Reproduction [Gra07]. Republic [Kod10c, GK10, Lar84]. Research [Sul84b, Tan87, Web85]. Resistance [Fal02]. response [Sha18b]. responsible [Gil82]. Retrospectus [Car01]. Reuben [Mar93]. Rev. [Wei02]. Reveal [Joh92]. Revew [Kor83]. Review [All83, Ano79, Ano11a, Ano11b, Ano12a, Ano12b, Ano18b, Ano18a, Ano18c, Ano18d, Ant08, Ant10, Arg92, Ash80, Auj14, Ber93, Ber16, Ber87, Bia87, Bid79, Bid82, Bid84, Bid92, Blu80, Bod93, Bro88a, Bro88b, Bro79, Car94, Cav84, Cor86, Dal18, Dav84, DeC84, DeC87, Dre87, Dre93, Due84, Due87, Due91, Dun88a, EK83, Far83a, Far83b, Fle86, Fra86b, Fra88, Fra89, Geh98, Geh09, Geh15, Gle86, Gob90, Gra88, Gra90, Gre81, Gre86, Gre87, Har80, Har81, Hen95, Hey85, Hil80, Hut89, Isa86, Joh91, Joh07a, Joh08, Joh07b, Joh07d, Joh07c, Joh09a, Joh09b, Joh09c, Joh09d, Joh10, Joh11a, Joh11b, Joy85, Kas82, Kel91, Kel98, Kel03, Kel08b, Kel09a, Kod84, Kod09a, Kod09b, Kod10a, Kod10b, Kod10c, Kod11, Kos10, Kra79a, La 14, Lar92, Law90b, Lie79]. Review [Lis80, Mat82, May87, McT84, McT85, Met81, Mif12a, Mif12b, Mif13, Mif15, Mif16a, Mif16b, Mif16c, Mif16d, Mif17a, Mif17b, Nes84, New84, Pal19a, Pal19b, Pan85, Pet85, Pre81, Rae88, Rai81, Rai82, Ruf17, Ruf19, Rum92b, Sax80a, Sax86, Sax07, Sch17, Sen17, Sha18a, Sil82a, Sil84, Smi88, Spa83, Sta81, Ste79a, Ste80a, Sti85, Sul83, Sul84a, Sul86, Sul87, Swo18, Szm86, Szm89, Tan82, Tan99, Tay90, Teb79, Teb86, Thé83, Tho80b, Tra79, Tra82, Tra84, Tra17, Tri84, Tur81, Van79, Van84, WP07a, WP07b, WP07c, WP07d, WP07e, WP07g, WP07f, WP07h, WP07k, WP07j, WP07l, WP07i, WP07o, WP07m, WP07n, WP07p, WP07q, WP07r, WP08, WP09a, WP09b, WP09k, WP09d, WP09c, WP09e, WP09f, WP09g, WP09j, WP09i, WP09h, Wal95, Wat84, Wee85, Wee88, Wel80, Win92, Win07, de 82]. review [Ano03, BR03, Sha18b]. reviews [Hel17, Mil89b]. Revision [Pan82]. Revisited [Cav88]. Revisiting [Amo90]. Revival [Auc01, Ber11, Sha17, Ano18d, Sha18b]. Revivals [Pet07a]. Revolution [BH80, Kas82]. Reynolds [WP07m]. RHM [Mid85]. Rhode [For90]. Richard [Tan99, Ant10, Due87, Mif16c, Ruf17, Sen17, Tur81, WP07o, WP07n, WP09g]. Richard-Gabriel [Ruf17]. Ricketts [WP07j, Wat01, Wat04]. Rinascimento [Pet07b, Geh09]. Rise [Gre87, La 14, Panxx, Sul87, You10b, Teb87, TM96, Wil84, You12]. Risks [Fli07]. Riverside [WP09c, Kel08c]. Rob [Shi10, WP09h]. Robert [BJS87, WP07g, BP85, Blu85, Cor86, DG93, EK83, Gen84, Gre86, Lan05, Mid85, Pal19a, Sav16, Sil82a, Sti85, WP09i]. Roberta [Kod10c]. Roberts [Kod84]. Robin [WP09j]. Robinson [Har87]. Rockies [Sul83, WCB82]. Roderick [Tri84]. Roger [Coa03, Wal95]. Rogers [WP09c, WP09g, Gen84, Gob98b, Kel08c, KBKHF16, LRS07, Tho80a, Tho83, Ano18c]. Role [Gra07]. Rollins [Wal82, Pan85]. Rollo [Lan83]. Roman [Ame92, Ano18c, KBKHF16, Lan05, Sme08]. Romano [Mif16a]. Romantic [DeC87, Kru86, Cra85, Cra05]. Rosebud [Har88]. Rosenwald [Hin82, Sil84]. Ross [WP09h]. Rounceval [Nee86]. Roundtable [CCS18]. Roy [Shi10]. Royal [Van84, Ano72, AN83]. Roylance [Smi88]. Rudge [Gli84, Bia87]. Ruggles [Sil79a]. Rummonds [Ruf17]. Russia [Isa86, Mar85]. Russian [BL^{+}83]. Russkie [BL^{+}83, Gle86]. Ryder [Blu80].

S [Ano12b, Gle86, Kel91, May87, Sta81, Wal88]. S.J [Kod11]. Sadek [Bod93]. Samuel [Yea12]. San [Har98]. Sandra [Sil84]. Sanford [Ano83]. Saul [Joh07a, HTBR05]. Saving [Hop09]. Saxe [Hen95, Rum92b]. sbornik [BL^{+}83, Gle86]. Sbrebrnik [Ber87]. Scale [Mak19]. Schaefer [SIvA87]. Schäfer [SIvA87, Hut89]. Schlesinger [Gob90]. Scholar [Fle93, Hut93]. Scholar-Printer [Fle93]. School [Her09]. Schreiber [Bid84, Dan02]. Schreyer [Sil82b]. Schriftgeissen [Due91]. Schriftgiessen [Rae88, Sch87]. scientific [BL^{+}83]. Scott [Ant08, Kod09b, Wel80]. Scott-Martin [Ant08]. Scoundrel [Haw17]. Scrappy [Sav16]. Screenplay [Sch18]. Scribes [Nes84, OW80, Dro83, Sta81]. Scribner [How80]. Scribner’s [Uza79]. Scripps [Mar07]. Script [Fle86, Gre09, HJB03, Bar85, Bar92]. Scythe [AB12, Geh15]. Sean [Sul87]. Search [Pet81]. Searching [CH83, Dav84]. Second [Auc01, CB18]. Secret [WP07p, Lox04]. Selected [Hil80, Joh09d, Kel98, Mac79, Dre94, LA94]. Selling [Far19]. Sem [LL06, WP07i]. Separately [Lar84, Sil82a, Win81]. Sequel [Lar92, Pet85, BC83]. Series [Fra86b, Nea79, Nea16]. Service [Sax80a, Com79]. set [Smi05b]. Settlement [Gra07]. Seven [Har03]. Seventeenth [Kor83, Hel00, Spu82]. Seventeenth-Century [Kor83, Spu82]. Seventeeth [Hel15]. Seventeeth-Century [Hel15]. Shakespeare [Ano18b, Ano18a, Smi15]. Shaw [Ano18d, Sha18b]. Sheets [Sav16]. Shenandoah [Gre86]. Sherman [La 13]. Sherwin [La 16a]. Shipwrecks [Win11]. Shoddimites [Dun79]. Shopfloor [Rum98]. Shops [Sil79b]. Short [CH83, Dav84, La 15, Rum14]. Shove [Gob98b]. Shown [Joh92]. Signed [La 16a]. Silicon [AAG87]. Silk [Mar93]. Silver [WP07c, Lan83]. Simon [Ash80, Geh15, Ame92, AB12, Mif16d, WP07p]. Since [Auc01]. Singer [Ste79b]. Sir [Ash80]. Sisson [Lie79]. Sisters [Cav88, Len88]. Sixteenth [Arg92, Joh09d, Sta81, OW80, Dav81, Hel15]. Sixteenth- [Hel15]. Sixteenth-Century [Arg92, Joh09d]. Sketch [McK79]. Sled [Har88]. Slimbach [WP07g]. Small [Sax94, Spu82, Kor83]. Smeijers [Joh07b]. Smith [Ano18b, Ano18a, WP07r, Ano89]. Smiths [Joh09b, Clo08]. Smithsonian [HS78, Har85, Wri92]. Social [Gre86, Lan83, Leu94, Tra84, Caz84, Chr82]. socialist [Cra85, Cra05, DeC87]. Society [Ant08, Kod10c, Mor91, GK10, Kos06, Mak19]. Some [Ano78, Bas18, BBRD99, Bid80, Bue82, Pre81, GL80, Ano79]. Somers [Bro88b]. Sorts [Ano79]. Sources [Bue82, Sta81, OW80]. South [Dav11]. Southall [WP07o]. Southern [Gre94]. Southwest [Ame79]. SP [Ant08, Kos06]. Spaces [Pri12]. Spanish [Ann85, Szm86]. Speak [Fal02]. Specimen [Anoxx, Due84, You10b, Wil64]. Specimens [Joh92, Joh07d, Lan04]. Speculator [Bar83]. Speech [Sch81]. Speed [Rum92a]. Spiritual [Pet07a, Mak19]. Spiritualist [Haw17]. Spread [Teb79, Whi92, Bar78]. Spufford [Kor83]. Sr [Kod10a, Kel08a]. St [Har03, Nee86, RB08]. Stallybrass [Win07].

Stanhope [BP85]. Stanley [Hil80]. Stansbury [Wei02]. Stansky [Dun88a]. Starling [Par94, BBRD99, BBRD99]. State [Ame79]. States [Day66, Har80, Kod09a, New84, Ste79a, Sul86, Bar76, Joy83, KR09, Lan16, Mil89a, Mil89b, Ste85, Teb78, Wal88, Whi92, Win83]. Stationer [Sin92]. Stationery [CB18]. STC [Pan82]. Steam [Lan83]. Steel [Hun80b, Rai82, Hun80a, Rai81]. Steel-engraved [Hun80b, Rai81]. Steele [Sav16]. Stempel [Sch87, Rae88]. Stephen [Hen95, Joh07a, Kod09b, Ano88b, Rum92b, Sil79a, Wat84]. Steps [SFRM83]. Stereotyping [Pom94]. Stern [Bid82, Sul86]. Stevens [Dun79]. Stichbild [Kru86]. Sticks [GL80, Pre81]. Still [Pan99, Sam14]. Stockdale [Cav88, Len88]. Stoddard [Coa03]. Stone [Mar93, WP07m, Wei89]. Stones [Pre81, GL80]. Story [Auj14, HH13, WB10]. Strahan [Ber87, Sre86]. Strasbourg [Tra84, Chr82]. Strategies [Rum98]. Streams [Swo18]. Streets [Hau06, WP07d]. Studies [Van84, AN83]. Study [Ano12a, Bod93, Kel10, SZ91, TM96]. Stuttgart [de 82, Ame79]. Style [Ame92]. Suarez [Kod11]. Subject [Har81, Ano80]. Subscription [Far19]. Südwesten [Ame79, de 82]. Sue [Mif15]. suit [Gre94]. Suppositions [BBRD99]. Surfaces [Her09]. Surviving [Mos08]. Susan [WP07k, McT84]. Sutcliffe [Teb79]. Sutra [Auj14, WB10]. Sylvia [Ano12b]. Symposium [CH83, Dav84]. Symth [Mif16b].

T [Joh10, Kod84, Lar92, Pet85]. T. [BC83, WP09k]. Table [Mif13, Lea10]. Tabor [Joh07a]. Tactile [Stu18]. Takashi [Fra87]. Tale [Pan99]. Talk [Mif13, Lea10]. Taste [Eck85]. Tastemakers [Joh09b, Clo08]. Tauchnitz [Fra89, TB88]. Teaching [CCS18, Her09, Law84]. Tebbel [Gre87, Har80]. Technique [Jan04, Joh08]. Techniques [Bid12, BW84, Ber16]. Technological [Fal02]. Technology [Lan83, Pre84]. Temper [Tay95]. Terminology [Dan02, ER82, Kod84]. Terms [Tan82, Gla79a, Gla79b]. texts [Bla00, Pea08, Mif12a]. Their [Hen95, Mif12a, WP09i, Ann75, Bid84, Old06, Pea08, SB82]. Them [Joh92]. Theo [Car94]. Theobald [Mar87]. Theodore [WP07e, Tic89, Tic05]. Theories [Car92]. These [Sto98]. Thin [Ano19b]. Thomas [Ber87, Gro01, Lar92, WP07f, Ano17, Bak92, Bla72, Bla00, Ugl06, Whi05]. Thompson [Gra90]. Thomson [Mif17a]. Thoughts [Mos07, Sax07, Tan87]. Thousand [Zla13]. Tibetan [AAG87]. Tichenor [WP07e]. Tidcombe [Kel03]. Tiller [Ano82, DeC84]. Time [Rum92a, Sha07, Sto98]. Tints [HF18]. Title [CH83, Dav84, Hel00, Hel17, Mor81]. Todd [Bak92, Fra89]. Tolbert [Hop12, Mif16c]. Tomlinson [Tan99]. torchi [Geh09, Pet07b]. Toronto [Thé83, Hul82]. Torre [Pan85]. Tory [Sme08]. Tracts [Mif17a, Maz15]. Trade [Ber16, Bid12, CC85, Gre86, Lan83, O’B03, Pre84, WP09k, Caz84, Dys84, Eve08, HH08, TM96, WP09b, Wee85]. Trades [Thé83, Hul82]. Tradesman [Law90c, Kel91]. Tradition [McT81]. Traditions [Har98]. Tramp [Pre84]. Transatlantic [Dre85]. Transcribed [La 15]. transcription [Sil79a]. Transformation [Gre87, Tra82, Eis94, Teb87]. translated [Ano91]. Transmission [AAG87]. Travel [Sul83, WCB82]. Treasury [McT84, Mey83]. Trianon [Gro01]. Tribute [Dwi80, La 16b, Pan85, Coa03]. trudov [BL^{+}83, Gle86]. Tschichold [Kel09a, Bur07, Dou06]. Tschicold [WP07n]. Tudor [WP09b, Eve08]. Tuer [Joh11b, You10a]. Twain [Gob98a, Sch81]. Twentieth [Ano12a, McT85, WP07o, Kel10, PM84, Sou05]. Twenty [Bel95, Zla13]. Two [Ash10, Dun79, Dun80, Lar92, Smi86, Teb78]. Type [Ann75, Ano80, Ano03, Ano18d, Anoxx, AN83, BS99, BBRD99, BR03, Eck87, Eli15, Fle86, Har81, HF06, Joh92, Joh07d, Kri17, Lie78, Lox04, Nes80, Pan99, Par94, Phe85, Sax80b, Sax86, Sax07, Sha17, Sha18b, Shi10, Sme03, Ste79b, Tra94, WP07o, WP09f, WP09j, Whi92, BS^{+}98, Bar85, Bar92, BBB^{+}02, Dod06, HS85, Lan04, Mac06, Mos07, Sou05, Bid79, Hen95, Joh07b, Van84, WP07p, Arg92, WP07q, Anoxx]. Typecasting [Reh93, Car94, Sax86]. Typeface [Cre06, Law90a, Kel91]. Typefaces [Ano79, Ano18d, La 13, Sha17, Sha18b, Ano78, Gra76, Lie78, RB08, Bid79, Joh09d, Van79, WP09h]. Typeforms [Bar07, Joh09a]. Typefounding [Lee89, Mos08, Rae88]. typefoundries [Lan04]. Typefoundry [Van84, AN83]. Typemaking [Cos94]. Types [Ame92, Ano18c, Joh11a, Kel91, Law71, Smi86, KBKHF16, Lan05, Ver10, Wil64, Due84]. Typesetter [Lab17]. typesetting [Hop12, SZ91, Bod93]. Typograph [Gob98b]. Typographer [McC95]. Typographic [Ann77, Ano65, BB82, Joh09b, Pet07a, Sax07, Tho79, Woo86, Clo08, Mos07, All83]. Typographica [Ber93, Bla83, Bla92]. Typographical [Rum98, Ste80b, Win92, DG93]. Typography [Gre09, Joh09d, Kel98, Lab17, Law84, Raf05, RB08, SZ91, WP07k, WP07q, WP07r, Ano03, BR03, Bur07, Day66, Dre94, MW05, Smi05a, Smi05b, Bod93, Kel09a, WP07a]. Typophile [Dre93, Rat92]. Typothetae [Sch81]. Typsetting [Mif16c]. Tyrus [Joh07a].

U. [Wal88]. U.S. [CB18, Bou08]. UCLA [Fle09, Gil82]. Uglo [WP07f]. Ulm [Tra17, Ame79, Bol16, de 82]. Uncial [HJB03]. Unconventional [Bas18]. Underground [EK83, Fra86c, Zla13, Dar82]. Unfolded [Pan07]. Unicode [BBB^{+}02]. Union [Rum98, Eck87]. Unions [Pre84]. United [Har80, Kod09a, New84, Ste79a, Sul86, Win83, Bar76, Day66, Joy83, KR09, Lan16, Mil89a, Mil89b, Ste85, Teb78, Wal88, Whi92]. Unity [WP07r, Smi05a, Smi05b]. University [Ano80, Bar78, Hal86, Har81, RB08, Sut78, Teb79, Teb86]. Unlikely [Tic89]. Unlost [Ano79, Ano78]. unrecorded [Nee86]. Unreported [Bre90]. Untangling [Sav16]. Uomini [Pet07b, Geh09]. Updike [Hil80, Hut93]. Use [Hel11, Phe85, TM96]. Used [Tan82, Gla79a, Gla79b]. Usher [Tra84].

V [Gle86, Gre81, BL^{+}83, Dan02]. Vale [WP07j, Wat01, Wat04]. Valley [Sax93, Gre86]. various [Bid84, SB82]. veka [BL^{+}83, Gle86]. Venice [Fle93]. Verlag [Fra86b]. Versus [Gob98b]. Vervliet [Joh09d, Joh11a, Pal19a]. Very [Far19]. Vesalius [Gor14]. Victor [McG94]. Victorian [Ber87, Kod10b, Mif13, WP09d, BF09, BB05, Dun79, Dun80, FB^{+}09, Lea10, Sre86]. Video [Szm89]. Videotape [Sax86]. View [Pan82]. Vignettes [Bum82]. Vincent [Pan85]. Vinne [Smi86, Tic89, Tic05]. Violent [Sil79b]. Virginia [Gre94, WP07r]. Visible [CB18, Pan07]. Vision [Mar87]. Visual [Leu94, Smi05b]. Voice [Bri03]. Voices [Dun79, Dun80]. vol [Kod09b]. Vol. [Eis94]. Volmer [Ben93]. Volume [Har80, KR09, Kod09a, Kod10c, Ame79, GK10, Teb78]. Vom [Sch87, Rae88].

W [Ano78, Ano79, Ano83, Kod09b, La 16a, La 16b, Lis80, Pan85, Swo18, Tra94, WP07h, Wal04, Wei83]. W. [Ano83, Dwi80, Par94, Wil64, Ano19a]. Wagner [Sul83]. Wakeman [Wee88].

Walker [Pan85, Dun88b]. Wallis [WP07h, Tra94]. Walter [Due91]. War [Auc01, CB18, Bou12, Bum82, CB18, Caz84, Gre86]. Warde [Lox08a, Lox08b, Mif16d]. Warren [Ste80b]. wars [Teb78]. Was [Bly15, Rus90, Har88]. Watch [Lox08a]. Water [Lan83, Bla72, Bla00, Gro01]. water-colour [Bla72, Gro01]. water-colours [Bla00]. Watermarks [Ber16, Bid12]. Watry [WP07j]. Watson [Ash10]. Way [Han00]. Weems [Gre94]. Weimerskirch [Ano91]. West [Sul83, Cav88, WCB82]. Western [Bre90]. White [You10a, Kel08a, Kod10a]. Whittingham [Pan85, Bue83]. Who [Bly15, O’B03]. Widow [Ano91, For90]. Wightman [WP07k]. Wilbur [BB88]. Wilde [Dun88b]. Wilentz [Sul87]. Wilhelm [Han93]. Wilkes [Due91]. will [Gre94]. Willes [WP09e]. William [Dun88a, Ano12b, Ber93, Bia87, Bla72, Cav84, Due84, Dwi80, Fra89, Gli84, Gro01, Har82, Haw17, May87, Nee86, Pei98, Pet81, Sch17, Sta85, Tan99]. Williams [Han00]. Winans [Sil82a]. Winchester [Bar83]. Wining [Sil82b]. Winkle [Kod10a]. Winship [Kod09b]. Winter [Sha18b]. Winterbottom [TM96]. Wise [Lar92, Pet85, BC83, Lar92]. without [Tic05, WP07e]. Wolfe [Tur81]. Women [Bas18, Cav88, Len88, LA94, Ste79a, Ste80a, Bar76, Hud78, Rum98, Wal95]. Wood [Auj14, Har88, Her09, Joh07c, Pom94, Shi10, Pom05]. Wood-Engraved [Her09]. Wood-engraver [Joh07c]. Woodblock [AAG87, McT81]. Woodblocks [Gor14]. Woodburytype [Han05]. Woodcut [Bue82]. Woodcuts [Leu94]. Wool [CC85]. Work [For90, Joh07b, Mat82, Mif16d, WP08, Bec07, Lox08b, MS80, Sme03]. Working [Sul87, Wil84]. workman [Gil82]. Works [Far83b, Pan85, Kri82, Wal82]. World [Auj14, CB18, Dun88a, Mif15, Sha07, Rai13, Sta85, Auc01, WB10]. Woudhuysen [Kod11]. Wright [BB88]. writer [GS06, Win07]. Writing [CB18]. Writings [Kel98, Dre94]. wrought [HH13]. Wuerttemberg [Ame79]. Württembergischen [Ame79, de 82].

XIII [BL^{+}83]. XIX [Gle86, BL^{+}83]. XVI [Gle86, BL^{+}83]. XVI-pervoi [BL^{+}83].

Years [All83, Ant08, Bel95, Cav82, Sto98, BB82, Dou06, Kos06, WP07n]. Yetsweirt [All87]. York [Sax80b, Sul87, Wil84, Keg16, Sin92, Sul84b, Sen17]. Yosef [Kos10]. Young [Joh11b, Wel80]. Yousi [La 14]. Youth [Tho80a].

Z [WP09f, Mac06]. Zainer [Tra17, Bol16]. Zapf [Kel09b]. Zealand [CC85]. Zenger [Joy85, RB84]. Zhukov [Bod93].

References

Alterman:1987:WSC

[AAG87] Benjamin Alterman, Deborah Alterman, and Laura L. Gewissler. From woodblock to silicon chip: The transmission of Tibetan language. Printing History, 9(1):??, ???? 1987.

Amert:2012:SRS

[AB12] Kay Amert and Robert Bringhurst. The Scythe and the Rabbit: Simon de Colines and the culture of the book in Renaissance Paris. Cary Graphic Arts Press, Rochester, NY, USA, 2012. ISBN 1-933360-56-9. 288 pp. LCCN Z232.C69 A44 2012.

Darnton:1979:BEP

[aD79] Robert autor Darnton. The business of enlightenment a publishing history of the Encyclopédie, 1775–1800. Belknap Press of Harvard University Press, Cambridge, MA, USA; London, UK, 1979. ISBN 0-674-08785-2. xiv + 624 pp. LCCN AE25.E6 D37.

Allen:1983:BRJ

[All83]  Sue Allen.  Book review:

Joseph Blumenthal, Typographic Years: A Printer’s Journey Through A Half Century, 1925–1975. Printing History, 5(2):??, ???? 1983. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-10-
0.

Allen:1987:JYC

[All87] Susan M. Allen. Jane Yetsweirt (1541–?): Claiming her place. Printing History, 9(2):??, ???? 1987.

Allen:1990:PDD

[All90] Susan M. Allen. The peculiar defense of Daniel Fowle. Printing History, 12(1):??, ???? 1990.

Amelung:1979:FID

[Ame79] Peter Amelung. Der Frühdruck im deutschen Südwesten, 1473-1500: eine Ausstellung der Württembergischen Landesbibliothek Stuttgart. (German) [On early printing in the German Southwest 1473– 1500: An exhibition of the Wuerttemberg State Library of Stuttgart. Volume 1: Ulm].

Württembergische Landesbibliothek, Stuttgart, Germany, 1979. ISBN 3-7772-7929-3 (vol. 1). ???? pp. LCCN Z240 .A48.

Amert:1991:OFO

[Ame92] Kay Amert. Origins of the French old-style: The roman and italic types of Simon de Colines. Printing History, 13/14 (2/1):??, ???? 1991–1992.

Amert:2003:DCL

[Ame03] Kay Amert. Digital comparison of letterforms. Printing History, 23(2):??, ???? 2003. ISSN 0192-9275.

                                        URL https://www.rit.edu/
                                        press/printing-history-46-
                                        0.

Amory:1990:GAN

[Amo90] Hugh Amory. ‘Gods altar needs not our pollishings’: Revisiting the Bay Psalm Book. Printing History, 13(2):??, ???? 1990.

Axel-Nilsson:1983:TSN

[AN83] Christian Axel-Nilsson. Type studies: the Norstedt collection of matrices in the typefoundry of the Royal Printing Office: a history and catalogue. Norstedts tryckeri, Stockholm, Sweden, 1983. ISBN 91-22-00613-3. 198 + 3 pp. LCCN Z250.A3 A9 1983. Translation from Swedish original by Roger C. Tanner.

Annenberg:1975:TFA

[Ann75]  Maurice Annenberg.  Type

foundries of America and their catalogs. Oak Knoll Press, New Castle, DE, USA, 1975. ISBN 1-884718-06-X. xxi + 286 pp. LCCN Z250.A1 A56 1994.

Annenberg:1977:TJT

[Ann77] Maurice Annenberg. A Typographic journey through the Inland printer, 1883–1900. Maran Press, Baltimore, MD, USA, 1977. ISBN 0-916526-04-6. xx + 731 pp. LCCN Z208 .T93.

Anninger:1985:SPC

[Ann85] Anne Anninger. Spanish and Portuguese 16th Century books in the Department of Printing and Graphic Arts: a description of an exhibition and a bibliographical catalogue of the collection. Houghton Library, Cambridge, MA, USA, 1985. x + 94 pp. LCCN Z1014 .H29 1985.

Anonymous:1865:TMH

[Ano65] Anonymous. Typographic messenger. J. Conner’s Sons, Publishers, New York, NY, USA, 1865–1865. LCCN Z119 .T98.

Anonymous:1885:AB

[Ano97] Anonymous. The American bookmaker. H. Lockwood and Co., New York, NY, USA, 1885–1897. ???? pp. LCCN Z119 .P94.

Anonymous:1927:AMG

[Ano27] Anonymous. Arts et métiers graphiques. Hachard et cie., Paris, France, 1927. ISSN 1278-1452. ???? pp. LCCN Z119 .A84.

Anonymous:1972:EJM

[Ano72] Anonymous. Edward Johnston, master calligrapher: centenary exhibition, 1872–1944, Royal College of Art Galleries, 3–25 Feb. 1972. British Library, London, UK, 1972. ISBN 0-7123-4927-8 (British Library), 1-58456-198-X (Oak Knoll Press). xx + 389 pp. LCCN Z43.42.J64 H65 2007.

Anonymous:1978:SUT

[Ano78] Anonymous. Some unlost typefaces of Frederic W. Goudy. Out of Sorts Letter Foundery, Larchmont, NY, USA, 1978. ???? pp. LCCN ????

Anonymous:1979:BRS

[Ano79] Anonymous. Book review: Some Unlost Typefaces of Frederic W. Goudy, Out of Sorts Press (Larchmont, NY). Printing History, 1(2):??, ???? 1979. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-2-

0. Originally printed as a keepsake for the Goudy Society Annual Dinner in 1978.

Anonymous:1980:HPB

[Ano80] Anonymous. The history of printing from its beginnings to 1930: the subject catalogue of the American Type Founders Company library in the Columbia University Libraries. Kraus International Publications, Millwood, NY, USA, 1980. ISBN 0-527-18763-1 (set). xxxvi + 2744 (4 volumes) pp. LCCN Z117 .C65 1980. US$295.00 (prepublication). With an introduction by Kenneth A. Lohf.

Anonymous:1982:TBD

[Ano82] Anonymous. Tiller: a bimonthly devoted to the Arts and Crafts movement. The Artsman, Bryn Mawr, PA, USA, 1982. ???? pp. LCCN N1 .T56.

Anonymous:1983:EHW

[Ano83] Anonymous. Elbert Hubbard to W. W. Denslow from The Sanford & Helen Berger Collection. Printing History, 5(1):??, ???? 1983. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-9-
0.

Anonymous:1988:CEP

[Ano88a] Anonymous. A catalogue of an exhibition of the Philip Hofer bequest in the Department of Printing and Graphic Arts: Philip Hofer Collection in the Houghton Library: Philip Hofer bequest. Harvard College Library, Cambridge, MA, USA, 1988. ISBN 0-914630-03-2. xiv + 218 pp. LCCN Z116.A3 H37 1988.

Anonymous:1988:MSH

[Ano88b] Anonymous. In memoriam Stephen Harvard. Printing History, 10(2):??, ???? 1988.

Anonymous:1989:MBB

[Ano89]  Anonymous.  In memo-

riam Bernard Brussel-Smith.

With contributions by James Fraser, Henry Pitz and Bernard Brussel-Smith. Printing History, 11(2):??, ???? 1989.

Anonymous:1991:LBW

[Ano91] Anonymous. A letter from Bodoni’s widow. (Edited and translated by Philip J. Weimerskirch). Printing History, 14 (1):??, ???? 1991.

Anonymous:1994:IPH

[Ano94] Anonymous. An index to Printing History, issues 1–

32. Printing History, 17 (2):??, ???? 1994. ISSN 0192-9275. URL https://

d3eb7xfyht02et.cloudfront.
net/wp-content/uploads/2011/
12/apha-ph-os-index1.pdfhttps:
//www.rit.edu/press/printing-
history-34-0.

Anonymous:2003:TTH

[Ano03] Anonymous. Type and typography: highlights from Matrix, the review for printers and bibliophiles. Mark Batty Publisher, West New York, NJ, USA, 2003. ISBN 0-9715687-6-6, 0-9724240-8-3 (special edition). xx + 374 pp. LCCN Z250.A2 T95 2003.

Anonymous:2011:BRA

[Ano11a] Anonymous. Book review: Alan Loney, The Books to Come. Printing History (New Series), (10):??, July 2011. ISSN 0192-9275.

Anonymous:2011:BRM

[Ano11b] Anonymous. Book review:

Martin Hopkinson, Ex Libris: The Art of Bookplates. Printing History (New Series), (10): ??, July 2011. ISSN 0192-9275.

Anonymous:2012:BRJ

[Ano12a] Anonymous. Book review:

Jerry Kelly, The Art of the Book in the Twentieth Century: A Study of Eleven Influential Book Designers from 1900 to 2000. Printing History (New Series), (11):??, January 2012. ISSN 0192-9275.

Anonymous:2012:BRW

[Ano12b] Anonymous. Book review:

William S. Peterson and Sylvia Holton Peterson, The Kelmscott Chaucer: A Census. Printing History (New Series), (11):??, January 2012. ISSN 0192-9275.

Anonymous:2017:IKT

[Ano17] Anonymous. An interview with Kseniya Thomas. Printing History (New Series), (21):??, Winter 2017. ISSN 0192-9275.

Anonymous:2018:BREb

[Ano18a] Anonymous. Book review:

Emma Smith, Shakespeare’s First Folio: Four Centuries of an Iconic Book. Printing History (New Series), (23):??, Winter 2018. ISSN 0192-9275.

Anonymous:2018:BREa

[Ano18b] Anonymous. Book review:

Emma Smith, The Making of Shakespeare’s First Folio. Printing History (New Series), (23):??, Winter 2018. ISSN 0192-9275.

Anonymous:2018:BRM

[Ano18c] Anonymous. Book review: Misha Beletsky and Jerry Kelly, The Noblest Roman: A History of the Centaur Types of Bruce Rogers. Printing History (New Series), (23):??, Winter 2018. ISSN 0192-9275.

Anonymous:2018:BRP

[Ano18d] Anonymous. Book review: Paul Shaw, Revival Type: Digital Typefaces Inspired by the Past. Printing History (New Series), (23):??, Winter 2018. ISSN 0192-9275.

Anonymous:2018:IDC

[Ano18e] Anonymous. An interview with Douglas Charles, journeyman printer. Printing History (New Series), (24):??, Summer 2018. ISSN 0192-9275.

Anonymous:2018:IHL

[Ano18f] Anonymous. Interview with Heiba Lamara. Printing History (New Series), (23):??, Winter 2018. ISSN 0192-9275.

Anonymous:2019:IBK

[Ano19a] Anonymous. An interview with Bruce Kennet, author of W. A. Dwiggins: A Life in Design, published by the Letterform Archive in 2017. Printing History (New Series), (25):??, Winter 2019. ISSN 0192-9275.

Anonymous:2019:ITI

[Ano19b] Anonymous. An interview with Thin Ice Press. Printing History (New Series), (26):??, Summer 2019. ISSN 0192-9275.

Anonymous:20xx:TPB

[Anoxx] Anonymous. Type to Print:

The Book & The Type Specimen Book. Printing History, 22(1–2):??, ???? 20xx. ISSN 0192-9275. URL https://www.

rit.edu/press/printing-history-
43--44-0.

Antonetti:2008:BRS

[Ant08] Martin Antonetti. Book review: Scott-Martin Kosofsky, ed. The SP Century: Boston’s Society of Printers through One Hundred Years of Change. Printing History (New Series), (3):??, January 2008. ISSN 0192-9275.

Antonetti:2010:BRR

[Ant10] Martin Antonetti. Book review: Richard Benson, The Printed Picture. Printing History (New Series), (8):??, July 2010. ISSN 0192-9275.

Archer-Parre:2017:LJP

[AP17] Caroline Archer-Parré. Leonard Jay: A pioneer of printing education. Printing History (New Series), (21):??, Winter 2017. ISSN 0192-9275.

Argetsinger:1991:AGR

[Arg92]  Mark Argetsinger.  Adobe

Garamond: A review: Jerry Kelly, Adobe Garamond: A New Adaptation of a Sixteenth-Century Type. Printing History, 13/14(2/1):??, ???? 1991–1992.

Ashby:1980:BRS

[Ash80] Anna Lou Ashby. Book review: Sir Francis Meynell and Herbert Simon, Fleuron Anthology. Printing History, 2(1):??, ???? 1980. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-3-
0.

Ashbrook:2010:TPC

[Ash10] Susan Ashbrook. Two portraits of Cape Cod: Amelia Watson and Clare Leighton. Printing History (New Series), (8):??, July 2010. ISSN 0192-9275.

Auchincloss:2001:SRF

[Auc01] Kenneth Auchincloss. The second revival: Fine printing since World War II. Printing History, 21(1):3–11, ???? 2001. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-41-
0.

Auji:2014:BRF

[Auj14]  Hala Auji.  Book review:

Frances Wood and Mark Barnard, The Diamond Sutra: The Story of the World’s Earliest Dated Printed Book. Printing History (New Series), (16): ??, July 2014. ISSN 0192-9275.

Baker:1992:EHT

[Bak92] Morton H. Baker. The early history of the Thomas Todd Company, 1864–1924. Printing History, 14(2):??, ???? 1992.

Barlow:1976:NWP

[Bar76] Marjorie Dana Barlow. Notes on women printers in Colonial America and the United States, 1639–1975. Hroswitha Club, New York, NY, USA, 1976. xi + 89 pp. LCCN Z231 .B37 1976a.

Barker:1978:OUP

[Bar78] Nicolas Barker. The Oxford University Press and the spread of learning, 1478–1978: an illustrated history. Clarendon Press, New York, NY, USA, 1978. ISBN 0-19-951086-5, 0-19-951085-7 (paperback). xiii + 69 + 103 pp. LCCN Z232.O98 B37.

Barnes:1983:JWS

[Bar83] James J. Barnes. Jonas Winchester: Speculator, medicine man. Printing History, 5(1):??, ???? 1983. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-9-
0.

Barker:1985:AMD

[Bar85] Nicolas Barker. Aldus Manutius and the development of Greek script and type in the fifteenth century. Chiswick Book Shop, Sandy Hook, CT, USA, 1985. xiv + 115 pp. LCCN ???? With original leaves from the 1st Aldine ed. of Aristotle, 1497, Crastonus’ Dictionarium Graecum, 1497, Euripides, 1503, and the Septuagint, 1518.

Barker:1992:AMD

[Bar92] Nicolas Barker. Aldus Manutius and the development of Greek script and type in the fifteenth century. Fordham University Press, New York, NY, USA, 1992. ISBN 0-8232-1247-5. xiii + 138 pp. LCCN Z232.M3 B37 1992.

                                        URL http://catalogue.bnf.
                                        fr/ark:/12148/cb35605350p.

Barrett:1993:FCP

[Bar93] Timothy Barrett. Fifteenthcentury papermaking. Printing History, 15(2):??, ???? 1993.

Bartram:2007:TH

[Bar07] Alan Bartram. Typeforms: a history. British Library, London, UK, 2007. ISBN 0-7123-0971-3 (British Library), 1-58456-222-6 (Oak Knoll). 128 pp. LCCN Z250 .B29 2007.

Baskin:2018:MSU

[Bas18] Lisa Unger Baskin. The myth of some unconventional women. Printing History (New Series), (23):??, Winter 2018. ISSN 0192-9275.

Baxter:2000:GPE

[Bax00]  Alfred W. Baxter.  Grab-

horn Press ephemera: The missing matter. Printing History, 20(2):??, ???? 2000. ISSN 0192-9275.

                                        URL https://www.rit.edu/
                                        press/printing-history-40-
                                        0.

Blumenthal:1982:TYP

[BB82] Joseph Blumenthal and Joseph Blumenthal. Typographic years: a printer’s journey through a half century, 1925–1975. Grolier Club, New York, NY, USA, 1982. ISBN 0-913720-38-0 (Frederic C. Beil). 153 + 22 pp. LCCN Z232.B654 B58 1982b.

Brunsman:1988:WWP

[BB88] Charlotte K. Brunsman and August E. Brunsman. Wright & Wright, printers: The “other” career of Wilbur and Orville. Printing History, 10 (1):??, ???? 1988.

Buchanan-Brown:2005:EVI

[BB05] John Buchanan-Brown. Early Victorian illustrated books: Britain, France and Germany, 1820–1860. British Library, London, UK, 2005. ISBN 0-7123-4794-1 (hardcover), 1-58456-169-6 (Oak Knoll Press). 304 pp. LCCN NC977 .B827 2005.

Berry:2002:LCT

[BBB^{+}02] John D. Berry, Mark Batty, Robert Bringhurst, Maxim Zhukov, Matthew Carter, John Hudson, Gayanch Bagdasaryan, Vincent Figgins, and Roger Robertson, editors. Language culture type: international type design in the age of Unicode. ATypI, New York, NY, USA, 2002. ISBN 1-932026-01-0. ix + 373 pp. LCCN Z250.A2 L36 2002.

Berliner:1998:SBN

[BBRD99] Harold Berliner, Nicolas Barker, Jim Rimmer, and John Dreyfus. Starling Burgess, no type designer: A rebuttal of some allegations and suppositions made by Mike Parker in his article “Starling Burgess, Type Designer?” in Printing History 31/32 (1994). Printing History, 19(1):3–22, ???? 1998–1999. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-37-

0. See [Par94].

Barker:1983:SEN

[BC83] Nicolas Barker and John (John F. R.) Collins. A sequel to An enquiry into the nature of certain nineteenth century pamphlets by John Carter and Graham Pollard: the forgeries of H. Buxton Forman and T. J. Wise re-examined. Scolar Press, London, UK, 1983. ISBN 0-85967-638-2, 0-85967-639-0 (set). 394 pp. LCCN Z1024 .C32 1983.

Beckwith:2007:IGL

[Bec07] Alice H. R. H. Beckwith. Illustrating the good life: the Pissarros’ Eragny Press, 1894– 1914: a catalogue of an exhibition of books, prints and drawings, related to the work of the press: [catalogue of an exhibition: The Grolier Club, February 20–April 28, 2007 …]. The Grolier Club, New York, NY, USA, 2007. ISBN 0-910672-71-7 (paperback). ix + 69 pp. LCCN Z232.P65 B435 2007.

Belanger:1995:TYA

[Bel95] Terry Belanger. Twenty years after. Printing History, 17(1): ??, ???? 1995. URL https://

d3eb7xfyht02et.cloudfront.
net/wp-content/uploads/2011/
12/33-Belanger1.pdf.

Benton:1993:CVN

[Ben93] Megan L. Benton. C. Volmer Nordlunde: The “grand old man” of modern Danish printing. Printing History, 15(1):??, ???? 1993.

Benson:2008:PP

[Ben08] Richard Benson. The printed picture. Museum of Modern Art, New York, NY, USA, 2008. ISBN 0-87070-721-3. 338 pp. LCCN N5300 .B54 2008.

Bernhardt:1987:BRP

[Ber87] Melissa M. Bernhardt. Book review: Patricia Thomas Sbrebrnik, Alexander Strahan: Victorian Publisher. Printing History, 9(2):??, ???? 1987.

Berger:1990:IDA

[Ber90] Sidney E. Berger. Innovation and diversity among the Green family of printers. Printing History, 12(1):??, ???? 1990.

Berger:1993:BRW

[Ber93] Sidney E. Berger. Book review: William Blades, Numismata Typographica; or the Medallic History of Printing.

Printing History, 15(1):??, ???? 1993.

Berkson:2011:RRC

[Ber11] William Berkson. Readability and revival: The case of Caslon. Printing History (New Series), (10):??, July 2011. ISSN 0192-9275.

Berger:2016:BRJ

[Ber16] Sydney Berger. Book review: John Bidwell, American Paper Mills, 1690-1832: A Directory of the Paper Trade with Notes on Products, Watermarks, Distribution Methods, and Manufacturing Techniques. Printing History (New Series), (19):??, January 2016. ISSN 0192-9275.

Belanger:1986:HOB

[BF86] Terry Belanger and Gary Frost. How to operate a book. Book Arts Press, Charlottesville, VA, USA, 1986. 30-minute video.

Barnhill:2009:PVA

[BF09] Georgia Brady Barnhill and Nancy Finlay. Picturing Victorian America: Prints By the Kellogg Brothers of Hartford, Connecticut, 1830–1880. Connecticut Historical Society, Hartford, CT, USA, 2009. ISBN 1-881264-09-2 (hardcover), 1-881264-10-6 (paperback), 0-8195-7125-3 (e-book). LCCN NE2312.K35 A4 2009.

Bailyn:1980:PAR

[BH80] Bernard Bailyn and John B. Hench, editors. The Press and the American Revolution.

American Antiquarian Society, Worcester, MA, USA, 1980. ISBN 0-912296-18-6. 383 pp. LCCN PN4861 .P7. With a foreword by Marcus A. McCorison and an afterword by James Russell Wiggins.

Bianchi:1987:BRW

[Bia87]  Daniel B. Bianchi.  Book

review: William J. Glick, William Edwin Rudge. Printing History, 9(2):??, ???? 1987.

Bidwell:1979:BRJ

[Bid79]  John Bidwell.  Book re-

view: J. Ben Lieberman, Type and Typefaces. Printing History, 1(1):??, ???? 1979. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-1-
0.

Bidwell:1980:SCO

[Bid80] John Bidwell. Some Caslon ornaments in some American books. Printing History, 2(2): ??, ???? 1980. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-4-
0.

Bidwell:1982:BRM

[Bid82] John Bidwell. Book review:

Madeleine B. Stern, ed., Publishers for Mass Entertainment in Nineteenth Century America. Printing History, 4(1–2): ??, ???? 1982. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-7-
-8-0.

Bidwell:1984:BRF

[Bid84] John Bidwell. Book review: Fred Schreiber, The Estiennes: an annotated catalogue of 300 highlights of their various presses. Printing History, 6(1): ??, ???? 1984. ISSN 0192-9275.

                                        URL https://www.rit.edu/
                                        press/printing-history-11-
                                        0.

Bidwell:1991:BRM

[Bid92] John Bidwell. Book review: Martin Hutner, The Making of the Book of Common Prayer of 1928. Printing History, 13/14 (2/1):??, ???? 1991–1992.

Bidwell:2012:APM

[Bid12] John Bidwell. American Paper Mills, 1690–1832: A Directory of the Paper Trade, with Notes on Products, Watermarks, Distribution Methods, and Manufacturing Techniques. Dartmouth College Press, Hanover, NH, USA, 2012. ISBN 1-58465-964-5 (hardcover), 1-61168-316-5 (ebook). LCCN TS1088 .B46 2012.

Blumenthal:1987:MDR

[BJS87] Joseph Blumenthal, Herbert H. Johnson, and Carl Schlesinger. In memoriam: Dr. Robert Lincoln Leslie, humanitarian and educator, 1885–1987. Printing History, 9(2):??, ???? 1987.

Bakhareva:1983:RKB

[BL^{+}83] G. V. Bakhareva, Sergeĭ Pavlovich Luppov, et al. Russkie knigi i biblioteki v XVI-pervoi polovine XIX veka: sbornik nauchnykh trudov. (Russian) [Russian books and libraries in XVI–first half XIII century: a collection of scientific papers]. Biblioteka Akademii nauk SSSR, Leningrad, USSR, 1983. 159 pp. LCCN Z166.R87 R87 1983.

Blades:1883:NTM

[Bla83] William Blades. Numismata typographica: Medallic history of printing. “Printers register” office, London, UK, 1883. xviii + 144 pp. LCCN Z234 .B63.

Blake:1972:WBW

[Bla72] William Blake. William Blake’s water-colour designs for the poems of Thomas Gray. Eyre Methuen, London, UK, 1972. ISBN 0-413-29260-6. iii–xx, 72 + 30 pp. LCCN ND1942.B55 K49 1972. With an introduction and commentary by Sir Geoffrey Keynes.

Blades:1992:NTM

[Bla92] William Blades. Numismata typographica: the medallic history of printing. Bird and Bull Press, Newtown, PA, USA, 1992. xvii + 144 pp. LCCN Z234 .B63 1992. With a foreword by Henry Morris.

Blake:2000:BWC

[Bla00] William Blake. Blake’s watercolours for the poems of Thomas Gray: with complete texts. Dover Publications, Inc., New York, NY, USA, 2000. ISBN 0-486-40944-9 (paperback). ???? pp. LCCN ND1942.B55 A4 2000.

Blumenthal:1977:PBAa

[Blu77a]  Joseph Blumenthal.  The

printed book in America. D. R. Godine, Boston, MA, USA, 1977. ISBN 0-87923-210-2. xvi + 250 pp. LCCN Z208 .B56.

Blumenthal:1977:PBAb

[Blu77b]  Joseph Blumenthal.  The

printed book in America. Scolar Press, London, UK, 1977. ISBN 0-85967-412-6. xvi + 250 pp. LCCN Z208 .B56 1977b.

Blumenthal:1980:BRJ

[Blu80]  Joseph Blumenthal.  Book

review: John Ryder, The Case for Legibility. Printing History, 2(1):??, ???? 1980. ISSN 0192-9275.

                                        URL https://www.rit.edu/
                                        press/printing-history-3-
                                        0.

Blumenthal:1985:RFH

[Blu85] Joseph Blumenthal. Robert Frost and his printers. W. Thomas Taylor, Austin, TX, USA, 1985. ISBN 0-614-24482-X, 99976-1-597-2. xi + 105 pp. LCCN PS3511.R94 Z553 1985.

Bly:2015:WWP

[Bly15] Daniel W. Bly. Who was Paul Messerschmidt? Printing History (New Series), (17):??, January 2015. ISSN 0192-9275.

Bodger:1993:BRG

[Bod93] Lowell Bodger. Book review:

George Sadek and Maxim Zhukov, Typography: Polyglot. A Comparative Study in Multilingual Typesetting. Printing History, 15(1):??, ???? 1993.

Bolton:2016:FCP

[Bol16] Claire Bolton. The Fifteenth-Century printing practices of Johann Zainer, Ulm, 1473– 1478, volume third series, volume viii; no. 18 of Oxford Bibliographical Society publications; Publication Printing Historical Society. Oxford Bibliographical Society, Oxford, UK, 2016. ISBN 0-901420-59-X hardback. xv + 289 pp. LCCN Z232.Z18 B65 2016.

Bonn:1983:FPB

[Bon83]  Thomas L. Bonn.  The

first pocket book. Printing History, 5(2):??, ???? 1983. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-10-
0.

Boudreau:2008:PUE

[Bou08] Joan Boudreau. Publishing the U.S. Exploring Expedition: The fruits of the glorious enterprise. Printing History (New Series), (3):??, January 2008. ISSN 0192-9275.

Boudreau:2012:PPF

[Bou12] Joan Boudreau. The portable press and field printing during the American Civil War. Printing History (New Series), (12): ??, July 2012. ISSN 0192-9275.

Bidwell:1985:RFP

[BP85] John Bidwell and Cynthia Owen Philip. Robert Fulton’s portrait of Lord Stanhope. Printing History, 7(2):??, ???? 1985.

Berry:2003:TTH

[BR03] John D. Berry and John Randle. Type and typography: highlights from Matrix, the review for printers and bibliophiles. Batty, New York, NY, USA, 2003. ISBN 0-9715687-6-6 (hardcover), 0-9724240-8-3. xvii + 374 pp. LCCN Z250.A2 T95 2003.

Brewer:1990:MAW

[Bre90] Frederic Brewer. Mr. Artzt’s Western Lever: A hitherto unreported printing press. Printing History, 12(1):??, ???? 1990.

Bringhurst:2003:VM

[Bri03]  Robert Bringhurst.  The

voice in the mirror. Printing History, 23(2):??, ???? 2003. ISSN 0192-9275.

                                        URL https://www.rit.edu/
                                        press/printing-history-46-
                                        0.

Bright:2005:NLI

[Bri05] Betty Bright. No Longer Innocent: Book Art in America: 1960–1980. Granary Books, New York, NY, USA, 2005. ISBN 1-887123-71-7. xix + 301 + 8 pp. LCCN Z1033.F5 B75 2005.

Bright:2012:HF

[Bri12] Betty Bright. History as fable.

Printing History (New Series), (11):??, January 2012. ISSN 0192-9275.

Brody:1979:BRE

[Bro79] Catherine T. Brody. Book review: Elizabeth Carroll Reilly, A Dictionary of Colonial American Printers’ Ornaments and Illustrations. Printing History, 1(2):??, ???? 1979. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-2-
0.

Brockman:1988:BRD

[Bro88a] William S. Brockman. Book review: Donald C. Dickinson, Dictionary of American Book Collectors. Printing History, 10 (1):??, ???? 1988.

Brockman:1988:BRJ

[Bro88b] William S. Brockman. Book review: Jeanne Somers, Index to “The Dolphin” and “The Fleuron”. Printing History, 10(1): ??, ???? 1988.

Bruccoli:1986:FMK

[Bru86] Matthew J. (Matthew Joseph)

Bruccoli. The fortunes of Mitchell Kennerley, bookman. Harcourt, Brace, Jovanovich, San Diego, CA, USA, 1986. ISBN 0-15-132671-1. xiv + 337 pp. LCCN Z473.K45 B78 1986.

Bruckner:1990:FG

[Bru90] D. J. R. Bruckner. Frederic Goudy. Masters of American design. Documents of American Design, New York, NY, USA, 1990. ISBN 0-8109-1035-7. 144 pp. LCCN Z232.G68 B78 1990.

Bain:1998:BTN

[BS^{+}98] Peter Bain, Paul Shaw, et al. Blackletter: type and national identity. Princeton Architectural Press, New York, NY, USA, 1998. ISBN 1-56898-125-2. 72 pp. LCCN Z250.5.G6 B58 1998.

Bain:1999:BTN

[BS99] Peter Bain and Paul Shaw. Blackletter: Type and National Identity, A Catalog of the Exhibition, March 3 to May 2, 1998. Printing History, 19/20 (2/1):??, ???? 1999. ISSN 0192-9275.

Buechler:1982:AOB

[Bue82] John Buechler. “Adapted from an old book”: Some sources for Chiswick Press woodcut initials. Printing History, 4(1–2): ??, ???? 1982. ISSN 0192-9275.

                                        URL https://www.rit.edu/
                                        press/printing-history-7-
                                        -8-0.

Buechler:1983:CWP

[Bue83]  John Buechler.  Charles

Whittingham, printer, 1795– 1876. University of Vermont, Burlington, VT, USA, 1983. 39 pp. LCCN ????

Bumgardner:1982:VPA

[Bum82] Georgia B. Bumgardner. Vignettes of the past: American historical broadsides through the War of 1812. Printing History, 4(1–2):??, ???? 1982. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-7-
-8-0.

Burke:2007:ALJ

[Bur07] Christopher Burke. Active literature: Jan Tschichold and new typography. Hyphen, London, UK, 2007. ISBN 0-907259-32-4 (hardcover). 335 pp. LCCN Z232.T863 .B875 2007.

Bridson:1984:PPP

[BW84] Gavin Bridson and Geoffrey Wakeman. Printmaking and picture printing: a bibliographical guide to artistic and industrial techniques in Britain, 1750–1900. The Plough Press, Oxford, UK, 1984. ISBN 0-902813-12-9. LCCN Z117 .B84 1984.

Bruntjen:1979:DCM

[BY79] Scott Bruntjen and Melissa L.

Young. Douglas C. McMurtrie, bibliographer and historian of printing, volume 4 of The great bibliographers series. Scarecrow Press, Metuchen, NJ, USA, 1979. ISBN 0-8108-1188-X. xiv + 206 pp. LCCN Z1005 .M35.

Candee:2000:IIN

[Can00] Richard M. Candee. Illustrating invention: Nineteenthcentury machine advertising for the Aikens of Franklin, New Hampshire. Printing History, 20(2):??, ???? 2000. ISSN 0192-9275.

                                        URL https://www.rit.edu/
                                        press/printing-history-40-
                                        0.

Carter:1991:TLC

[Car92] Matthew Carter. Theories of letterform construction. Part 1. Printing History, 13/14(2/1): ??, ???? 1991–1992.

Carr:1994:BRT

[Car94] Dan Carr. Book review: Theo Rehak, Practical Typecasting. Printing History, 16(1–2):??, ???? 1994.

Carter:2001:RLP

[Car01] Sebastian Carter. The Rampant Lions Press — retrospectus and prospectus. Printing History, 21(1):??, ???? 2001. ISSN 0192-9275.

                                        URL https://www.rit.edu/
                                        press/printing-history-41-
                                        0.

Casper:2007:IB

[Cas07] Scott E. Casper, editor. The industrial book, 1840–1880, volume 3 of A history of the book in America. Published in association with the American Antiquarian Society by the University of North Carolina Press, Chapel Hill, NC, USA, 2007. ISBN 0-8078-3085-2 (hardcover). xix + 539 pp. LCCN Z473 .I53 2007.

Cave:1982:LYG

[Cav82] Roderick Cave. The lost years of the Golden Cockerel Press. Printing History, 4(1–2):??, ???? 1982. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-7-
-8-0.

Cave:1983:PP

[Cav83] Roderick Cave. The Private Press. R.R. Bowker, New York, NY, USA, second edition, 1983. ISBN 0-8352-1695-0. xvi + 389 pp. LCCN Z231.5.P7 C37 1983.

Cave:1984:BRW

[Cav84]  Roderick Cave.  Book re-

view: William Morris, The Ideal Book. Essays and Lectures on the Arts of the Book. Printing History, 6(1): ??, ???? 1984. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-11-
0.

Cave:1988:SSR

[Cav88] Roderick Cave. The Stockdale sisters revisited: Women printers and editors in the West Indies. [With a rebuttal by John A. Lent]. Printing History, 10 (2):??, ???? 1988. See [Len88].

Cazden:1984:SHG

[Caz84] Robert E. (Robert Edgar) Cazden. A social history of the German book trade in America to the Civil War, volume 1 of Studies in German literature, linguistics, and culture. Camden House, Columbia, SC, USA, 1984. ISBN 0-938100-09-2. xvii + 801 pp. LCCN Z479 .C39 1984.

Cameron:2018:IWM

[CB18] James T. Cameron and George D. Barnum. Invisible writing made visible: The U.S. Government Printing Office and prisoner of war stationery in the Second World War. Printing History (New Series), (23):??, Winter 2018. ISSN 0192-9275.

Cave:1985:GWT

[CC85] Roderick Cave and Kathleen Coleridge. For gospel and wool trade: Early printing in New Zealand. Printing History, 7 (1):??, ???? 1985.

Chen:2018:RDT

[CCS18] Amy Hildreth Chen, Michelle Chesner, and Shannon K. Supple. Roundtable discussion: Teaching printing history with Codex Conquest. Printing History (New Series), (23):??, Winter 2018. ISSN 0192-9275.

Crump:1983:SEC

[CH83] M. (Michael) Crump and Michael Harris, editors. Searching the eighteenth century: papers presented at the Symposium on the Eighteenth Century Short Title Catalogue in July 1982. British Library in association with the Department of Extra-Mural Studies, University of London, London, UK, 1983. ISBN 0-7123-0020-1. LCCN Z699.4.E17 S95 1982.

Chartier:1990:HEF

[Cha90] Roger Chartier. Histoire de l’édition française. Printing History, 12(1):??, ???? 1990.

Chrisman:1982:LCL

[Chr82] Miriam Usher Chrisman. Lay culture, learned culture: books and social change in Strasbourg, 1480–1599. Yale University Press, New Haven, CT, USA, 1982. ISBN 0-300-02530-0. xxx + 401 pp. LCCN DC801.S577 C47 1982.

Clouse:2008:MSJ

[Clo08]  Doug Clouse.  Mackellar,

Smiths and Jordan: typographic tastemakers of the late nineteenth century. Oak Knoll Press, New Castle, DE, USA, 2008. ISBN 1-58456-232-3 (hardcover). 174 + 2 pp. LCCN Z250.A2 C58 2008.

Coakley:2003:HHP

[Coa03] J. F. Coakley. Homan Hallock, punchcutter. Roger Stoddard, for Hugh from Roger. [Tribute for Hugh Amory]. Printing History, 23(1):??, ???? 2003. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-45-
0.

Comparato:1979:CGF

[Com79] Frank E. Comparato. Chronicles of genius and folly: R. Hoe and Company and the printing press as a service to democracy.

Labyrinthos, Culver City, CA, USA, 1979. xv + 846 pp. LCCN Z249 .C65 1979.

Corbin:1986:BRJ

[Cor86] Harold Corbin. Book review: Joseph Blumenthal, Robert Frost and His Printers. Printing History, 8(1):??, ???? 1986.

Cost:1994:LBB

[Cos94] Patricia A. Cost. Linn Boyd Benton, Morris Fuller Benton, and typemaking at ATF. Printing History, 16(1–2):??, ???? 1994. URL https://

                                        d3eb7xfyht02et.cloudfront.
                                        net/wp-content/uploads/2011/
                                        12/31-32-Cost-Benton1.pdf.

Carter:1934:ENC

[CP34] John Carter and Graham Pollard. An enquiry into the nature of certain nineteenth century pamphlets. Constable and Co., Ltd., London, UK, 1934. ISBN 0-8383-1261-6. xii + 400 pp. LCCN Z1024 .C32.

Carter:1971:ENC

[CP71] John Carter and Graham Pollard. An enquiry into the nature of certain nineteenth century pamphlets. Haskell House Publishers, New York, NY, USA, 1971. ISBN 0-8383-1261-6. viii + 400 pp. LCCN Z1024 .C32 1971.

Carter:1992:ENC

[CP92] John Carter and Graham Pollard, editors. An enquiry into the nature of certain nineteenth century pamphlets: with an epilogue. Scolar Press, Aldershot, Hants, England, second edition, 1992. ISBN 0-85967-639-0 (set), 0-938768-31-X. 400 + 41 pp. LCCN Z989.W8 C37 1992. With an epilogue by Nicolas Barker and John Collins.

Crawford:1985:CRA

[Cra85] Alan Crawford. C. R. Ashbee: architect, designer and romantic socialist. Yale University Press, New Haven, CT, USA, 1985. ISBN 0-300-03467-9. 499 pp. LCCN N6797.A76 C7 1985. URL http:/

/www.gbv.de/dms/ilmenau/
toc/013632698.PDF.

Crawford:2005:CRA

[Cra05] Alan Crawford. C. R. Ashbee:

architect, designer and romantic socialist. Yale University Press, New Haven, CT, USA, second edition, 2005. ISBN 0-300-10939-3 (paperback). 499 pp. LCCN N6797.A76 C7 2005.

Creesy:2006:MHT

[Cre06] Charles Creesy. Monticello:

The history of a typeface. Printing History, 25(1):??, ???? 2006. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-49-
0.

Cyr:1993:JMA

[Cyr93] Paul A. Cyr. Joel Munsell, Aldus’s disciple in Albany. Printing History, 15(2):??, ????

Dalal:2018:BRH

[Dal18]  Radha Dalal.  Book re-

view: Hala Auji, Printing Arab Modernity: Book Culture and the American Press in Nineteenth-Century Beirut. Printing History (New Series), (24):??, Summer 2018. ISSN 0192-9275.

Dane:1980:BR

[Dan80] William J. Dane. A Bradley reminiscence. Printing History, 2(1):??, ???? 1980. ISSN 0192-

                                        9275. URL https://www.rit.
                                        edu/press/printing-history-
                                        3-0.

Dane:1994:CMP

[Dan94] Joseph A. Dane. The curse of the mummy paper. Printing History, 17(2):??, ???? 1994. ISSN 0192-9275.

                                        URL https://www.rit.edu/
                                        press/printing-history-34-
                                        0.

Dane:2002:HAB

[Dan02] Joseph A. Dane. The Huntington apocalypse blockbook (Schreiber editions IV/V) with a note on terminology. Kay Amert, A Renaissance Font: Paris, 1516. Printing History, 21(2):??, Spring 2002. ISSN 0192-9275.

                                        URL https://www.rit.edu/
                                        press/printing-history-42-
                                        0.

Darnton:1982:LUO

[Dar82] Robert Darnton. The literary underground of the Old Regime. Harvard University Press, Cambridge, MA, USA, 1982. ISBN 0-674-53656-8. ix + 258 pp. LCCN DC133.3 .D37 1982.

Davis:1981:MMP

[Dav81] Natalie Zemon Davis. Misprint and Minerva: Printers’ journeymen in Sixteenth Century Lyon. Printing History, 3(1): ??, ???? 1981. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-5-
0.

Davis:1984:BRM

[Dav84] Stephen Paul Davis. Book review: M. Crump and M. Harris, eds. Searching the Eighteenth Century: Papers Presented at the Symposium on the Eighteenth Century Short Title Catalogue in July 1982. Printing History, 6(2):??, ???? 1984.

Davids:2011:PLB

[Dav11] Betsy Davids. From palm leaf to book: A South Asia quest. Printing History (New Series), (10):??, July 2011. ISSN 0192-9275.

Day:1966:BTE

[Day66] Kenneth Day, editor. Book typography, 1815–1965: in Europe and the United States of America. University of Chicago Press, Chicago, IL, USA, 1966. xxiii + 401 + 192 pp. LCCN Z116.A2 D3913 1966b.

deMarezOyens:1982:BRP

[de 82] Felix B. de Marez Oyens. Book review: Peter Amelung, Der Frühdruck im deutschen Südwesten 1473–1500: eine Ausstellung der Württembergischen Landesbibliothek Stuttgart. Band 1: Ulm. Printing History, 4(1– 2):??, ???? 1982. ISSN 0192-

                                        9275. URL https://www.rit.
                                        edu/press/printing-history-
                                        7--8-0.

DeCandido:1984:BRT

[DeC84] GraceAnne Andreassi DeCandido. Book review: Tiller: A Bimonthly Devoted to the Arts and Crafts Movement. Printing History, 6(2):??, ???? 1984.

DeCandido:1987:BRA

[DeC87] GraceAnne A. DeCandido. Book review: Alan Crawford, C. R. Ashbee: Architect, Designer Romantic Socialist. Printing History, 9(2):??, ???? 1987.

DeVinne:1981:HCH

[DeV81] Theodore Low DeVinne. How he collected his bill. Printing History, 3(2):??, ???? 1981. ISSN 0192-9275.

                                        URL https://www.rit.edu/
                                        press/printing-history-5-
                                        0.

Devine:1985:PIL

[Dev85] Warren D. Devine, Jr. The printing industry as a leader in electrification, 1883–1930. Printing History, 7(2):??, ???? 1985.

Dreyfus:1993:TMA

[DG93] John Dreyfus and Eric Gill. A typographical masterpiece: an account by John Dreyfus of Eric Gill’s collaboration with Robert Gibbings in producing the Golden Cockerel Press edition of “The four Gospels” in 1931. Bain and Williams, London, UK, 1993. ISBN 1-873359-00-4. xii + 105 pp. LCCN Z232.G63 D7 1993.

Dickinson:1986:DAB

[Dic86] Donald C. Dickinson. Dictionary of American book collectors. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 1986. ISBN 0-313-22544-3 (hardcover). xvi + 383 pp. LCCN Z989.A1 D53 1986.

Doggett:1979:LIP

[DM79] Marguerite V. Doggett and Karin N. Mango, editors. Long Island printing, 1791–1830: a checklist of imprints. Long Island Historical Society, Brooklyn, NY, USA, 1979. ISBN 0-89062-071-7. xxv + 191 pp. LCCN Z209.L57 D63.

Dodd:2006:GOI

[Dod06] Robin Dodd. From Gutenberg to OpenType: an illustrated history of type from the earliest letterforms to the latest digital fonts. Hartley and Marks, Vancouver, BC, Canada, 2006. ISBN 0-88179-210-1 (paperback). 192 pp. LCCN Z250 A2 D63 2006.

Doubleday:2006:JTD

[Dou06] Richard B. Doubleday. Jan Tschichold, designer: the Penguin years. Oak Knoll Press, New Castle, DE, USA, 2006. ISBN 1-58456-178-5 (Oak Knoll Press), 0-85331-946-4 (Lund Humphries). xxi + 218 pp. LCCN Z232.T863 D68 2006.

Dreyfus:1981:HNP

[Dre81] John Dreyfus. A history of the Nonesuch Press. Nonesuch Press, London, UK, 1981. ISBN 0-370-30397-0. xv + 319 pp. LCCN Z232.N82 D74 1981. With an introduction by Geoffrey Keynes and a descriptive catalogue by David McKitterick and Simon Rendall and John Dreyfus.

Dreyfus:1985:TIP

[Dre85] John Dreyfus. A transatlantic involvement with printing history. Printing History, 7(1):??, ???? 1985.

Dreyfus:1987:BRP

[Dre87] John Dreyfus. Book review: Paul Needham, The Printer and the Pardoner. Printing History, 9(2):??, ???? 1987.

Dreyfus:1993:BRJ

[Dre93] John Dreyfus. Book review: John Rathé, Bibliography of the Typophile Chap Books, 1935– 1992. Printing History, 15(1): ??, ???? 1993.

Dreyfus:1994:PSW

[Dre94] John (John Gustave) Dreyfus.

Into print: selected writings on printing history, typography and book production. British Library, London, UK, 1994. ISBN 0-7123-0343-X. x + 339 pp. LCCN Z124 .I55 1994.

Drogin:1983:AMS

[Dro83] Marc Drogin. Anathema!: medieval scribes and the history of book curses. Allanheld and Schram, Totowa, NJ, USA, 1983. ISBN 0-8390-0301-3. xx + 137 pp. LCCN Z6 .D76 1983.

DeHamel:2005:DDL

[DSC^{+}05] Christopher De Hamel, Joel Silver, John P. Chalmers, Daniel Wayne Mosser, and Michael Thompson. Disbound and dispersed: the leaf book considered. Caxton Club, Chicago, IL, USA, 2005. ISBN 1-58456-161-0. ???? pp. LCCN Z1033.L4 D4 2005.

Duensing:1984:BRW

[Due84] Paul Hayden Duensing. Book review: William Caslon, A Specimen of Printing Types. Introduction and notes by James Mosley. Printing History, 6(2): ??, ???? 1984.

Duensing:1987:BRR

[Due87] Paul Hayden Duensing. Book review: Richard E. Huss, The Printer’s Composition Matrix. Printing History, 9(2):??, ????

Duensing:1991:BRW

[Due91] Paul Hayden Duensing. Book review: Walter Wilkes, Das Schriftgeissen — Eine Dokumentation. Printing History, 14 (1):??, ???? 1991.

Dunlap:1979:TVV

[Dun79]  Joseph R. Dunlap.  Two

Victorian voices advocating good book design: I. Henry Stevens and the Shoddimites. Printing History, 1(1):??, ???? 1979. ISSN 0192-9275.

                                        URL https://www.rit.edu/
                                        press/printing-history-1-
                                        0.

Dunlap:1980:TVV

[Dun80]  Joseph R. Dunlap.  Two

Victorian voices advocating good book design: II. Charles Kegan Paul, perceptive publisher. Printing History, 2(1): ??, ???? 1980. ISSN 0192-9275.

                                        URL https://www.rit.edu/
                                        press/printing-history-3-
                                        0.

Dunlap:1988:BRP

[Dun88a] Joseph Dunlap. Book review: Peter Stansky, Redesigning the World: William Morris, the 1880s, and the Arts and Crafts. Printing History, 10(2):??, ???? 1988.

Dunlap:1988:EWO

[Dun88b] Joseph Dunlap. Emery Walker, Oscar Wilde and May Morris, Emery Walker and “letterpress printing”: A centennial celebration. Printing History, 10 (2):??, ???? 1988.

Dwiggins:1980:TWD

[Dwi80] W. A. (William Addison)

Dwiggins. A Tribute to W. A. Dwiggins on the hundredth anniversary of his birth: Tribute to William Addison Dwiggins. Priv. print. for friends of Hermann Püterschein at the Inkwell Press, New York, NY, USA, 1980. 153 pp. LCCN Z116.A3 T74 1983.

Dyer:1982:BJP

[Dye82] Alan Dyer. A biography of James Parker, colonial printer. Whitston Pub. Co., Troy, NY, USA, 1982. ISBN 0-87875-202-1. xvii + 421 pp. LCCN Z232.P24 D93 1982.

Dyson:1984:PPN

[Dys84] Anthony Dyson. Pictures to print: the nineteenth-century engraving trade. Farrand Press, London, UK, 1984. ISBN 1-05-083001-0, 1-05-083002-9 (limited ed.), 1-85083-001-0 (jacket). xxx + 234 pp. LCCN Z326 .D9 1984.

Easton:1983:MHB

[Eas83] Phoebe Jane Easton. Marbling:

a history and a bibliography. Dawson’s Book Shop, Los Angeles, CA, USA, 1983. ISBN 0-87093-180-6. xiii + 190 + 21 pp. LCCN Z271 .E2 1983.

Eckman:1985:GPT

[Eck85] James Eckman. A genial philippic on taste. Printing History, 7(2):??, ???? 1985.

Eckman:1987:UTF

[Eck87] James Eckman. The Union Type Foundry of Chicago, 1884–1892. Printing History, 9(1):??, ???? 1987.

Eisenstein:1981:EPR

[Eis81] Elizabeth E. Eisenstein. The early printer as “Renaissance Man”. Printing History, 3(1): ??, ???? 1981. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-5-
0.

Eisenstein:1994:PPA

[Eis94] Elizabeth L. Eisenstein. The printing press as an agent of change: communications and cultural transformation in early-modern Europe. Vol. I, II. Cambridge University Press, Cambridge, UK, reprint edition, 1994. ISBN 0-521-21967-1 (vol. I, hardcover), 0-521-21969-8 (vol. II, hardcover), 0-521-22044-0 (set of two hardcover volumes), 0-521-29955-1 (combined paperback edition). xxi + 794 (2 volumes) pp. LCCN Z124 .E37 1980.

Ebeling-Koning:1983:BRR

[EK83] Blanche T. Ebeling-Koning. Book review: Robert Darnton, The Literary Underground of the Old Regime. Printing History, 5(2):??, ????

1983. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-10-
0.

Eliason:2015:HHT

[Eli15] Craig Eliason. A history of the “Humanist” type classification. Printing History (New Series), (18):??, July 2015. ISSN 0192-9275.

Etherington:1982:BCB

[ER82] Don Etherington and Matt Roberts. Bookbinding and the conservation of books: a dictionary of descriptive terminology. Library of Congress, 1982. LCCN Z266.7.

Erickson:2017:AHO

[Eri17]  Jesse Ryan Erickson.  An

aesthetic history of the ouija board. Printing History (New Series), (22):??, Summer 2017. ISSN 0192-9275.

Evenden:2008:PPP

[Eve08] Elizabeth Evenden. Patents, pictures and patronage: John Day and the Tudor book trade, volume 2008: 2 of St Andrews studies in Reformation history. Ashgate, Aldershot, UK, 2008. ISBN 0-7546-5480-X. x + 220 pp. LCCN Z232.D277 E94 2008.

Fallon:2002:WIS

[Fal02] Peter K. Fallon. Why the Irish speak English: The consequences of one culture’s resistance to technological change. Printing History, 21(2):??, Spring 2002. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-42-
0.

Farmer:1983:BRJ

[Far83a]  David Farmer.  Book re-

view: John Dreyfus, A History of the Nonesuch Press. Printing History, 5(2):??, ???? 1983. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-10-
0.

Farrell:1983:BRN

[Far83b] Mark Farrell. Book review: Nati Krivatsy, Bibliography of the Works of Gregorio Leti. Printing History, 5(2):??, ???? 1983. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-10-
0.

Farrington:2019:VGB

[Far19] Lynne Farrington. A very good book indeed: Selling bibles by subscription in Nineteenth-Century America. Printing History (New Series), (25):??, Winter 2019. ISSN 0192-9275.

Finlay:2009:PVA

[FB^{+}09] Nancy Finlay, Georgia Brady Barnhill, et al., editors. Picturing Victorian America: prints by the Kellogg brothers of Hartford, Connecticut, 1830– 1880. Connecticut Historical Society, Hartford, CT, USA, 2009. ISBN 1-881264-09-2 (hardcover), 1-881264-10-6 (paperback). xv + 233 pp. LCCN NE2312.K35 A4 2009.

Feather:1986:DBH

[Fea86]  John Feather.  A Dictio-

nary of Book History. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1986. ISBN 0-19-520520-0. 278 pp. LCCN Z1006 .F38

1986. URL http://www.loc.
gov/catdir/enhancements/
fy0635/86012860-d.html.

Farren:2009:BBF

[FIbibBLL09] Donald Farren, August A.

Imholtz, Linda F. Children’s books in bygone Baltimore Lapides, and Linda F. Lapides, editors. The Baltimore Bibliophiles at fifty, 1954–2004: with, Children’s books in bygone Baltimore: an essay and a catalogue by Linda F. Lapides. Baltimore Bibliophiles, Baltimore, MD, USA, 2009. ISBN 1-58456-251-X. 175 + 1 pp. LCCN Z1008.B25 B25 2009.

Fletcher:1986:BRN

[Fle86] H. George Fletcher. Book review: Nicolas Barker, Aldus Manutius and the Development of Greek Script & Type in the Fifteenth Century. Printing History, 8(2):??, ???? 1986.

Fletcher:1993:IHS

[Fle93] H. George Fletcher. The ideal of the humanist scholar-printer: Aldus in Venice. Printing History, 15(2):??, ???? 1993.

Fletcher:2009:AUM

[Fle09] H. George Fletcher. Aldus, UCLA, and me. Printing History (New Series), (5):??, January 2009. ISSN 0192-9275.

Flint:2007:GIA

[Fli07] Richard Flint. ‘A great industrial art’: Circus posters, business risks, and the origins of color letterpress printing in America. Printing History, 25(2):??, ???? 2007. ISSN 0192-9275.

                                        URL https://www.rit.edu/
                                        press/printing-history-50-
                                        0.

Ford:1990:WWA

[For90] Margaret Lane Ford. A widow’s work: Ann Franklin of Newport, Rhode Island. Printing History, 13(2):??, ???? 1990.

Franklin:1986:AP

[Fra86a] Colin Franklin. The Ashendene Press. Bridwell Library, Southern Methodist University, Dallas, TX, USA, 1986. xiv + 256 + 24 pp. LCCN Z232.A76 F73 1986.

Fraser:1986:BRP

[Fra86b] James H. Fraser. Book review: The Printing History Series of Verlag Renate Raecke. Printing History, 8(2):??, ???? 1986.

Fraser:1986:UPE

[Fra86c] Sibylle Fraser. Underground printing in Europe, 1933– 1945 an historical observation.

Printing History, 8(1):??, ???? 1986.

Fraser:1987:PTK

[Fra87] James Fraser. Profile: Takashi Kono at eighty. Printing History, 9(1):??, ???? 1987.

Fraser:1988:BRC

[Fra88] James H. Fraser. Book review:

Colin Franklin, The Ashendene Press. Printing History, 10(1): ??, ???? 1988.

Fraser:1989:BRW

[Fra89]  James Fraser.  Book re-

view: William B. Todd and Ann Bowden, Tauchnitz International Editions in English 1841–1955; A Bibliographical History. Printing History, 11 (1):??, ???? 1989.

Garvey:1991:LAP

[Gar91] Eleanor M. Garvey. Leaves from an album of printing and graphic arts. Printing History, 14(1):??, ???? 1991.

Gehl:1998:ECW

[Geh98] Paul F. Gehl. The Europeans are coming! Or, what’s new in Continental history of the book: A review essay. Printing History, 18(1):??, March 1998. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-35-
0.

Gehl:2009:BRG

[Geh09]  Paul F. Gehl.  Book re-

view: Giancarlo Petrella, Uomini, torchi e libri nel Rinascimento. Printing History (New Series), (6):??, July 2009. ISSN 0192-9275.

Gehl:2015:BRK

[Geh15] Paul Gehl. Book review: Kay Amert, The Scythe and the Rabbit: Simon de Colines and the Culture of the Book in Renaissance Paris. Printing History (New Series), (18):??, July 2015. ISSN 0192-9275.

Genz:1984:BRB

[Gen84] Marcella Genz. BR to Robert Ballou. [Bruce Rogers letter]. Printing History, 6(1):??, ???? 1984. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-11-
0.

Gill:1982:PHG

[Gil82] Eric Gill. Printed by Hague and Gill: a checklist prepared in conjunction with the exhibit, A responsible workman, observing Eric Gill’s centenary, UCLA Library, 26 April–16 July, 1982. Regents of the University of California, Los Angeles, CA, USA, 1982. 48 pp. LCCN Z232.H143 P74 1982.

Goldman:2006:HPA

[GK06] Yosef Goldman and Ari (Ari G. M.) Kinsberg. Hebrew printing in America, 1735–1926: a history and annotated bibliography. YG Books, Brooklyn, NY, USA, 2006. ISBN 1-59975-685-4 (hardcover). xxvii + vii + 1188 pp. LCCN Z7070 .G595 2006.

Gross:2010:HBA

[GK10] Robert A. Gross and Mary Kelley, editors. A history of the book in America. Volume 2, An extensive republic: print, culture, and society in the new nation, 1790–1840. American Antiquarian Society by The University of North Carolina Press, Chapel Hill, NC, USA, 2010. ISBN 0-8078-3339-8. xix + 697 pp. LCCN Z2010 629186.

Gibson:1980:SSS

[GL80] John Gibson and Laurie Lewis.

Sticks and stones: some aspects of Canadian printing history. Toronto Typographic Association, Toronto, ON, Canada, 1980. 6 + 120 + 1 pp. LCCN Z206 .S74 1980.

Glaister:1979:GGBa

[Gla79a] Geoffrey Ashall Glaister. Glaister’s glossary of the book: terms used in papermaking, printing, bookbinding and publishing with notes on illuminated manuscripts and private presses. Allen and Unwin, London, UK, second edition, 1979. ISBN 0-04-010006-5. xv + 551 + 8 pp. LCCN Z118 .G457 1979.

Glaister:1979:GGBb

[Gla79b] Geoffrey Ashall Glaister. Glaister’s glossary of the book: terms used in papermaking, printing, bookbinding and publishing with notes on illuminated manuscripts and private presses. University of California Press, Berkeley, CA, USA, second edition, 1979. ISBN 0-520-03364-7. 551 pp. LCCN Z118 .G55 1979.

Gleboff:1986:BRG

[Gle86] Serge Gleboff. Book review: G. V Bakhareva and S. P. Luppov, Russkie knigi i biblioteki v XVI — pervoî polovine XIX veka: sbornik nauchnykh trudov. Printing History, 8(1):??, ???? 1986.

Glick:1984:WER

[Gli84] William J. Glick. William Edwin Rudge, volume 57 of Typophile chap book. Typophiles, New York, NY, USA, 1984. x + 91 + 32 pp. LCCN Z232.R89 G55 1984.

Goble:1990:BRC

[Gob90] Corban Goble. Book review: Carl Schlesinger, ed., The Biography of Ottmar Mergenthaler; A New Edition. Printing History, 13(2):??, ???? 1990.

Goble:1998:MTN

[Gob98a] Corban Goble. Mark Twain’s nemesis: The Paige Compositor. Printing History, 18(2):2– 16, ???? 1998. ISSN 0192-9275.

                                        URL http://www.codex99.
                                        com/typography/124.html;
                                         https://circuitousroot.
                                        com/artifice/letters/press/
                                        noncastcomp/paige/index.
                                        html;   https://www.rit.
                                        edu/press/printing-history-
                                        36-0.

Goble:1998:RTV

[Gob98b] Corban Goble. Rogers’s Typograph versus Mergenthaler’s Linotype: The push and shove 0f patents and priority in the 1890s. Printing History, 18(1):??, March 1998. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-35-
0.

Godburn:2016:NCD

[God16] Mark R. Godburn. Nineteenthcentury dust-jackets. Private Libraries Association, Ravelston, London, UK, 2016. ISBN 1-58456-347-8 (Oak Knoll Press), 0-900002-88-3 (Private Libraries Association). 215 pp. LCCN NC973 .G59 2016.

Goree:2014:WVP

[Gor14] John Goree. The woodblocks of Vesalius and the printings: From the Renaissance to the Modern Era. Printing History (New Series), (16):??, July 2014. ISSN 0192-9275.

Gray:1976:NCO

[Gra76]  Nicolete Gray.  Nineteenth

century ornamented typefaces. University of California Press, Berkeley, CA, USA, 1976. ISBN 0-520-03074-5. 238 pp. LCCN Z250.A2 G7 1976. With a chapter on Ornamented types in America by Ray Nash.

Grannis:1988:BRM

[Gra88] Chandler B. Grannis. Book review: Matthew J. Bruccoli, The Fortunes of Mitchell Kennerley, Bookman. Printing History, 10(1):??, ???? 1988.

Grannis:1990:BRB

[Gra90] Chandler Grannis. Book review: Bradbury Thompson, The Art of Graphic Design. Printing History, 12(1):??, ???? 1990.

Graalfs:2007:PRR

[Gra07] Gregory Graalfs. Photography in reproduction: Its role in the settlement of California. Printing History (New Series), (1): ??, January 2007. ISSN 0192-9275.

Green:1981:BRM

[Gre81]  James Green.  Book re-

view: Marguerite V. Doggett, Long Island Printing 1791– 1830. Printing History, 3(2): ??, ???? 1981. ISSN 0192-9275.

                                        URL https://www.rit.edu/
                                        press/printing-history-5-
                                        0.

Green:1986:BRR

[Gre86] James Green. Book review: Robert E. Cazden, A Social History of the German Book Trade in America to the Civil War; Christopher L. Dolmetsch, The German Press of the Shenandoah Valley. Printing History, 8(2):??, ???? 1986.

Green:1987:BRJ

[Gre87] James N. Green. Book review: John Tebbel, Between Covers: The Rise and Transformation of Book Publishing in America. Printing History, 9(2):??, ???? 1987.

Green:1994:CKB

[Gre94]  James N. Green.  ‘The

Cowl knows best what will suit in Virginia’: Parson Weems on Southern readers. Printing History, 17(2):??, ???? 1994. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-34-
0.

Green:2009:DAS

[Gre09] Nile Green. The development of Arabic-script typography in Georgian Britain. Printing History (New Series), (5):??, January 2009. ISSN 0192-9275.

Grossman:2001:TPW

[Gro01] Carol Grossman. The Trianon Press’s William Blake’s Water-colour Designs for the Poems of Thomas Gray. Printing History, 21(1):??, ???? 2001. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-41-
0.

Green:2006:BFW

[GS06] James N. Green and Peter Stallybrass. Benjamin Franklin: writer and printer. Oak Knoll, New Castle, DE, USA, 2006. ISBN 1-58456-187-4 (Oak Knoll Press), 0-7123-4938-3 (British Library). x + 179 pp. LCCN Z232.F8 G74

Hall:1986:HUP

[Hal86] Max Hall. Harvard University Press: a history. Harvard University Press, Cambridge, MA, USA, 1986. ISBN 0-674-38080-0, 0-674-38081-9 (paperback). 257 pp. LCCN Z473.H34.

Hanson:1993:BFW

[Han93] David A. Hanson. Baron Frederick Wilhelm von Egloffstein: Inventor of the first commercial halftone process in America. Printing History, 15(1):??, ???? 1993.

Hanff:2000:WWP

[Han00]  Peter E. Hanff.  Way &

Williams, Publishers, Chicago, 1895–98. Printing History, 20 (2):??, ???? 2000. ISSN 0192-

                                        9275. URL https://www.rit.
                                        edu/press/printing-history-
                                        40-0.

Hanson:2005:JCW

[Han05]  David A. Hanson.  John

Carbutt and the Woodburytype in America. Printing History, 24(2):??, ???? 2005. ISSN 0192-9275.

                                        URL https://www.rit.edu/
                                        press/printing-history-48-
                                        0.

Harlan:1980:BRJ

[Har80] Robert D. Harlan. Book review: John Tebbel, A History of Book Publishing in the United States, Volume III. Printing History, 2(1):??, ???? 1980. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-3-
0.

Harris:1981:BRH

[Har81] Elizabeth Harris. Book review:

The History of Printing from its Beginnings to 1930: The Subject Catalog of the American Type Founders Library in the Columbia University Libraries. Printing History, 3(2): ??, ???? 1981. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-5-

0. Introduction by Kenneth A. Lohf.

Harlan:1982:MWM

[Har82] Robert D. Harlan. The memoirs of William M. Cubery. Printing History, 4(1–2):??, ???? 1982. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-7-
-8-0.

Harris:1985:PMS

[Har85] Elizabeth M. Harris. Printing and a museum [the Smithsonian Institution Graphic Arts Division]. Printing History, 7 (2):??, ???? 1985.

Harris:1986:IPB

[Har86] Elizabeth M. Harris. Inventing printing for the blind. Printing History, 8(2):??, ???? 1986.

Harmsen:1987:RJH

[Har87] Tyrus G. Harmsen. Robinson Jeffers and his printers. Printing History, 9(2):??, ???? 1987.

Harvey:1988:BRW

[Har88] Clifford A. Harvey. Before Rosebud was a sled: Documentation and reprinting of early 19th Century commercial wood engravings from the GramLee Collection. Printing History, 10 (2):??, ???? 1988.

Harris:1989:PBP

[Har89]  Elizabeth Harris.  Press-

builders in Philadelphia, 1776– 1850. Printing History, 11(2): ??, ???? 1989.

Harris:1993:RP

[Har93] Elizabeth Harris. The rail presses. Printing History, 15 (2):??, ???? 1993.

Harlan:1998:OSF

[Har98]  Robert D. Harlan.  Ori-

gins of San Francisco fine printing traditions. Printing History, 18(2):??, ???? 1998. ISSN 0192-9275.

                                        URL https://www.rit.edu/
                                        press/printing-history-36-
                                        0.

Harlan:2003:SLC

[Har03]  Robert D. Harlan.  Seven

letters from C. H. St. John Hornby to John Henry Nash. Printing History, 24(1):??, ???? 2003. ISSN 0192-9275.

                                        URL https://www.rit.edu/
                                        press/printing-history-47-
                                        0.

Hauser:2006:PLS

[Hau06]  Don Hauser.  Printers of

the Lanes and Streets of Melbourne (1837–1975). Nondescript Press, Melbourne, VIC, Australia, 2006. ISBN 0-9758424-0-4 (hardcover). xi + 171 pp. LCCN Z221 .H38 2006.

Hawley:2017:WBP

[Haw17] E. Haven Hawley. William Berry: Publisher, scoundrel, and spiritualist. Printing History (New Series), (22):??, Summer 2017. ISSN 0192-9275.

Hanson:2013:CPC

[HB13] David Hanson and Edaward Bierstadt. Color photography and color printing. Printing History (New Series), (13):??, January 2013. ISSN 0192-9275.

Harvard:1981:ICC

[HC81] Stephen Harvard and Bernardino Cataneo. An italic copybook: the Cataneo manuscript, volume 5 of Studies in the history of calligraphy. Published for the Dept. of Printing and Graphic Arts of the Houghton Library, Harvard University and the Newberry Library by Taplinger Pub. Co., New York, NY, USA, 1981. ISBN 0-8008-4286-3. 55 pp. LCCN Z43.A3 H33.

Heller:2000:MIM

[Hel00] Marvin J. Heller. Mirrorimage monograms as printers’ devices on title pages of Hebrew books printed in the Seventeenth and Eighteenth Centuries. Printing History, 20(2): ??, ???? 2000. ISSN 0192-9275.

URL https://www.rit.edu/

                                        press/printing-history-40-
                                        0.

Heller:2011:BYB

[Hel11] Marvin J. Heller. Behold, you are beautiful, my love: The use of ornamental frames in Hebrew incunabula. Printing History (New Series), (10):??, July 2011. ISSN 0192-9275.

Heller:2015:EMS

[Hel15] Marvin J. Heller. The eagle motif in Sixteenth- and Seventeeth-Century Hebrew books. Printing History (New Series), (17):??, January 2015. ISSN 0192-9275.

Heller:2017:LME

[Hel17] Marvin J. Heller. The lion motif on early Hebrew title pages and pressmarks reviews. Printing History (New Series), (22):??, Summer 2017. ISSN 0192-9275.

Henry:1995:BRM

[Hen95]  Barbara Henry.  Book re-

view: Maurice Annenberg, Type Foundries of America and Their Catalogs, (with additions and an introduction by Stephen O. Saxe). Printing History, 17(1):??, ???? 1995. ISSN 0192-9275.

                                        URL https://www.rit.edu/
                                        press/printing-history-33-
                                        0.

Heritage:2009:CLJ

[Her09] Barbara Heritage. Collecting litho jam jar labels and teaching wood-engraved elephants:

Rare Book School’s Printing Surfaces Collection. Printing History (New Series), (6):??, July 2009. ISSN 0192-9275.

Heyeck:1985:BRP

[Hey85] Robin Heyeck. Book review:

Phoebe Jane Easton, Marbling. Printing History, 7(1):??, ???? 1985.

Hugill-Fontanel:2006:AMG

[HF06] Amelia Hugill-Fontanel. Arts et Métiers Graphiques: The graphic design magazine of the deberny et peignot type foundry. Printing History, 25 (1):??, ???? 2006. ISSN 0192-

9275. URL https://www.rit.
edu/press/printing-history-
49-0.

Hugill-Fontanel:2018:MTI

[HF18] Amelia J. Hugill-Fontanel.

Multitudinous tints: An inventor’s pursuit of instantaneous multicolor printing. Printing History (New Series), (24):??, Summer 2018. ISSN 0192-9275.

Harris:2006:MPC

[HH06] Moira F. Harris and Leo J.

Harris. Minnesota on paper: collecting our printed history. University of Minnesota Press, Minneapolis, MN, USA, 2006. ISBN 0-8166-4554-X, 0-8166-4555-8 (paperback). xxi + 221 pp. LCCN Z1029.5 .H37 2006.

Howard-Hill:2008:BBT

[HH08] T. H. (Trevor Howard) Howard-Hill. The British book trade, 1475–1890: a bibliography.

British Library, London, UK, 2008. ISBN 0-7123-5059-4 (set: British Library), 1-58456-255-2 (set: Oak Knoll Press). 2 volumes pp. LCCN Z324 .H69 2009.

Hoyle:2013:GGP

[HH13] John T. (John Thomas) Hoyle and Richard L. (Richard Lee) Hopkins. Going with Goudy to Philadelphia: a story of what man has wrought, volume 21 of Typographic curiosities. Hill and Dale Private Press and Typefoundry, Terra Alta, WV, USA, 2013. 19 pp. LCCN Z232.G68 H687 2013.

Hidy:2007:CLD

[Hid07] Lance Hidy. Calligraphy and letterpress in design education. Printing History (New Series), (2):??, July 2007. ISSN 0192-9275.

Hill:1980:BRD

[Hil80] Jonathan A. Hill. Book review: David McKitterick, ed., Stanley Morison & D. B. Updike: Selected Correspondence. Printing History, 2(2): ??, ???? 1980. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-4-
0.

Rosenwald:1982:EIB

[Hin82] Sandra Hindman, editor. The Early illustrated book: essays in honor of Lessing J. Rosenwald. Library of Congress, Washington, DC, USA, 1982. ISBN 0-8444-0398-9. xv + 260 + 8 pp. LCCN Z1023 .E18 1982.

Holliday:2007:EJM

[HJ07] Peter Holliday and Edward Johnston. Edward Johnston: master calligrapher. British Library, London, UK, 2007. ISBN 0-7123-4927-8 (British Library), 1-58456-198-X (Oak Knoll). xx + 389 pp. LCCN Z43.42.J64.

Holliday:2003:PAH

[HJB03] Peter Holliday, Edward Johnston, and Robert Bridges. 1901—1926: A phonetic alphabet in the half uncial script. Printing History, 24(1): ??, ???? 2003. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-47-
0.

Hodnett:1978:FBF

[Hod78] Edward Hodnett. Francis Barlow: first master of English book illustration. Scolar, London, UK, 1978. ISBN 0-85967-350-2. 237 pp. LCCN NC978.5.B37 H63 1978.

Hopkins:2009:SPH

[Hop09] Richard L. Hopkins. Saving printing history outside the box. Printing History (New Series), (6):??, July 2009. ISSN 0192-9275.

Hopkinson:2011:ELA

[Hop11] Martin Hopkinson. Ex Libris: the art of bookplates. The British Museum Press, London, UK, 2011. ISBN 0-7141-2674-8. 111 pp. LCCN Z993 .H67 2011.

Hopkins:2012:TLM

[Hop12] Richard L. (Richard Lee) Hopkins. Tolbert Lanston and the Monotype: the origin of digital typesetting. University of Tampa Press, Tampa, FL, USA, 2012. ISBN 1-59732-100-1 (hardcover), 1-59732-101-X (paperback). xvii + 214 pp. LCCN Z126.5 .H67 2012.

Howells:1980:CPR

[How80] William Dean Howells. The country printer. [reprint from Scribner’s Magazine]. Printing History, 2(2):??, ???? 1980. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-4-
0.

Harris:1978:CPB

[HS78] Elizabeth M. Harris and Clinton Sisson. The common press: being a record, description and delineation of the early eighteenth-century handpress in the Smithsonian Institution, with a history and documentation of the press. D. R. Godine, Boston, MA, USA, 1978. ISBN 0-87923-211-0. 62 pp. LCCN Z249 .H28.

Herdrich:1985:PPT

[HS85] Peter Herdrich and Terry Stevick. From punch to printing type: the art and craft of hand punching. Columbia University, New York, NY, USA, 1985. LCCN VBF 0752. One 60-minute videocassette.

Harmsen:2005:PPS

[HTBR05] Tyrus G. Harmsen, Stephen Tabor, Mariana Blau, and Patrick Reagh. The Plantin Press of Saul and Lillian Marks: a bibliography, 1930–1985. Press of Patrick Reagh, Sebastopol, CA, USA, 2005. xviii + 151 + 16 pp. LCCN Z232.P73 H35 2005.

Hudak:1978:EAW

[Hud78] Leona M. Hudak. Early American women printers and publishers, 1639–1820. Scarecrow Press, Metuchen, NJ, USA, 1978. ISBN 0-8108-1119-7. xxi + 813 pp. LCCN Z231 .H83.

Hudson:2008:DPE

[Hud08] Graham Spence Hudson. The design and printing of Ephemera in Britain and America, 1720– 1920. British Library, London, UK, 2008. ISBN 0-7123-4904-9 (British Library), 1-58456-224-2 (Oak Knoll press, rel). 159 pp. LCCN ????

Hulse:1982:DTP

[Hul82] Elizabeth Hulse. A dictionary of Toronto printers, publishers, booksellers and the allied trades: 1798–1900. Anson-Cartwright, Toronto, ON, Canada, 1982. ISBN 0-919974-08-2. xxii + 311 pp. LCCN Z488.6.T67 H84 1982.

Hunnisett:1980:DBS

[Hun80a] Basil Hunnisett. A dictionary of British steel engravers. F. Lewis, Leigh-on-Sea, England, 1980. ISBN 0-9531706-7-3. 148 pp. LCCN NE628.3 .H84.

Hunnisett:1980:SEB

[Hun80b] Basil Hunnisett. Steel-engraved book illustration in England. D. R. Godine, Boston, MA, USA, 1980. ISBN 0-87923-322-2. xvi + 263 pp. LCCN NE628.3 .H86 1980.

Huss:1985:PCM

[Hus85] Richard E. Huss. The printer’s composition matrix: a history of its origin and development. Oak Knoll series on the history of the book. Oak Knoll Books, New Castle, DE, USA, 1985. ISBN 0-938768-09-3. x + 56 + 8 pp. LCCN Z253 .H88 1985.

Hutner:1989:BRF

[Hut89]  Martin Hutner.  Book re-

view: Fünf Jahrhunderte Buchillustration, Meisterwerke der Buchgrapik, aus der Bibliothek Otto Schäfer. Printing History, 11(2):??, ???? 1989.

Hutner:1990:MBC

[Hut90] Martin Hutner. The making of the Book of Common Prayer of 1928. Chiswick Book Shop, Southbury, CT, USA, 1990. xvi + 75 + 24 pp. LCCN BX5945 .H88 1990.

Hutner:1993:DBU

[Hut93] Martin W. Hutner. Daniel Berkeley Updike: Humanist, scholar, printer. Printing History, 15(2):??, ???? 1993.

Isaac:1986:BRG

[Isa86] Elliot S. Isaac. Book review: Gary Marker, Publishing, Printing and the Origins of Intellectual Life in Russia, 1700– 1800. Printing History, 8(1):??, ???? 1986.

Janssen:2004:TDH

[Jan04] Frans (Franciscus Antonius Wilhelmus Gerard) Anton Janssen. Technique and design in the history of printing: 26 essays. Hes and de Graaf, ’t-Goy-Houten, The Netherlands, 2004. ISBN 90-6194-289-6. 380 pp. LCCN ????

Kecskes:1986:CII

[jK86] Lily Chia jen Kecskes. Chinese ink and inkmaking. Printing History, 8(1):??, ???? 1986.

Johnson:1979:AAN

[Joh79] Diane Johnson. American art nouveau. H. N. Abrams, New York, NY, USA, 1979. ISBN 0-8109-0678-3. 311 pp. LCCN N6510.5.A77 J63.

Johnson:1991:BRC

[Joh91] Herbert H. Johnson. Book review: A Catalogue of an Exhibition of the Philip Hofer Bequest. Printing History, 14(1): ??, ???? 1991.

Johnston:1991:GMC

[Joh92] Alastair Johnston. ‘Guard the mysteries! Constantly reveal them!’ The history of printing as shown in type specimens.

Printing History, 13/14(2/1): ??, ???? 1991–1992.

Johnston:1993:RCM

[Joh93] Alastair Johnston. Reflections on the centenary of the Merrymount Press. Printing History, 15(1):??, ???? 1993.

Johnson:2007:BRT

[Joh07a] Alastair Johnson. Book review: Tyrus Harmsen and Stephen Tabor, The Plantin Press of Saul and Lillian Marks: A Bibliography. Printing History (New Series), (2):??, July 2007. ISSN 0192-9275.

Johnston:2007:BRF

[Joh07b] Alastair Johnston. Book review: Fred Smeijers, Type Now: A Manifesto, plus Work So Far. Printing History (New Series), (1):??, January 2007. ISSN 0192-9275.

Johnston:2007:BRJb

[Joh07c] Alastair Johnston. Book review: Jane Pomeroy, Alexander Anderson: Wood-engraver and Illustrator. Printing History (New Series), (1):??, January 2007. ISSN 0192-9275.

Johnston:2007:BRJa

[Joh07d] Alastair Johnston. Book review: John A. Lane, Early Type Specimens in the Plantin– Moretus Museum. Printing History (New Series), (1):??, January 2007. ISSN 0192-9275.

Johnson:2008:BRF

[Joh08] Alastair Johnson. Book review: Frans A Janssen. Technique and Design in the History of Printing. Printing History (New Series), (3):??, January 2008. ISSN 0192-9275.

Johnston:2009:BRA

[Joh09a] Alastair Johnston. Book review: Alan Bartram, Typeforms: A History. Printing History (New Series), (5):??, January 2009. ISSN 0192-9275.

Johnston:2009:BRD

[Joh09b] Alastair Johnston. Book review: Doug Clouse, MacKellar, Smiths & Jordan: Typographic Tastemakers of the Late Nineteenth Century. Printing History (New Series), (6):??, July 2009. ISSN 0192-9275.

Johnston:2009:BRG

[Joh09c]  Alastair Johnston.  Book

review: Graham Hudson, The Design and Printing of Ephemera in Britain and America 1720–1890. Printing History (New Series), (5):??, January 2009. ISSN 0192-9275.

Johnston:2009:BRH

[Joh09d] Alastair Johnston. Book review: Hendrik D. L. Vervliet, The Paleo-Typography of the French Renaissance: Selected Papers on Sixteenth-Century Typefaces. Printing History (New Series), (6):??, July 2009. ISSN 0192-9275.

Johnston:2010:BRJ

[Joh10] Alastair Johnston. Book review: Jay T. Last, The Color Explosion: Nineteenth Century American Lithography. Printing History (New Series), (7): ??, January 2010. ISSN 0192-9275.

Johnston:2011:BRH

[Joh11a] Alastair Johnston. Book review: Hendrik D. L. Vervliet, French Renaissance Printing Types: A Conspectus. Printing History (New Series), (10):??, July 2011. ISSN 0192-9275.

Johnston:2011:BRM

[Joh11b] Alastair Johnston. Book review: Matthew McLennan Young, Field & Tuer, the Leadenhall Press. Printing History (New Series), (10):??, July 2011. ISSN 0192-9275.

Johnston:2011:GFC

[Joh11c] Alastair M. Johnston. A glance at the first century of California printing. Printing History (New Series), (9):??, January 2011. ISSN 0192-9275.

Joyce:1983:GBC

[Joy83] Donald F. Joyce. Gatekeepers of black culture: blackowned book publishing in the United States, 1817–1981, volume 70 of Contributions in Afro-American and African studies. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 1983.

ISBN 0-313-23332-2 (hardcover). ISSN 0069-9624. xiv + 249 pp. LCCN Z471 .J69 1983. URL http://

lcweb.loc.gov/catdir/toc/
becites/business/afro/82009227.
idx.html;   http://lcweb.
loc.gov/catdir/toc/becites/
business/afro/82009227.toc.
html;   http://lcweb.loc.
gov/rr/business/guide4.html.

Joyce:1985:BRJ

[Joy85] William L. Joyce. Book review:

James P. Bell, ed., John Peter Zenger and Freedom of the Press. Printing History, 7(2): ??, ???? 1985.

Kaser:1982:BRB

[Kas82] John Kaser. Book review:

Bernard Bailyn and John B. Hench, eds., The Press & the American Revolution. Printing History, 4(1–2):??, ???? 1982. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-7-
-8-0.

Kelly:2016:NRH

[KBKHF16] Jerry Kelly, Misha Beletsky, Jerry Kelly, and Amelia Hugill-Fontanel. The noblest roman: a history of the Centaur types of Bruce Rogers, volume 236 of Book Club of California publication. The Book Club of California, San Francisco, CA, USA, 2016. ISBN 0-9819597-9-

2. 120 + 16 + 2 pp. LCCN Z232.R67 K45 2016.

Kegler:2016:APE

[Keg16] Richard Kegler. The Aries Press of Eden, New York. RIT Press, Rochester, NY, USA, 2016. ISBN 1-939125-21-9 (hardcover), 1-939125-22-7 (deluxe edition), 1-939125-25-1 (e-book). viii + 85 pp. LCCN Z232.A69 K44 2016.

Kelly:1991:BRA

[Kel91] Jerry Kelly. Book review: Alexander S. Lawson, Anatomy of a Typeface; Printing Types: An Introduction; The Compositor as Artist, Craftsman, and Tradesman. Printing History, 14(1):??, ???? 1991.

Kelly:1998:BRJ

[Kel98] Jerry Kelly. Book review: John Dreyfus, Into Print: Selected Writings on Printing History, Typography and Book Production. Printing History, 18(1):??, March 1998. ISSN 0192-9275.

                                        URL https://www.rit.edu/
                                        press/printing-history-35-
                                        0.

Kelly:2003:BRM

[Kel03] Jerry Kelly. Book review: Marianne Tidcombe, The Doves Press. Printing History, 24(1): ??, ???? 2003. ISSN 0192-9275.

                                        URL https://www.rit.edu/
                                        press/printing-history-47-
                                        0.

Keller:2008:PBB

[Kel08a] Kate Van Winkle Keller. Printers of ballads, books, and newspapers: biographical notes and checklists for Nathaniel Coverly, Sr., Nathaniel Coverly, Jr., and Joseph White. American Antiquarian Society, Worcester, MA, USA, 2008. ISBN 1-929545-54-1. 162 pp. LCCN Z232.T4.

Kelly:2008:BRP

[Kel08b] Jerry Kelly. Book review: Peter Holliday, Edward Johnston: Master Calligrapher. Printing History (New Series), (3):??, January 2008. ISSN 0192-9275.

Kelly:2008:FFB

[Kel08c] Jerry Kelly. The first flowering: Bruce Rogers at the Riverside Press, 1896-1912; with a checklist of the Riverside Press editions. David R. Godine, Boston, MA, USA, 2008. ISBN 1-56792-374-7. 54 + 20 pp. LCCN Z232.R67.

Kelly:2009:BRC

[Kel09a] Jerry Kelly. Book review:

Christopher Burke, Active Literature: Jan Tschichold and New Typography. Printing History (New Series), (5):??, January 2009. ISSN 0192-9275.

Kelly:2009:HZN

[Kel09b] Jerry Kelly. Hermann Zapf at ninety. Printing History (New Series), (6):??, July 2009. ISSN 0192-9275.

Kelly:2010:ABT

[Kel10] Jerry Kelly. The art of the book in the twentieth century: a study of eleven influential book designers from 1900 to 2000. RIT Cary Graphic Arts Press, Rochester, NY, USA, 2010. ISBN 1-933360-46-1. xviii + 179 pp. LCCN Z246 .K45 2010.

Koda:1984:BRM

[Kod84]  Paul S. Koda.  Book re-

view: Matt T. Roberts and Don Etherington, Bookbinding and the Conservation of Books: A Dictionary of Descriptive Terminology. Printing History, 6(1):??, ???? 1984. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-11-
0.

Koda:2009:BRC

[Kod09a] Paul S. Koda. Book review: Carl F. Kaestle and Janice A. Radway, eds., A History of the Book in America, Volume 4: Print in Motion: The Expansion of Publishing and Reading in the United States, 1880– 1940. Printing History (New Series), (6):??, July 2009. ISSN 0192-9275.

Koda:2009:BRS

[Kod09b]  Paul S. Koda.  Book re-

view: Scott E. Casper, Jeffrey D. Groves, Stephen W. Nissenbaum, and Michael Winship, eds., A History of the Book in America (vol. 3). Printing History (New Series), (5):??, January 2009. ISSN 0192-9275.

Koda:2010:BRK

[Kod10a]  Paul S. Koda.  Book re-

view: Kate van Winkle Keller, Printers of Ballads, Books, and Newspapers: Biographical Notes and Checklists for Nathaniel Coverly, Sr., Nathaniel Coverly, Jr., and Joseph White. Printing History (New Series), (7):??, January 2010. ISSN 0192-9275.

Koda:2010:BRN

[Kod10b] Paul S. Koda. Book review:

Nancy Finlay, ed. Picturing Victorian America: Prints by the Kellogg Brothers of Hartford, Connecticut, 1830–1880. Printing History (New Series), (7):??, January 2010. ISSN 0192-9275.

Koda:2010:BRR

[Kod10c] Paul S. Koda. Book review:

Roberta A. Gross and Mary Kelley, eds., A History of the Book in America. Volume 2: An Extensive Republic: Print, Culture, and Society in the New Nation, 1790–1840. Printing History (New Series), (8):??, July 2010. ISSN 0192-9275.

Koda:2011:BRM

[Kod11] Paul S. Koda. Book review:

Michael F. Suarez, S.J., and H. R. Woudhuysen, eds. The Oxford Companion to the Book. Printing History (New Series), (10):??, July 2011. ISSN 0192-

Korey:1983:BRM

[Kor83] Marie Elena Korey. Book revew: Margaret Spufford, Small Books and Pleasant Histories: Popular Fiction and Its Readership in Seventeenth-Century England. Printing History, 5(2):??, ???? 1983. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-10-
0.

Kosofsky:2006:SCB

[Kos06] Scott-Martin Kosofsky. The SP century: Boston’s Society of Printers through one hundred years of change. Oak Knoll Books, New Castle, DE, USA, 2006. ISBN 1-58456-194-7. xx + 266 pp. LCCN Z120.S73 S65 2006.

Kosofsky:2010:BRY

[Kos10] Scott-Martin Kosofsky. Book review: Yosef Goldman, Hebrew Printing in America 1735–1926: A History and Annotated Bibliography. Printing History (New Series), (8):??, July 2010. ISSN 0192-9275.

Kaestle:2009:HBA

[KR09] Carl F. Kaestle and Janice A. Radway. A History of the Book in America: Volume 4: Print in Motion: The Expansion of Publishing and Reading in the United States, 1880–1940. University of North Carolina Press, Chapel Hill, NC, USA, 2009. ISBN 0-8078-3186-7. xvii + 669 pp. LCCN Z473 .P75 2009.

Kraus:1979:BRJ

[Kra79a] Joe W. Kraus. Book review: Joseph Blumenthal, The Printed Book in America. Printing History, 1(1):??, ???? 1979. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-1-
0.

Kraus:1979:MCD

[Kra79b] Joe Walker Kraus. Messrs.

Copeland and Day, 69 Cornhill, Boston, 1893–1899. G. S. MacManus Co., Philadelphia, PA, USA, 1979. ISBN ???? xii + 179 pp. LCCN Z473.C76 K7 1979.

Krivatsy:1982:BWG

[Kri82] Nati Krivatsy. Bibliography of the works of Gregorio Leti. Oak Knoll Books, New Castle, DE, USA, 1982. ISBN 0-938768-06-9. v + 117 + 5 pp. LCCN Z8504.2 .K74 1982 DG465.7.L4.

Kristensen:2010:BBB

[Kri10] John Kristensen. The book [broadside, bookplate, business card, and birth announcement] beautiful. Printing History (New Series), (7):??, January 2010. ISSN 0192-9275.

Kristensen:2017:MJT

[Kri17] John Kristensen. The Merrymount Janson type and matrices. Printing History (New Series), (21):??, Winter 2017. ISSN 0192-9275.

Krummel:1986:CMP

[Kru86] D. W. Krummel. Clarifying the musical page: The romantic Stichbild. Printing History, 8(2):??, ???? 1986.

LaMoy:2013:FFS

[La 13] William T. La Moy. Frederic Fairchild Sherman and his Goudy typefaces. Printing History (New Series), (13):??, January 2013. ISSN 0192-9275.

LaMoy:2014:BRK

[La 14] William T. La Moy. Book review: Keri Yousi, Balzac, Grandville, and the Rise of Book Illustration. Printing History (New Series), (16):??, July 2014. ISSN 0192-9275.

LaMoy:2015:SEL

[La 15] William T. La Moy. A short essay and a letter by Christopher Morely on publishing matters transcribed and annotated. Printing History (New Series), (17):??, January 2015. ISSN 0192-9275.

LaMoy:2016:ALS

[La 16a] William T. La Moy. An autographed letter (signed) from Frederic W. Goudy to [Alpheus] Sherwin Cody dated 18 August 1935. Printing History (New Series), (19):??, January 2016. ISSN 0192-9275.

LaMoy:2016:FWG

[La 16b] William T. La Moy. Frederic W. Goudy’s formal tribute to Bertha M. Goudy. Printing History (New Series), (19):??, January 2016. ISSN 0192-9275.

Levenson:1994:WPN

[LA94] Roger Levenson and Susan Acker. Women in printing: Northern California, 1857– 1890: selected excerpts. Feathered Serpent, ????, CA, USA, 1994. 4 unnumbered pp. LCCN ????

Labovitz:2017:ETO

[Lab17] John Labovitz. The electric typesetter: The origins of computing in typography. Printing History (New Series), (21):??, Winter 2017. ISSN 0192-9275. URL http:/

/web.mit.edu/Saltzer/www/
publications/CC-244.html;
 http://www.dpbsmith.com/
tj2.html;   https://johnlabovitz.
com/publications/The-electric-
typesetter--The-origins-
of-computing-in-typography.
pdf;   https://printinghistory.
org/publications/printing-
history/ns/#21.

Lancaster:1983:NCA

[Lan83] John Lancaster. Nineteenth Century America: Book trade technology and social history. [introduction]. Rollo G. Silver, The Power of the Press: Hand, Horse, Water, and Steam. Printing History, 5(1): ??, ???? 1983. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-9-

Lange:2002:FOM

[Lan02] Gerald Lange. Fraternal offspring: Matthew Carter’s Manutius/Miller. Printing History, 21(2):??, Spring 2002. ISSN 0192-9275.

                                        URL https://www.rit.edu/
                                        press/printing-history-42-
                                        0.

Lane:2004:ETS

[Lan04] John A. Lane. Early type specimens in the Plantin–Moretus Museum: annotated descriptions of the specimens to ca. 1850 (mostly from the Low Countries and France) with preliminary notes on the typefoundries and printing offices. Oak Knoll Press, New Castle, DE, USA, 2004. ISBN 1-58456-139-4. 344 pp. LCCN Z250.A2 L36 2004.

Lane:2005:GPP

[Lan05] John A. Lane. Garamond premier pro: a contemporary adaptation: modeled on the roman types of Claude Garamond and the italic types of Robert Granjon. Adobe Systems, San Jose, CA, USA, 2005. 44 pp. LCCN ????

Lankiewicz:2016:MKA

[Lan16]  Donald Lankiewicz.  Mein

Kampf in America: How Hitler came to be published in the United States. Printing History (New Series), (20):??, July 2016. ISSN 0192-9275.

Larson:1984:SPE

[Lar84] Judy L. Larson. Separately published engravings in the Early Republic: An introduction to copperplate engraving and printing in America through 1820. Printing History, 6(1):??, ???? 1984. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-11-
0.

Larson:1992:BRJ

[Lar92] Jennifer E. Larson. Book review: John Carter and Graham Pollard, An Enquiry into the Nature of Certain Nineteenth Century Pamphlets, 2d ed. with new material; Nicholas Barker and John Collins, A Sequel to an Enquiry: The Book Forgeries of H. Buxton Forman and T. J. Wise Re-examined; John Collins, The Two Forgers: A Biography of Harry Buxton Forman & Thomas James Wise. Printing History, 14(2): ??, ???? 1992.

Last:2005:CEN

[Las05]  Jay T. Last.  The Color

Explosion: Nineteenth-Century American Lithography. Hillcrest Press, Santa Ana, CA, USA, 2005. ISBN 0-914589-11-3. 316 pp. LCCN NE2303.7 .L37 2005.

Lawson:1971:PTI

[Law71] Alexander S. Lawson. Printing Types: an Introduction. Beacon Press, Boston, MA, USA, 1971. ISBN 0-8070-6658-3. 119 pp. LCCN Z250 .L35.

Lawson:1984:TT

[Law84] Alexander Lawson. Teaching typography. Printing History, 6(2):??, ???? 1984.

Lawson:1990:AT

[Law90a] Alexander S. Lawson. Anatomy of a Typeface. Godine, Boston, MA, USA, 1990. ISBN 0-87923-332-X (hardcover), 0-87923-333-8 (paperback). 428 pp. LCCN Z250 .L34 1990.

Lawson:1990:BRD

[Law90b] Alexander S. Lawson. Book review: D. J. R. Bruckner, Frederic Goudy. Printing History, 13(2):??, ???? 1990.

Lawson:1990:CAC

[Law90c] Alexander S. Lawson. The Compositor as Artist, Craftsman, and Tradesman. Press of the Nightowl, Athens, GA, USA, 1990. ISBN 0-912960-17-5. 33 pp. LCCN Z242.C73 L3 1990.

Leary:2010:PBT

[Lea10] Patrick Leary. The Punch Brotherhood: table talk and print culture in Mid-Victorian London. The British Library, London, UK, 2010. ISBN 0-7123-0923-3, 0-7123-6366-1 (ebook). x + 197 + 24 pp. LCCN PN5130.P8 L43 2010eb.

Lee:1989:OIM

[Lee89] Jennifer B. Lee. “Our infant manufactures”: Early typefounding in Philadelphia. Printing History, 11(2):??, ???? 1989.

Lee:20xx:EBL

[Leexxa]  Jennifer B. Lee.  Editor,

The Bullen Letters. Printing History, 22(1–2):??, ???? 20xx. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-43-
-44-0.

Lee:20xx:IE

[Leexxb]  Jennifer B. Lee.  Intro-

duction to the Exhibition. Printing History, 22(1–2):??, ???? 20xx. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-43-
-44-0.

Lent:1988:PWE

[Len88] John A. Lent. Pioneer women editors — the Stockdale sisters of Bermuda. Printing History, 10(1):??, ???? 1988. See comments [Cav88].

Leuthold:1994:BPV

[Leu94] Steven Leuthold. The book and the peasant: Visual representation and social change in German woodcuts, 1521– 1525. Printing History, 17(2): ??, ???? 1994. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-34-
0.

Lieberman:1978:TT

[Lie78] J. Ben Lieberman. Type and typefaces. A Treasury of typography book. Myriade Press, New Rochelle, NY, USA, second edition, 1978. ISBN 0-918142-01-6 (hardcover), 0-918142-02-4 (paperback). 142 pp. LCCN Z250 .L6 1977.

Lieberman:1979:BRE

[Lie79] J. Ben Lieberman. Book review: Elizabeth Harris and Clinton Sisson, The Common Press. Printing History, 1(2): ??, ???? 1979. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-2-
0.

Liska:1980:BRJ

[Lis80] Christine Liska. Book review: Joe W. Kraus, Messrs. Copeland & Day. Printing History, 2(2):??, ???? 1980. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-4-
0.

Lommen:2006:SHM

[LL06] Mathieu Lommen and Nynke Leistra. Sem Hartz and the making of Linotype Juliana. Inferno Press, Vancouver, BC, Canada, 2006. ISBN 0-9732652-4-8. 13 unnumbered pp. LCCN ????

Loxley:2004:TSH

[Lox04] Simon Loxley. Type: the secret history of letters. I. B. Tauris, London, UK, 2004. ISBN 1-85043-397-6, 1-84511-028-5 (paperback), 1-4175-5698-6 (ebook), 0-85771-761-8, 1-282-69501-0, 0-85773-017-7. vii + 248 pp. LCCN Z250.A2 L6 2004.

Loxley:2008:FWC

[Lox08a] Simon Loxley. Frederic Warde, Crosby Giage, and the Watch Hill Press. Printing History (New Series), (4):??, July 2008. ISSN 0192-9275.

Loxley:2008:PDL

[Lox08b] Simon Loxley. Printer’s devil:

the life and work of Frederic Warde. David R. Godine, Boston, MA, USA, 2008. ISBN 1-56792-367-4 (hardcover). ???? pp. LCCN Z246.W357 L69 2008.

Loney:2010:BC

[LQS10] Alan Loney, Jenni Quilter, and Kyle Schlesinger. The books to come. Cuneiform Press, Victoria, TX, USA, 2010. ISBN 0-9827926-2-X (hardcover). 134 pp. LCCN PR9639.3.L6 B66 2010.

Landon:2007:HLB

[LRS07] Richard Landon, Bruce Rogers, and Thomas T. Schweitzer. Humane letters: Bruce Rogers, designer of books and artist. Thomas Fisher Rare Book Library, Toronto, ON, Canada, 2007. ISBN 0-7727-6063-2. 112 pp. LCCN Z232.R67 L35 2007.

MacKitterick:1979:SC

[Mac79] David J. (David John) MacKitterick. Selected correspondence. Scolar Press, London, UK, 1979. ISBN 0-85967-589-0 (Scolar Press), 0-89679-001-0 (Moretus). xxvi + 217 pp. LCCN Z232.M85 A36.

Macmillan:2006:ZTD

[Mac06] Neil Macmillan. An A–Z of type designers. Yale University Press, New Haven, CT, USA, 2006. ISBN 0-300-11151-7. 200 pp. LCCN Z250.A2 A19 2006.

Makala:2019:SMA

[Mak19] Jeffrey Makala. ‘Spiritual machinery : The American Bible Society and the mechanisms of large-scale printing in the Early Nineteenth Century. Printing History (New Series), (25):??, Winter 2019. ISSN 0192-9275.

Marker:1985:PPO

[Mar85]  Gary Marker.  Publishing,

printing, and the origins of intellectual life in Russia, 1700– 1800. Princeton University Press, Princeton, NJ, USA, 1985. ISBN 0-691-05441-X. xiii + 301 pp. LCCN Z368 .M27 1985.

Margolin:1987:PTC

[Mar87] Victor Margolin. Paul Theobald & Company: Publisher with a new vision. Printing History, 9 (2):??, ???? 1987.

Markham:1993:MMS

[Mar93] Sandra J. Markham. Memento mori on silk and stone: Reuben Manley, printer, 1818– 42. Printing History, 15(1):??, ???? 1993.

Mardersteig:2001:CAD

[Mar01] Martino Mardersteig. The Cento Amici del Libro. Printing History, 21(1):??, ???? 2001. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-41-
0.

Maryatt:2007:EGP

[Mar07] Kitty Maryatt. Experience Gutenberg Project: Printing Beorum II at the Scripps College Press. Printing History (New Series), (1):??, January 2007. ISSN 0192-9275.

Mattson:1982:BRG

[Mat82] Francis O. Mattson. Book review: Giovanni Mardersteig, The Officina Bodoni: An Account of the Work of a Hand Press, 1923–1977. Printing History, 4(1–2):??, ???? 1982. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-7-
-8-0.

Mayo:1987:BRW

[May87] Hope Mayo. Book review:

William S. Peterson, A Bibliography of the Kelmscott Press. Printing History, 9(1):??, ???? 1987. MazurThomson:2015:ATI

[Maz15] Ellen Mazur Thomson. Aesthetic tracts. Innovation in late-nineteenth-century book design. Oak Knoll Books, New Castle, DE, USA, 2015. ISBN 1-58456-336-2. xiv + 177 + 16 pp. LCCN Z4-8.

McCorison:1995:JMI

[McC95] Marcus A. McCorison. John Mycall — the ingenious typographer of Newburyport. Printing History, 17(1):??, ???? 1995. ISSN 0192-9275.

                                        URL https://www.rit.edu/
                                        press/printing-history-33-
                                        0.

McGuinne:1994:VHI

[McG94] Dermot McGuinne. Victor Hammer — an Irish connection. Printing History, 16(1–2): ??, ???? 1994.

McKinstry:1979:SBP

[McK79] E. Richard McKinstry. A sketch of the brief printing career of David Cree, 1784– 1786. Printing History, 1(2): ??, ???? 1979. ISSN 0192-9275.

                                        URL https://www.rit.edu/
                                        press/printing-history-2-
                                        0.

McTigue:1981:EBC

[McT81]  Bernard McTigue.  En-

glish bookbinding and the Continental woodblock tradition. Printing History, 3(2):??, ???? 1981. ISSN 0192-9275.

                                        URL https://www.rit.edu/
                                        press/printing-history-5-
                                        0.

McTigue:1984:BRS

[McT84] Bernard McTigue. Book review: Susan E. Meyer, A Treasury of the Great Children’s Book Illustrators. Printing History, 6(1):??, ???? 1984. ISSN 0192-9275.

McTigue:1985:BRB

[McT85] Bernard McTigue. Book review: Brigid Peppin and Lucy Micklethwait, Book Illustrators of the Twentieth Century. Printing History, 7(1):??, ???? 1985.

Metzger:1981:BRB

[Met81] Philip A. Metzger. Book review: The Business of Enlightenment: A Publishing History of the Encyclopedie, 1775– 1800. Printing History, 3(2): ??, ???? 1981. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-5-
0.

Meyer:1983:TGC

[Mey83] Susan E. Meyer. A treasury of the great children’s book illustrators. Abrams, New York, NY, USA, 1983. ISBN 0-8109-0782-8. 272 pp. LCCN NC965 .M49 1983.

Middleton:1985:RRH

[Mid85] R. Hunter (Robert Hunter)

Middleton. RHM, Robert Hunter Middleton, the man and his letters: eight essays on his life and career. Caxton Club, Chicago, IL, USA, 1985. ISBN 0-8405-5008-3. ix + 100 pp. LCCN Z250.A2 M657 1985.

Mifflin:2012:BRD

[Mif12a] Jeffrey Mifflin. Book review:

David Pearson, Books as History: The Importance of Books Beyond Their Texts. Printing History (New Series), (12):??, July 2012. ISSN 0192-9275.

Mifflin:2012:BRM

[Mif12b] Jeffrey Mifflin. Book review: Martin Hopkinson, Ex Libris: The Art of Bookplates. Printing History (New Series), (12): ??, July 2012. ISSN 0192-9275.

Mifflin:2013:BRP

[Mif13] Jeffrey Mifflin. Book review: Patrick Leary, The “Punch” Brotherhood: Table Talk and Print Culture in Mid-Victorian London. Printing History (New Series), (13):??, January 2013. ISSN 0192-9275.

Mifflin:2015:BRS

[Mif15] Jeffrey Mifflin. Book review: Sue Rainey, Creating A World on Paper: Harry Fenn’s Career in Art. Printing History (New Series), (17):??, January 2015. ISSN 0192-9275.

Mifflin:2016:BRF

[Mif16a] Jeffrey Mifflin. Book review: Frank Romano, History of the Linotype Company. Printing History (New Series), (19):??, January 2016. ISSN 0192-9275.

Mifflin:2016:BRG

[Mif16b]  Jeffrey Mifflin.  Book re-

view: Gill Partington and Adam Symth, Ed, Book Destruction from the Medieval to the Contemporary. Printing History (New Series), (20):??, July 2016. ISSN 0192-9275.

Mifflin:2016:BRR

[Mif16c]  Jeffrey Mifflin.  Book re-

view: Richard L. Hopkins, Tolbert Lanston and the Monotype: The Origin of Digital Typsetting. Printing History (New Series), (19):??, January 2016. ISSN 0192-9275.

Mifflin:2016:BRS

[Mif16d] Jeffrey Mifflin. Book review:

Simon Loxley, Printer’s Devil: The Life and Work of Frederic Warde. Printing History (New Series), (20):??, July 2016. ISSN 0192-9275.

Mifflin:2017:BRE

[Mif17a] Jeffrey Mifflin. Book review:

Ellen Mazur Thomson, Aesthetic Tracts: Innovation in Late-Nineteenth-Century Book Design. Printing History (New Series), (21):??, Winter 2017. ISSN 0192-9275.

Mifflin:2017:BRM

[Mif17b] Jeffrey Mifflin. Book review:

Mark R. Godburn, Nineteenth-Century Dust Jackets. Printing History (New Series), (22):??, Summer 2017. ISSN 0192-9275.

Miller:1989:PDCa

[Mil89a] Sarah Jordan Miller. Producing documents for Congress and the nation: Government printing in the United States, past and present. Part I. Printing History, 11(1):??, ???? 1989.

Miller:1989:PDCb

[Mil89b] Sarah Jordan Miller. Producing documents for Congress and the nation: Government printing in the United States, past and present. Part II. Reviews.

Printing History, 11(2):??, ???? 1989.

Mortimer:1981:DRT

[Mor81] Ruth Mortimer. The dimensions of the Renaissance title page. Printing History, 3(1): ??, ???? 1981. ISSN 0192-9275.

                                        URL https://www.rit.edu/
                                        press/printing-history-5-
                                        0.

Mortimer:1991:BSA

[Mor91] Ruth Mortimer. The Bibliographical Society of America: Publishing about printing. Printing History, 14(1):??, ???? 1991.

Mosley:2007:HTT

[Mos07] James Mosley. Handmade type: Thoughts on the preservation of typographic materials. Incline Press, Oldham, UK, 2007. ii + 23 pp. LCCN Z250.A3 M854 2007. The First Justin Howes Memorial Lecture given at the St Bride Foundation on 21 February 2006.

Mosley:2008:MTL

[Mos08] James Mosley. The materials of typefounding: A list of surviving collections. Printing History (New Series), (4):??, July 2008. ISSN 0192-9275.

Morris:1982:IBE

[MP82] William Morris and William S. Peterson. The ideal book: essays and lectures on the arts of the book. University of California Press, Berkeley, CA, USA, 1982. ISBN 0-520-04563-7. xxxix + 134 pp. LCCN Z116.A3 M65425 1982.

Meynellcomp:1973:FA

[MS73] Francis Meynell and Herbert Simon. Fleuron: anthology. E. Benn, London, UK, 1973. ISBN 0-510-41850-3, 0-8020-2080-1. xv + 359 pp. LCCN Z123 .M47.

Mardersteig:1980:OBA

[MS80] Giovanni Mardersteig and Hans Schmoller, editors. The Officina Bodoni: an account of the work of a hand press, 1923– 1977. Edizioni Valdonega, Verona, Italy, 1980. ISBN 88-85033-05-9. lix + 285 pp. LCCN Z232.O33 M3713 1980b.

Mitchell:2005:BTD

[MW05] Michael Mitchell and Susan Wightman. Book typography: a designer’s manual. Libanus Press, Marlborough, Wiltshire, 2005. ISBN 0-948021-66-7. xi + 434 pp. LCCN Z246 .M57 2005.

Neavill:1979:MLS

[Nea79]  Gordon B. Neavill.  The

modern library series: Format and design, 1917–1977. Printing History, 1(1):??, ???? 1979. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-1-
0.

Neavill:2016:IML

[Nea16] Gordon Neavill. The illustrated modern library series. Printing History (New Series), (20):??, July 2016. ISSN 0192-9275.

Needham:1986:PPU

[Nee86] Paul Needham. The printer and the pardoner: an unrecorded indulgence printed by William Caxton for the Hospital of St. Mary Rounceval, Charing Cross. Library of Congress, Washington, DC, USA, 1986. ISBN 0-8444-0508-6. 101 + 2 pp. LCCN Z232.C38 N44 1986.

Nesbitt:1980:LTL

[Nes80] Alexander Nesbitt. A life with type and letters. Printing History, 2(2):??, ???? 1980. ISSN 0192-9275.

                                        URL https://www.rit.edu/
                                        press/printing-history-4-
                                        0.

Nesbitt:1984:BRM

[Nes84] Alexander Nesbitt. Book review: Marc Drogin. Anathema: Medieval Scribes and the History of Book Curses. Printing History, 6(2):??, ???? 1984.

Newman:1984:BRD

[New84] Richard Newman. Book review: Donald Franklin Joyce. Gatekeepers of Black Culture: Black-Owned Book Publishing in the United States, 1817– 1981. Printing History, 6(2):??, ???? 1984.

OBrien:2003:ENC

[O’B03] Donald C. O’Brien. The early Nineteenth-Century Boston engraving trade and the engravers who developed it. Printing History, 23(1):??, ???? 2003. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-45-
0.

OBrien:2014:EPP

[O’B14] Donald C. O’Brien. The engravers of Philadelphia’s Port Folio magazine. Printing History (New Series), (15):??, January 2014. ISSN 0192-9275.

Oldham:2006:FGN

[Old06] Robert W. Oldham. A field guide to North American hand presses and their manufacturers. Ad Lib Press, Doswell, VA, USA, 2006. ISBN 1-4243-2901-9. iv + 128 pp. LCCN Z249 .O43 2006.

Osley:1980:SSH

[OW80] A. S. Osley and Berthold Wolpe. Scribes and sources: handbook of the chancery hand in the sixteenth century. David R. Godine, Boston, MA, USA, 1980. ISBN 0-87923-297-8. 291 pp. LCCN Z43.A3.

Palmieri:2019:BRH

[Pal19a] Brooke Sylvia Palmieri. Book review: Hendrik D. L. Vervliet, Robert Granjon, Letter-Cutter, 1513–1590: Oeuvre Catalogue. Printing History (New Series), (26):??, Summer 2019. ISSN 0192-9275.

Palmieri:2019:BRM

[Pal19b] Brooke Sylvia Palmieri. Book review: Michalis Pichler, Publishing Manifestos. Printing History (New Series), (25):??, Winter 2019. ISSN 0192-9275.

Pantzer:1982:ESR

[Pan82] Katharine F. Pantzer. Ephemera in the STC revision: A housekeeper’s view. Printing History, 4(1–2):??, ???? 1982. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-7-
-8-0.

Pankow:1985:BRJ

[Pan85]  David Pankow.  Book re-

view: John Buechler, Charles Whittingham, Printer, 1795– 1876; James Davis, Printed by Hague and Gill; Vincent Torre, A Tribute to W. A. Dwiggins on the Hundredth Anniversary of His Birth; Gay Walker, The Works of Carl P. Rollins. Printing History, 7(1):??, ???? 1985.

Pankow:1988:DDB

[Pan88] David Pankow. Dungeons and dragon’s blood: The development of late 19th and early 20th Century platemaking processes. Printing History, 10(1): ??, ???? 1988.

Pankow:1998:FON

[Pan99] David Pankow. A face by any other name is still my face: A tale of type piracy. Printing History, 19(1):??, ???? 1998–1999. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-37-
0.

Pankow:2007:FUV

[Pan07] David Pankow. Facing the unfolded and visible book of the future. Printing History, 25(2): ??, ???? 2007. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-50-
0.

Pankow:20xx:EIR

[Panxx] David Pankow. Editor’s introduction: The rise and fall of ATF. Printing History, 22(1–2):??, ???? 20xx. ISSN 0192-9275.

URL  http://everything.
explained.today/American_
Type_Founders/#Later%20mergers%
20and%20acquisitions;   https:
//www.rit.edu/press/printing-
history-43--44-0.

Parker:1994:WSB

[Par94] Mike Parker. W. Starling Burgess, type designer? Printing History, 16(1–2):??, ???? 1994. See rebuttal [BBRD99].

Pearson:2008:BHI

[Pea08] David Pearson. Books as history: the importance of books beyond their texts. British Library, London, UK, 2008. ISBN 1-58456-233-1 (hardcover), 0-7123-4923-5. 207 pp. LCCN Z4 .P43 2008.

Peich:1998:WEF

[Pei98] Michael Peich. William Everson: Fine printer. Printing History, 18(2):??, ???? 1998. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-36-
0.

Perry:2015:EDP

[Per15] Timothy P. J. Perry. Early depictions of the printing press. Printing History (New Series), (18):??, July 2015. ISSN 0192-9275.

Peterson:1981:WMD

[Pet81] William S. Peterson. William Morris and the “damned chemists”: The search for an ideal ink at the Kelmscott Press. Printing History, 3(2): ??, ???? 1981. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-5-
0.

Peterson:1984:BKP

[Pet84] William S. Peterson. A bibliography of the Kelmscott Press, volume 24 of The Soho bibliographies. Clarendon Press, New York, NY, USA, 1984. ISBN 0-19-818199-X. xliii + 217 pp. LCCN Z232.M87 P44 1984.

Peterson:1985:BRJ

[Pet85] William S. Peterson. Book review: John Carter and Graham Pollard, An Enquiry into the Nature of Certain Nineteenth Century Pamphlets, 2d ed.; Nicolas Barker and John Collins, A Sequel to an Enquiry into the Nature of Certain Nineteenth Century Pamphlets by John Carter and Graham Pollard: The Forgeries of H. Buxton Forman and T. J. Wise Re-examined. Printing History, 7(1):??, ???? 1985.

Petro:1987:GGR

[Pet87] Pamela Petro. Gwasg Gregynog: The reincarnation of a press. Printing History, 9(2): ??, ???? 1987.

Peterson:2007:NCR

[Pet07a] William S. Peterson. Nineteenth-Century revivals, typographic and spiritual. Printing History, 25(2):??, ???? 2007. ISSN 0192-9275. URL https://www.

rit.edu/press/printing-history-
50-0.

Petrella:2007:UTL

[Pet07b] Giancarlo Petrella. Uomini, torchi e libri nel Rinascimento. (Italian) [Men, presses and books in the Renaissance], volume 18 of Libri e biblioteche. Forum, Udine, Italy, 2007. ISBN 88-8420-404-6. 355 pp. LCCN Z8.I8 P47 2007.

Phelps:1985:FPU

[Phe85] C. Deirdre Phelps. The first publication to use Americanmade type. Printing History, 7 (1):??, ???? 1985.

Peppin:1984:BIT

[PM84] Brigid Peppin and Lucy Micklethwait. Book illustrators of the twentieth century. Arco Publishing, New York, NY, USA, 1984. ISBN 0-668-05670-3 (hardcover). 336 pp. LCCN NC978 .P46 1984.

Pomeroy:1994:CMW

[Pom94] Jane R. Pomeroy. On the changes made in wood engravings in the stereotyping process. Printing History, 17(2): ??, ???? 1994. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-34-
0.

Pomeroy:2005:AAW

[Pom05] Jane R. Pomeroy. Alexander Anderson (1775–1870): wood engraver and illustrator: an annotated bibliography. Oak Knoll Press, New Castle, DE, USA, 2005. ISBN 1-58456-162-9. lxxxvi + 2514 (3 volumes) pp. LCCN NE1112.A52 P66 2005; Z8035.35 .P65 2005.

Peterson:2011:KCC

[PP11] William S. (William Samuel) Peterson and Sylvia Holton Peterson. The Kelmscott Chaucer: a census. Oak Knoll Press, New Castle, DE, USA, 2011. ISBN 1-58456-289-7. xvi + 280 pp. LCCN Z8164 .P47 2011.

Pretzer:1981:BRC

[Pre81] William S. Pretzer. Book review: Comp. and ed. by John Gibson and Laurie Lewis, Sticks and Stones: Some Aspects of Canadian Printing History. Printing History, 3(2):??, ???? 1981. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-5-
0.

Pretzer:1984:TPC

[Pre84] William S. Pretzer. Tramp printers: Craft culture, trade unions, and technology. Printing History, 6(2):??, ???? 1984.

Price:2012:S

[Pri12] Gary H. Price. Spaces. Printing History (New Series), (11):??, January 2012. ISSN 0192-9275.

Price:2013:PM

[Pri13] Gary H. Price. Printing mathematics. Printing History (New Series), (14):??, July 2013. ISSN 0192-9275.

Partington:2014:BDM

[PS14] Gill Partington and Adam Smyth, editors. Book destruction from the medieval to the contemporary. New directions in book history. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire, UK, 2014. ISBN 1-137-36765-2 (hardcover). xi + 216 pp. LCCN Z659 .B66 2014.

Raecke:1988:BRS

[Rae88] Renate Raecke. Book review:

Vom Schriftgiessen [Of Typefounding]. Porträt der Firma D. Stempel. Printing History, 10(1):??, ???? 1988.

Rafaeli:2005:BT

[Raf05] Ari Rafaeli. Book Typography.

Oak Knoll Press, New Castle, DE, USA, 2005. ISBN 1-58456-157-2 (Oak Knoll Press), 0-7123-0693-5 (British Library). 160 pp. LCCN Z246 .R25 2005.

Rainwater:1981:BRB

[Rai81]  Robert Rainwater.  Book

review: Basil Hunnisett, Steel-engraved Book Illustration in England. Printing History, 3(2):??, ???? 1981. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-5-
0.

Rainwater:1982:BRB

[Rai82] Robert Rainwater. Book review: Basil Hunnisett, A Dictionary of British Steel Engravers. Printing History, 4(1– 2):??, ???? 1982. ISSN 0192-

9275. URL https://www.rit.
edu/press/printing-history-
7--8-0.

Rainey:2013:CWP

[Rai13] Sue Rainey. Creating a world on paper: Harry Fenn’s career in art. University of Massachusetts Press, Boston, MA, USA, 2013. ISBN 1-55849-979-2, 1-61376-222-4 (e-book). xi + 392 pp. LCCN N6537.F39 R35 2013.

Rathe:1992:BTC

[Rat92] John F. Rathé. Bibliography of the Typophile chap books, 1935–1992, volume 60 of Typophile chap books. Typophiles, New York, NY, USA, 1992. ISBN 0-945074-02-6. 94 pp. LCCN ????

Ray:1982:AFI

[Ray82] Gordon N. (Gordon Norton) Ray. The art of the French illustrated book 1700–1914. Pierpont Morgan Library, New York, NY, USA, 1982. ISBN 0-8014-1535-7. xxxii + 557 + 16 pp. LCCN NC980.

Rutherfurd:1984:JPZ

[RB84] Livingston Rutherfurd and James B. Bell, editors. John Peter Zenger and freedom of the press. A. Colish, Mount Vernon, NY, USA, 1984. xx + 71 + 13 pp. LCCN Z232.Z5 R925 1984. With an introduction by Tom Wicker.

Ross:2008:NLT

[RB08] Fiona Ross and Rob Banham, editors. Non-Latin typefaces at St Bride Library, London, and Department of Typography and Graphic Communication, University of Reading. St Bride Library, London, UK, 2008. ISBN 0-9504161-5-0. 63 pp. LCCN Z250.A2 N66 2008.

Rehak:1993:PT

[Reh93] Theo Rehak. Practical Typecasting. Oak Knoll Books, New Castle, DE, USA, 1993. ISBN 0-938768-33-6. xvii + 221 pp. LCCN Z250 .R33 1993.

Reilly:1975:DCA

[Rei75] Elizabeth Carroll Reilly. A dictionary of colonial American printers’ ornaments and illustrations. American Antiquarian Society, Worcester, MA, USA, 1975. ISBN 0-912296-06-2. xxxvi + 515 pp. LCCN Z208 .R43.

Reardon:1991:CPP

[RK91] Tom Reardon and Kent Kirby. Collotype: Prince of the printing processes. Printing History, 14(1):??, ???? 1991.

Romano:2003:PE

[Rom03] Frank Romano. The phototypesetting era. Printing History, 23(2):??, ???? 2003. ISSN 0192-9275.

                                        URL https://www.rit.edu/
                                        press/printing-history-46-
                                        0.

Romano:2014:HLC

[Rom14] Frank J. Romano. History of the Linotype Company. RIT Press, Rochester, NY, USA, 2014. ISBN 1-939125-03-0 (hard cover), 1-933360-60-7 (paperback), 1-933360-75-5 (ebook). xv + 463 pp. LCCN Z253 .R736 2014.

Roylance:1986:EGA

[Roy86]  Dale Roylance.  European

graphic arts: the art of the book from Gutenberg to Picasso. Princeton University Library, Princeton, NJ, USA, 1986. vii + 189 pp. LCCN Z116.A3 R68 1986.

Ruffin:2017:BRR

[Ruf17] Katherine M. Ruffin. Book review: Richard-Gabriel Rummonds, Fantasies and Hard Knocks: My Life as a Printer. Printing History (New Series), (21):??, Winter 2017. ISSN 0192-9275.

Ruffin:2019:BRJ

[Ruf19] Katherine M. Ruffin. Book review: James Raven, What is the History of the Book? Printing History (New Series), (26): ??, Summer 2019. ISSN 0192-9275.

Rumble:1992:TGS

[Rum92a]  Walker Rumble.  A time

of giants: Speed composition in Nineteenth-Century America. Printing History, 14(2):??, ???? 1992.

Rummonds:1992:BRS

[Rum92b] Richard Gabriel Rummonds.

Book review: Stephen O. Saxe, American Iron Hand Presses. Printing History, 14(2):??, ???? 1992.

Rumble:1998:SSI

[Rum98] Walker Rumble. Strategies of shopfloor inclusion: The gender politics of Augusta Lewis and Women’s Typographical Union No. 1, 1868–1872. Printing History, 18(1):??, March 1998. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-35-
0.

Rumble:2014:JDS

[Rum14] Walker Rumble. Jacobs, Dwiggins and the short life of Linotype Charter. Printing History (New Series), (15):??, January 2014. ISSN 0192-9275.

Rummonds:2015:FHK

[Rum15] Richard-Gabriel Rummonds.

Fantasies and hard knocks: my life as a printer. Ex Ophidia Press, Port Townsend, WA, USA, 2015. ISBN 0-692-36404-8 (hardcover). xviii + 795 pp. LCCN Z232.R94 A3 2015.

Rushmore:1990:WNW

[Rus90] Arthur W. Rushmore. When Noah was a pup. Printing History, 12(1):??, ???? 1990.

Ryder:1979:CL

[Ryd79] John Ryder. The case for legibility. Bodley Head, London, UK, 1979. ISBN 0-370-30158-7 (hardcover). 77 pp. LCCN Z116.A3 R96.

Samuelson:2014:SL

[Sam14] Todd Samuelson. Still life. Printing History (New Series), (16):??, July 2014. ISSN 0192-9275.

Savage:2016:MSF

[Sav16] Elizabeth Savage. The mystery of the “scrappy fragments”: Untangling Robert Steele’s discovery of frisket sheets. Printing History (New Series), (19): ??, January 2016. ISSN 0192-9275.

Saxe:1980:BRF

[Sax80a]  Stephen O. Saxe.  Book

review: Frank E. Comparato, Chronicles of Genius and Folly: R. Hoe & Company and the Printing Press as a Service to Democracy. Printing History, 2(2):??, ???? 1980. ISSN 0192-9275.

URL https://www.rit.edu/

press/printing-history-4-
0.

Saxe:1980:TFN

[Sax80b]  Stephen O. Saxe.  The

type founders of New York City, 1840–1900. Printing History, 2(1):??, ???? 1980. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-3-
0.

Saxe:1981:BHG

[Sax81] Stephen O. Saxe. A brief history of Golding & Co. Printing History, 3(2):??, ???? 1981. ISSN 0192-9275. URL https:/

/d3eb7xfyht02et.cloudfront.
net/wp-content/uploads/2011/
12/saxe-golding-ph-v6-pp12-
191.pdfhttps://www.rit.edu/
press/printing-history-5-
0.

Saxe:1986:BRP

[Sax86] Stephen O. Saxe. Book review: From Punch to Printing Type: The Art and Craft of Hand Punchcutting and Typecasting. Videotape. Printing History, 8(1):??, ???? 1986.

Saxe:1989:RPP

[Sax89] Stephen O. Saxe. Ramage Proof Press. Printing History, 11(2):??, ???? 1989.

Saxe:1990:FCP

[Sax90] Stephen O. Saxe. “Franklin”

Common Press. Printing History, 12(1):??, ???? 1990.

Saxe:1991:GCP

[Sax91] Stephen O. Saxe. The Goodman Common Press: The oldest American-made press. Printing History, 14(1):??, ???? 1991.

Saxe:1993:LVM

[Sax93] Stephen O. Saxe. The Landis Valley Museum Ramage Press. Printing History, 15(1):??, ???? 1993.

Saxe:1994:SOP

[Sax94] Stephen O. Saxe. “A small old printing press”. Printing History, 16(1–2):??, ???? 1994.

Saxe:2007:BRJ

[Sax07] Stephen O. Saxe. Book review: James Mosley, Handmade Type: Thoughts on the Preservation of Typographic Materials. Printing History (New Series), (2):??, July 2007. ISSN 0192-9275.

Schreiber:1982:EAC

[SB82] Fred Schreiber and Nicolas Barker. The Estiennes: an annotated catalogue of 300 highlights of their various presses. E. K. Schreiber, New York, NY, USA, 1982. ISBN ???? ix + 284 pp. LCCN Z232.E8 S36 1982.

Schreyer:1981:MTO

[Sch81]  Alice D. Schreyer.  Mark

Twain and “The Old-Fashioned Printer” [speech by Twain to the Typothetae]. Printing History, 3(2):??, ???? 1981. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-5-
0.

Schmets:1987:SPF

[Sch87] Ronald Schmets. Vom Schriftgiessen: Porträt der Firma D. Stempel, Frankfurt am Main. (German) [From the font: Portrait of the company D. Stempel, Frankfurt am Main]. Technische Hochschule Darmstadt, Darmstadt, West Germany, 1987. ISBN 3-88607-050-6. 16 + 84 pp. LCCN Z250.A2 D12 1987.

Schlesinger:1989:BOM

[Sch89] Carl Schlesinger. The biography of Ottmar Mergenthaler, inventor of the Linotype. Oak Knoll series on the history of the book. Oak Knoll Books, New Castle, DE, USA, 1989. ISBN 0-938768-12-3, 0-938768-13-1 (special edited). xviii + 125 pp. LCCN Z253 .M48 1989. With added historical notes based on recent findings, researched and edited by Carl Schlesinger ; introduction by Elizabeth Harris.

Schwenk:2017:BRW

[Sch17] Kim Schwenk. Book review:

William J. Kiesel, Printers Devices in Esoteric Publishing. Printing History (New Series), (22):??, Summer 2017. ISSN 0192-9275.

Schreiner:2018:PSH

[Sch18] Erin Schreiner. Printing the screenplay in Hollywood and beyond. Printing History (New Series), (24):??, Summer 2018. ISSN 0192-9275.

Schneider:2019:PIE

[Sch19] Nina Schneider. The printed intarsia of Erasmus Loya. Printing History (New Series), (26): ??, Summer 2019. ISSN 0192-9275.

Saxe:1991:AIH

[SD91] Stephen O. Saxe and John DePol. American iron hand presses. Yellow Barn Press, Council Bluffs, IA, USA, 1991. xii + 108 pp. LCCN Z249 .S23 1991.

Saxe:1992:AIH

[SD92] Stephen O. Saxe and John DePol. American iron hand presses. Oak Knoll Books, New Castle, DE, USA, 1992. ISBN 0-938768-35-2, 0-938768-36-0. xiv + 108 pp. LCCN Z249 .S23 1992.

Senchyne:2017:BRR

[Sen17] Jonathan Senchyne. Book review: Richard Kegler, The Aries Press of Eden, New York. Printing History (New Series), (21):??, Winter 2017. ISSN 0192-9275.

Saxe:1983:SG

[SFRM83] Stephen O. Saxe, Vincent Fitzpatrick, Frederick N. Rasmussen, and H. L. Mencken. In the steps of Gutenberg. Printing History, 5(1):??, ???? 1983. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-9-
0.

Shaw:2007:KTA

[Sha07] Matthew J. Shaw. Keeping time in the age of Franklin: Almanacs and the Atlantic world. Printing History (New Series), (2):??, July 2007. ISSN 0192-9275.

Shaw:2017:RTD

[Sha17] Paul Shaw. Revival Type:

Digital Typefaces Inspired by the Past. Yale University Press, New Haven, CT, USA, 2017. ISBN 0-300-21929-6 (hardcover). vii + 247 pp. LCCN Z250 .S533 2017.

Shaw:2018:BRC

[Sha18a] Paul Shaw. Book review: Caroline Archer-Parré and Malcolm Dick, eds. John Baskerville: Art and Industry of the Enlightenment. Printing History (New Series), (24):??, Summer 2018. ISSN 0192-9275.

Shaw:2018:PSR

[Sha18b] Paul Shaw. Paul Shaw’s response to Jerry Kelly’s review of Revival Type: Digital Typefaces Inspired by the Past (2017) published in Printing History 23 (Winter 2018). Printing History (New Series), (24):??, Summer 2018. ISSN 0192-9275.

Shields:2010:CRR

[Shi10] David Shields. Considering Rob Roy Kelly’s American Wood Type Collection. Printing History (New Series), (7): ??, January 2010. ISSN 0192-9275.

Silver:1979:ASP

[Sil79a] Rollo G. Silver. The autobiography of Stephen P. Ruggles, [partial transcription of an 1870 autobiography by inventor Ruggles (1808– 1880).]. Printing History, 1(1): ??, ???? 1979. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-1-
0.

Silver:1979:VAA

[Sil79b] Rollo G. Silver. Violent assaults on American printing shops, 1788–1860. Printing History, 1(2):??, ???? 1979. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-2-
0.

Silver:1982:BRR

[Sil82a] Rollo G. Silver. Book review: Robert B. Winans, A Descriptive Checklist of Book Catalogues Separately Printed in America, 1693– 1800. Printing History, 4(1–2): ??, ???? 1982. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-7-
-8-0.

Silver:1982:CPD

[Sil82b] Rollo G. Silver. The convivial printer: Dining, wining, and marching, 1825–1860. Alice D. Schreyer, the permanence of ephemera, [introduction]. Printing History, 4(1–2): ??, ???? 1982. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-7-
-8-0.

Silver:1984:BRS

[Sil84] Larry Silver. Book review: Sandra Hindman, ed. The Early Illustrated Book: Essays in Honor of Lessing J. Rosenwald. Printing History, 6(2):??, ???? 1984.

Singerman:1980:CDA

[Sin80] Robert Singerman. A curious document attributing a Jewish background to Johann Gutenberg. Printing History, 2(1):??, ???? 1980. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-3-
0.

Singerman:1992:NJN

[Sin92] Robert Singerman. Naphtali Judah, New York bookseller and stationer. Printing History, 14(2):??, ???? 1992.

Schafer:1987:FJB

[SIvA87] Otto Schäfer, Eduard Isphording, and Manfred von Arnim. Fünf Jahrhunderte Buchillustration: Meisterwerke der Buchgraphik aus der Bibliothek Otto Schäfer. (German) [Five Centuries of Book Illustration: Masterpieces of the book graphic from the Otto Schaefer Library], volume 42 of Ausstellungskataloge des Germanischen Nationalmuseums.

Serie A; Ausstellungskataloge / Bayerische Staatsbibliothek. Germanisches Nationalmuseum Nürnberg, Nürnberg, West Germany, 1987. ISBN 3-88607-053-0. xlviii + 188 + 30 pp. LCCN NC977 .F86 1987.

Smeijers:2003:TNM

[Sme03] Fred Smeijers. Type now: a manifesto: plus work so far. Hyphen Press, London, UK, 2003. ISBN 0-907259-24-3 (paperback). 142 pp. LCCN Z250.A2 S64 2003.

Smedresman:2008:GTC

[Sme08] Gabe Smedresman. Geofroy Tory’s Champ Fleury in the context of the Renaissance reconstruction of the Roman capital alphabet. Printing History (New Series), (3):??, January 2008. ISSN 0192-9275.

Smith:1986:LLT

[Smi86] Virginia Smith. Longevity and legibility: Two types from the De Vinne Press and how they have fared. Printing History, 8 (1):??, ???? 1986.

Smith:1988:BRD

[Smi88]  Virginia Smith.  Book re-

view: Dale Roylance, European Graphic Arts: The Art of the Book From Gutenberg to Picasso. Printing History, 10(1): ??, ???? 1988.

Smith:2005:FMU

[Smi05a] Virginia Smith. Forms in modernism: the unity of typography, architecture and the design arts 1920s–1970s. Random House Inc., New York, NY, USA, 2005. LCCN ????

Smith:2005:FMV

[Smi05b] Virginia (Virginia Grace St. George)

Smith. Forms in modernism: a visual set: the unity of typography, architecture and the design arts. Watson-Guptill Publications, 1515 Broadway, New York, NY 10036, USA, 2005. ISBN 0-8230-5624-4. 160 pp. LCCN NK1390 .S626 2005.

Smith:2015:MSF

[Smi15] Emma (Emma Josephine)

Smith. The Making of Shakespeare’s First Folio. Bodleian Library, University of Oxford, Oxford, UK, 2015. ISBN 1-85124-442-5 (hardcover). ix + 180 + 32 pp. LCCN PR3071 .S52 2015.

Somers:1986:IDF

[Som86] Jeanne Somers. Index to The Dolphin and The Fleuron, volume 9 of Art reference collection. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 1986. ISBN 0-313-25436-2. ISSN 0193-6867. xvii + 147 pp. LCCN Z119 .S614 1986.

Southall:2005:PTT

[Sou05] Richard Southall. Printer’s type in the twentieth century: manufacturing and design methods. British Library, London, UK, 2005. ISBN 0-7123-4812-3 (British Library: hardcover), 1-58456-155-6 (Oak Knoll Press). xv + 238 pp. LCCN Z250.A2 S68 2005.

Sparling:1983:BRG

[Spa83] Tobin A. Sparling. Book review: Gordon N. Ray, The Art of the French Illustrated Book, 1700–1914. Printing History, 5 (2):??, ???? 1983. ISSN 0192-9275.

Spufford:1982:SBP

[Spu82]  Margaret Spufford.  Small

Books and Pleasant Histories: Popular Fiction and Its Readership in Seventeenth-Century England. Cambridge University Press, Cambridge, UK, 1982. ISBN 0-8203-0595-2. xxi + 275 + 2 pp. LCCN PR972 .S65 1982.

Srebrnik:1986:ASV

[Sre86] Patricia Thomas Srebrnik. Alexander Strahan, Victorian publisher. University of Michigan Press, Ann Arbor, MI, USA, 1986. ISBN 0-472-10072-6. viii + 269 + 4 pp. LCCN Z325.S887 S68 1986.

Standard:1981:BRO

[Sta81]  Paul Standard.  Book re-

view: A. S. Osley, Scribes and Sources: A Handbook for the Chancery Hand of the Sixteenth Century. Printing History, 3(2):??, ???? 1981. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-5-
0.

Stansky:1985:RWW

[Sta85] Peter Stansky. Redesigning the world: William Morris, the 1880s, and the Arts and Crafts. Princeton University Press, Princeton, NJ, USA, 1985. ISBN 0-691-06616-7. xvi + 293 pp. LCCN NK1142 .S7 1985.

Stauffer:2010:LMP

[Sta10] Andrew M. Stauffer. Legends of the mummy paper. Printing History (New Series), (8): ??, July 2010. ISSN 0192-9275.

Stern:1979:BRM

[Ste79a] Madeleine B. Stern. Book review: Marjorie Dana Barlow, comp., Notes on Women Printers in Colonial America and the United States, 1639– 1975. Printing History, 1(1): ??, ???? 1979. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-1-
0.

Stern:1979:IMS

[Ste79b] Madeleine B. Stern. Isaac M. Singer’s type machine. Printing History, 1(2):??, ???? 1979. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-2-
0.

Stern:1980:BRL

[Ste80a] Madeleine B. Stern. Book review: Leona M. Hudak, Early American Women Printers and Publishers, 1639– 1820. Printing History, 2(1):

??, ???? 1980. ISSN 0192-9275.

                                        URL https://www.rit.edu/
                                        press/printing-history-3-
                                        0.

Stern:1980:EMH

[Ste80b] Madeleine B. Stern. Every man his own printer: The typographical experiments of Josiah Warren. Printing History, 2(2): ??, ???? 1980. ISSN 0192-9275.

                                        URL https://www.rit.edu/
                                        press/printing-history-4-
                                        0.

Stern:1980:PME

[Ste80c] Madeleine B. Stern, editor. Publishers for mass entertainment in nineteenth century America. G. K. Hall, Boston, MA, USA, 1980. ISBN 0-8161-8471-2. xxii + 358 pp. LCCN Z473 .S8575.

Stern:1985:ABU

[Ste85] Madeleine B. Stern. Antiquarian bookselling in the United States: a history from the origins to the 1940s. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 1985. ISBN 0-313-24729-3 (hardcover). xviii + 246 pp. LCCN Z479 .S79 1985.

Stillwell:1985:BRB

[Sti85] Dennis Stillwell. Book review: Bruce Beck, ed. Robert Hunter Middleton: The Man and His Letters. Printing History, 7(2): ??, ???? 1985.

Stinehour:1990:JB

[Sti90] Roderick Stinehour. Joseph Blumenthal. Printing History, 13(2):??, ???? 1990.

Stoddard:1987:MBA

[Sto87] Roger E. Stoddard. Morphology and the book from an American perspective. Printing History, 9(1):??, ???? 1987.

Stoddard:1998:OMJ

[Sto98] Roger E. Stoddard. “Oh, Mr.

Jefferson — after all these years, why do we know so little about the books 0f your time?”. Printing History, 18(1):??, March 1998. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-35-
0.

Stone:2005:APE

[Sto05] Reynolds Stone. The Albion Press, volume 16 of Printing Historical Society publication. Printing Historical Society, London, UK, 2005. ISBN 0-900003-13-8. 59–99 pp. LCCN ???? An essay first published in the Journal of the Printing Historical Society, no. 2, 1966, with supplements of 1967 and 1971. With an introductory note by James Mosley.

Stuckey:2018:TL

[Stu18] Amanda Stuckey. Tactile literacy. Printing History (New Series), (24):??, Summer 2018. ISSN 0192-9275.

Sullivan:1983:BRH

[Sul83]  Larry E. Sullivan.  Book

review: Henry R. Wagner & Charles L. Camp, The Plains & the Rockies, A Critical Bibliography of Exploration, Adventure and Travel in The American West, 1800– 1865. Printing History, 5(2): ??, ???? 1983. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-10-
0.

Sullivan:1984:BRA

[Sul84a] Larry E. Sullivan. Book review: Alan Dyer. A Biography of James Parker, Colonial Printer. Printing History, 6(2): ??, ???? 1984.

Sullivan:1984:OND

[Sul84b] Larry E. Sullivan. Old and new directions: Doctoral research on New York printing history. Printing History, 6(1): ??, ???? 1984. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-11-
0.

Sullivan:1986:BRM

[Sul86] Larry E. Sullivan. Book review: Madeleine B. Stern, Antiquarian Bookselling in the United States: A History from the Origins to the 1940s. Printing History, 8(1):??, ???? 1986.

Sullivan:1987:BRS

[Sul87] Larry E. Sullivan. Book review: Sean Wilentz, Chants Democratic: New York City and the Rise of the American Working Class. Printing History, 9(2): ??, ???? 1987.

Sutcliffe:1978:OUP

[Sut78] Peter H. Sutcliffe. The Oxford University Press: an informal history. Clarendon Press, New York, NY, USA, 1978. ISBN 0-19-951084-9. xxviii + 303 pp. LCCN Z325.O94.

Suarez:2010:OCB

[SW10] Michael F. Suarez and H. R.

Woudhuysen, editors. The Oxford Companion to the Book. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2010. ISBN 0-19-860653-2 (set). lxv + 1327 (2 volumes) pp. LCCN Z4 .O89 2010.

Swanson:2013:FKA

[Swa13] Mathilde Kredel Brown Swanson. Fritz Kredel: Artist, illustrator, and grandfather. Printing History (New Series), (14): ??, July 2013. ISSN 0192-9275.

Sword:2018:BRD

[Swo18] Stephen Sword. Book review: Douglas W. Charles, Bed & Platen Book Printing Machines: American and British Streams of Ingenious Regression in the Quest for Print Quality. Printing History (New Series), (24):??, Summer 2018. ISSN 0192-9275.

Sadek:1991:TPC

[SZ91] George Sadek and Maxim Zhukov. Typography, polyglot: a comparative study in multilingual typesetting. Center for Design and Typography, New York, NY, USA, 1991. 35 pp.

Szmuk:1986:BRA

[Szm86] Szilvia E. Szmuk. Book review: Anne Anninger, Spanish and Portuguese 16th Century Books in the Department of Printing and Graphic Arts. Printing History, 8(2):??, ???? 1986.

Szmuk:1989:BRH

[Szm89] Szilvia Szmuk. Book review: How to Operate a Book. Video. Printing History, 11(1):??, ???? 1989.

Tanselle:1981:ABR

[Tan81] G. Thomas Tanselle. Analytical bibliography and Renaissance printing history. Printing History, 3(1):??, ???? 1981. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-5-
0.

Tanselle:1982:BRG

[Tan82] G. Thomas Tanselle. Book review: Geoffrey Ashall Glaister, Glaister’s Glossary of the Book: Terms Used in Papermaking, Printing, Bookbinding and Publishing with Notes on Illuminated Manuscripts and Private Presses. Printing History, 4(1–2):??, ???? 1982. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-7-
-8-0.

Tanselle:1987:TRP

[Tan87] G. Thomas Tanselle. Thoughts on research in printing history. Printing History, 9(2):??, ???? 1987.

Tanselle:1998:BRW

[Tan99] G. Thomas Tanselle. Book review: William Tomlinson and Richard Masters, Bookcloth 1823–1980. Printing History, 19(1):??, ???? 1998–1999. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-37-
0.

Taylor:1990:BRJ

[Tay90] Marvin J. Taylor. Book review:

John Feather, A Dictionary of Book History. Printing History, 12(1):??, ???? 1990.

Taylor:1995:TP

[Tay95] W. Thomas Taylor. The temper of the present. Printing History, 17(1):??, ???? 1995. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-33-
0.

Todd:1988:TIE

[TB88] William B. (William Burton)

Todd and Ann Bowden. Tauchnitz international editions in English, 1841–1955: a bibliographical history. Bibliographical Society of America, New York, NY, USA, 1988. ISBN 0-914930-12-5. xxiv + 1078 pp. LCCN Z321.3.B47 T63 1988.

Tebbel:1978:HBP

[Teb78] John William Tebbel. A history of book publishing in the United States. Volume III, The Golden age between two wars, 1920– 1940. R. R. Bowker Company, 1180 Avenue of the Americas, New York, NY 10036, USA, 1978. ISBN 0-8352-0498-7 (hardcover). xiii + 774 pp. LCCN Z473 .T42.

Tebbel:1979:BRP

[Teb79] John Tebbel. Book review: Peter Sutcliffe, The Oxford University Press: An Informal History [and] Nicholas Barker, The Oxford University Press and the Spread of Learning, 1478– 1978. Printing History, 1(2): ??, ???? 1979. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-2-
0.

Tebbel:1986:BRM

[Teb86] John Tebbel. Book review: Max Hall, Harvard University Press: A History. Printing History, 8(2):??, ???? 1986.

Tebbel:1987:BCR

[Teb87] John William Tebbel. Between covers: The rise and transformation of book publishing in America. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1987. ISBN 0-19-504189-5. xi + 514 pp. LCCN Z473 .T4

1987. URL http://www.loc.
gov/catdir/enhancements/
fy0637/86012859-d.html.

Theriault:1983:BRE

[Thé83] Michel Thériault. Book review: Elizabeth Hulse, A Dictionary of Toronto Printers, Publishers, Booksellers and Allied Trades, 1798– 1900. Printing History, 5(2): ??, ???? 1983. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-10-
0.

Thompson:1979:EJI

[Tho79] Susan Thompson. Early journals of interest to APHA, excerpts from five earlier printing related journals: The Typographic Messenger, The Inland Printer, The American Bookmaker, The Engraver and Printer, and The Printing Art. Printing History, 1(1):??, ???? 1979. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-1-
0.

Thompson:1980:NBR

[Tho80a] Neville Thompson. A note on Bruce Rogers in youth and age. Printing History, 2(1):??, ???? 1980. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-3-
0.

Thompson:1980:BRD

[Tho80b] Susan O. Thompson. Book review: Diane Chalmers Johnson, American Art Nouveau. Printing History, 2(2):??, ???? 1980. ISSN 0192-9275.

URL https://www.rit.edu/

                                        press/printing-history-4-
                                        0.

Thompson:1983:BRL

[Tho83]  Neville Thompson.  Bruce

Rogers looks back. Printing History, 5(1):??, ???? 1983. ISSN 0192-9275.

                                        URL https://www.rit.edu/
                                        press/printing-history-9-
                                        0.

Thompson:1988:AGD

[Tho88] Bradbury Thompson. The art of graphic design. Yale University Press, New Haven, CT, USA, 1988. ISBN 0-300-04301-5. xii + 218 pp. LCCN Z116.A3 T45 1988. With contributions by noteworthy designers and critics and art historians.

Thomson:2006:GFL

[Tho06] Ellen Mazur Thomson. The Graphic Forms Lectures. Printing History, 25(1):??, ???? 2006. ISSN 0192-9275.

                                        URL https://www.rit.edu/
                                        press/printing-history-49-
                                        0.

Thomson:2011:CB

[Tho11] Ellen Mazur Thomson. The Cornhill Booklet, 1900–1914. Printing History (New Series), (9):??, January 2011. ISSN 0192-9275.

Thomson:2019:FOE

[Tho19] Ellen Mazur Thomson. To feast with one’s eyes: Colour printing in Nineteenth-Century cookbooks. Printing History (New Series), (26):??, Summer 2019. ISSN 0192-9275.

Tichenor:1989:TVU

[Tic89] Irene Tichenor. Theodore De Vinne: Unlikely leader. Printing History, 11(1):??, ???? 1989.

Tichenor:2005:NAC

[Tic05] Irene Tichenor. No art without craft: the life of Theodore Low De Vinne, printer. David R. Godine, Publisher, Jaffrey, NH, USA, 2005. ISBN 1-56792-286-4. 329 pp. LCCN Z232.D52 T53 2005.

Tidcombe:2002:DP

[Tid02]  Marianne Tidcombe.  The

Doves Press. British Library, London, UK, 2002. ISBN 0-7123-4708-9 (British Library), 1-58456-084-3 (Oak Knoll Press). xii + 258 + 8 pp. LCCN Z232.C65 T53 2002.

Tomlinson:1996:BSE

[TM96] William Tomlinson and Richard Masters. Bookcloth 1823– 1980: a study of early use and the rise of manufacture, Winterbottom’s dominance of the trade in Britain and America, production methods and costs and the identification of qualities and designs. D. Tomlinson, Stockport, Cheshire, England, 1996. ISBN 0-9527736-0-0. xvi + 143 + 11 pp. LCCN Z269.3.P8 T66 1996.

Traister:1979:BRE

[Tra79] Daniel Traister. Book review: Edward Hodnett, Francis Barlow: First Master of English book Illustration. Printing History, 1(2):??, ???? 1979. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-2-
0.

Traister:1981:RBI

[Tra81]  Daniel Traister.  The Re-

naissance Book, introduction to conference papers. Printing History, 3(1):??, ???? 1981. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-5-
0.

Traister:1982:BRE

[Tra82] Daniel Traister. Book review: Elizabeth L. Eisenstein, The Printing Press as an Agent of Change: Communications and Cultural Transformation in Early-Modern Europe. Printing History, 4(1–2): ??, ???? 1982. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-7-
-8-0.

Traister:1984:BRM

[Tra84] Daniel Traister. Book review: Miriam Usher Chrisman. Lay Culture, Learned Culture: Books and Social Change in Strasbourg, 1480-1599 and Bibliography of Strasbourg Imprints, 1480–1599. Printing History, 6(2):??, ???? 1984.

Tracy:1994:WEL

[Tra94] Walter Tracy. Why Egyptian? Lawrence W. Wallis, type designs by George W. Jones for the Linotype machine. Printing History, 16(1–2):??, ???? 1994.

Traister:2017:BRC

[Tra17] Daniel Traister. Book review: Claire M. Bolton, The Fifteenth-Century Printing Practices of Johann Zainer, Ulm, 1473–1478. Printing History (New Series), (21):??, Winter 2017. ISSN 0192-9275.

Trissel:1984:BRR

[Tri84] James Trissel. Book review:

Roderick Cave. The Private Press. Printing History, 6(2): ??, ???? 1984.

Twain:1968:OFP

[TS68] Mark Twain and Robert Stowell. The old fashioned printer. The Green Leaf Press, South Diamond Harbour, ??, ??, 1968. 8 unnumbered pp. LCCN ????

Turner:1981:BRR

[Tur81]  J. Rigbie Turner.  Book

review: Richard J. Wolfe, Early American Music Engraving and Printing: A History of Music Publishing in America from 1787 to 1825. Printing History, 3(2): ??, ???? 1981. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-5-
0.

Uglow:2006:NEL

[Ugl06] Jennifer S. Uglow. Nature’s engraver: a life of Thomas Bewick. Faber and Faber, London, UK, 2006. ISBN 0-571-22374-5 (hardcover), 0-571-22375-3. xix + 458 + 16 pp. LCCN NE1147.6.B47 U35 2006.

Uzanne:1979:EBR

[Uza79]  Octave Uzanne.  The end

of books [reprinted from Scribner’s (1896)]. Printing History, 1(2):??, ???? 1979. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-2-
0.

VanWingen:1979:BRN

[Van79] Peter M. Van Wingen. Book review: Nicolette Gray, Nineteenth Century Ornamented Typefaces. Printing History, 1 (1):??, ???? 1979. ISSN 0192-

9275. URL https://www.rit.
edu/press/printing-history-
1-0.

VanWingen:1984:BRC

[Van84] Peter M. VanWingen. Book review: Christian Axel-Nilsson. Type Studies: The Norstedt Collection of Matrices in the Typefoundry of the Royal Printing Office; a History and Catalogue. Printing History, 6 (2):??, ???? 1984.

Vervliet:2010:FRP

[Ver10] Hendrik D. L. Vervliet. French Renaissance printing types: a conspectus. Oak Knoll Press, New Castle, DE, USA, 2010. ISBN 1-58456-271-4 (Oak Knoll.), 0-948170-18-2 (Bible Society), 0-900003-14-6 (PHS).

471 pp. LCCN Z250.A2 V46 2010.

Voremberg:2019:CJI

[Vor19] Fuchsia Voremberg. Chinook jargon and itinerant mimeography in the Pacific Northwest. Printing History (New Series), (26):??, Summer 2019. ISSN 0192-9275.

Walker:1982:WCP

[Wal82] Gay Walker. The works of Carl P. Rollins. Yale University Library, New Haven, CT, USA, 1982. xii + 200 pp. LCCN Z232.R68 W34 1982.

Walker:1988:PUS

[Wal88] Gay Walker. Printing for the United States: Meriden gravure and the U. S. Government. Printing History, 10(2): ??, ???? 1988.

Walkup:1995:BRR

[Wal95] Kathleen A. Walkup. Book review: Roger Levenson, Women in Printing. Northern California, 1857–1890. Printing History, 17(1):??, ???? 1995. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-33-
0.

Wallis:2004:GWJ

[Wal04] L. W. Wallis. George W. Jones, printer laureate. Mark Batty Publisher, West New York, NJ, USA, 2004. ISBN 0-9725636-7-9. 128 + 8 pp. LCCN Z232.J726 W35 2004.

Waters:1984:BRS

[Wat84] Sheila Waters. Book review: Stephen Harvard. The Cataneo Manuscript: An Italic Copybook. Printing History, 6(2):??, ???? 1984.

Watry:2001:VPC

[Wat01] Maureen Watry. The Vale Press: Charles Ricketts, a publisher in earnest. British Library, London, UK, 2001. ISBN 0-7123-4724-0. 304 + 8 pp. LCCN Z325.V34 W38 2004.

Watry:2004:VPC

[Wat04] Maureen M. Watry. The Vale Press: Charles Ricketts, a publisher in earnest. Oak Knoll Press, New Castle, DE, USA, 2004. ISBN 1-58456-072-X. xxiv + 266 + 8 pp. LCCN Z325.V34 W38 2004.

Wood:2010:DSS

[WB10] Frances Wood and Mark Barnard. The Diamond Sutra: the story of the World’s earliest dated printed book. The British Library, London, UK, 2010. ISBN 0-7123-5090-X (hardcover). 112 pp. LCCN Z186.C5 W66 2010.

Wagner:1982:PRC

[WCB82] Henry R. (Henry Raup) Wagner, Charles Lewis Camp, and Robert H. Becker. The Plains and the Rockies: a critical bibliography of exploration, adventure, and travel in the American West, 1800–1865. J. Howell-Books, San Francisco, CA, USA, fourth edition, 1982. ISBN 0-910760-11-X. xx + 745 + 32 pp. LCCN F591; Z1251.W5 W2 1982.

Weber:1985:DRP

[Web85] Renée I. Weber. Doctoral research in printing history, 1970–1984. Printing History, 7 (2):??, ???? 1985.

Wees:1985:BRA

[Wee85] J. Dustin Wees. Book review:

Anthony Dyson, Pictures to Print: The Nineteenth-century Engraving Trade. Printing History, 7(2):??, ???? 1985.

Wees:1988:BRG

[Wee88] J. Dustin Wees. Book review: Gavin Bridson and Geoffrey Wakeman, Printmaking & Picture Printing: A Bibliographical Guide. Printing History, 10(1):??, ???? 1988.

Weimerskirch:1983:EWR

[Wei83] Philip J. Weimerskirch. Ed., W. R. Chambers & Co., The Externals of Books [reprint]. Printing History, 5(2):??, ???? 1983. ISSN 0192-9275.

Weimerskirch:1989:LSA

[Wei89] Philip J. Weimerskirch. Lithographic stone in America. Printing History, 11(1):??, ???? 1989.

Weimerskirch:2002:RAS

[Wei02] Philip J. Weimerskirch. The Rev. Abraham O. Stansbury and the Stansbury Press.

Printing History, 21(2):??, Spring 2002. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-42-
0.

Weimerskirch:2005:BCP

[Wei05] Philip J. Weimerskirch. The beginning of color printing in America. Printing History, 24(2):??, ???? 2005. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-48-
0.

Wells:1980:BRS

[Wel80] James Wells. Book review: Scott Bruntjen and Melissa L. Young, Douglas McMurtrie: Bibliographer and Historian of Printing. Printing History, 2(1):??, ???? 1980. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-3-
0.

Whiteman:1991:ISH

[Whi92] Maxwell Whiteman. The introduction and spread of Hebrew type in the United States. Printing History, 13/14(2/1): ??, ???? 1991–1992.

Whitesell:2005:TJB

[Whi05] David R. Whitesell. Thomas Jefferson and the book arts. Printing History, 24(2):??, ???? 2005. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-48-
0.

WCS:1764:SPT

[Wil64] William Caslon and Son. A specimen of printing types, by W. Caslon and Son, letter founders, in London. Dryden Leach, London, UK, 1764. 39 pp. LCCN Z250 .W736 1764.

Wilentz:1984:CDN

[Wil84] Sean Wilentz. Chants democratic: New York City and the rise of the American working class, 1788–1850. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1984. ISBN 0-19-503342-6 (hardcover), 0-19-504012-0 (paperback). xii + 446 + 12 pp. LCCN HD8085.N53 W54 1984.

Willes:2008:RMF

[Wil08] Margaret Willes. Reading matters: five centuries of discovering books. Yale University Press, New Haven, CT, USA, 2008. ISBN 0-300-12729-4 (hardcover), 0-300-14236-6 (ebook). xvi + 295 pp. LCCN Z987.5.G7 W55 2008.

Winans:1981:DCB

[Win81] Robert B. Winans. A descriptive checklist of book catalogues separately printed in America, 1693–1800. American Antiquarian Society, Worcester, MA, USA, 1981. ISBN 0-912296-47-X. xxxi + 207 pp. LCCN Z1029 .W56.

Winship:1983:PPN

[Win83]  Michael Winship.  Print-

ing with plates in the Nineteenth Century United States. Printing History, 5(2):??, ???? 1983. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-10-
0.

Wingen:1991:BRJ

[Win92] Peter Van Wingen. Book review: John Dreyfus, A Typographical Masterpiece: An Account of the Golden Cockerel Press Edition of ‘The Four Gospels’ in 1931. Printing History, 13/14(2/1):??, ???? 1991–1992.

Winship:1995:APH

[Win95] Michael Winship. The art preservative: From the history of the book back to printing history. Printing History, 17(1):??, ???? 1995. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-33-
0.

Winton:2007:BRJ

[Win07] Calhoun Winton. Book review: James N. Green and Peter Stallybrass, Benjamin Franklin, Printer and Writer. Printing History (New Series), (1):??, January 2007. ISSN 0192-9275.

Winship:2011:PSC

[Win11] Michael Winship. Pirates, shipwrecks, and comic almanacs: Charles Ellms packages books in Nineteenth-Century America. Printing History (New Series), (9):??, January 2011. ISSN 0192-9275.

Wolfe:1980:EAM

[Wol80] Richard J. Wolfe. Early American music engraving and printing: a history of music publishing in America from 1787 to 1825 with commentary on earlier and later practices. Music in American life. University of Illinois Press, Urbana, IL, USA, 1980. ISBN 0-252-00726-3. xix + 321 pp. LCCN ML112 .W64.

Woodward:1986:MMP

[Woo86] David Woodward. Maps, music, and the printer: Graphic or typographic? Printing History, 8(2):??, ???? 1986.

WP:2007:BRA

[WP07a] WP. Book review: Ari Rafaeli, Book Typography. Printing History (New Series), (1):??, January 2007. ISSN 0192-9275.

WP:2007:BRB

[WP07b] WP. Book review: Betty Bright, No Longer Innocent: Book Art in America. Printing History (New Series), (2): ??, July 2007. ISSN 0192-9275.

WP:2007:BRC

[WP07c] WP. Book review: Christopher de Hamel and Joel Silver, Disbound and Dispersed: The Leaf Book Considered. Printing History (New Series), (1):??, January 2007. ISSN 0192-9275.

WP:2007:BRD

[WP07d] WP. Book review: Don Hauser, Printers of the Streets and the Lanes of Melbourne. Printing History (New Series), (2):??, July 2007. ISSN 0192-9275.

WP:2007:BRI

[WP07e] WP. Book review: Irene Tichenor, No Art without Craft: The Life of Theodore Low DeVinne, Printer. Printing History (New Series), (1): ??, January 2007. ISSN 0192-9275.

WP:2007:BRJb

[WP07f] WP. Book review: Jenny Uglo, Nature’s Engraver: A Life of Thomas Bewick. Printing History (New Series), (2):??, July 2007. ISSN 0192-9275.

WP:2007:BRJa

[WP07g] WP. Book review: John A. Lane and Robert Slimbach, Garamond Premier Pro: A Contemporary Adaptation. Printing History (New Series), (1):??, January 2007. ISSN 0192-9275.

WP:2007:BRL

[WP07h] WP. Book review: Lawrence Wallis, George W. Jones: Printer Laureate. Printing History (New Series), (1):??, January 2007. ISSN 0192-9275.

WP:2007:BRMd

[WP07i] WP. Book review: Mathieu Lommen, Sem Hartz and the Making of Linotype Juliana.

Printing History (New Series), (2):??, July 2007. ISSN 0192-9275.

WP:2007:BRMb

[WP07j] WP. Book review: Maureen Watry, The Vale Press: Charles Ricketts, A Publisher in Ernest. Printing History (New Series), (1):??, January 2007. ISSN 0192-9275.

WP:2007:BRMa

[WP07k] WP. Book review: Michael Mitchell and Susan Wightman, Book Typography: A Designer’s Manual. Printing History (New Series), (1):??, January 2007. ISSN 0192-9275.

WP:2007:BRMc

[WP07l] WP. Book review: Moira F.

Harris and Leo J. Harris, Minnesota on Paper: Collecting our Printed History. Printing History (New Series), (2):??, July 2007. ISSN 0192-9275.

WP:2007:BRRb

[WP07m] WP. Book review: Reynolds Stone, The Albion Press. Printing History (New Series), (1): ??, January 2007. ISSN 0192-9275.

WP:2007:BRRc

[WP07n] WP. Book review: Richard B. Doubleday, Jan Tschicold, Designer: The Penguin Years. Printing History (New Series), (2):??, July 2007. ISSN 0192-

WP:2007:BRRa

[WP07o] WP. Book review: Richard Southall, Printer’s Type in the Twentieth Century: Manufacturing and Design Methods. Printing History (New Series), (1):??, January 2007. ISSN 0192-9275.

WP:2007:BRS

[WP07p] WP. Book review: Simon Loxley, Type: The Secret History of Letters. Printing History (New Series), (2):??, July 2007. ISSN 0192-9275.

WP:2007:BRT

[WP07q] WP. Book review: Type and Typography: Highlights from Matrix, the Review for Printers and Bibliophiles. Printing History (New Series), (1):??, January 2007. ISSN 0192-9275.

WP:2007:BRV

[WP07r] WP. Book review: Virginia Smith, Forms in Modernism: The Unity of Typography, Architecture and the Design Arts. Printing History (New Series), (2):??, July 2007. ISSN 0192-9275.

WP:2008:BRA

[WP08] WP. Book review: Alice H. R. H. Beckwith, Illustrating the Good Life: The Pissarros’ Eragny Press, 1894–1914. A Catalogue of an Exhibition of Books, Prints and Drawings Related to the Work of the Press. Printing History (New Series), (3):??, January 2008. ISSN 0192-9275.

WP:2009:BRD

[WP09a] WP. Book review: Donald Farren and August A. Imholtz, Jr., eds., The Baltimore Bibliophiles at Fifty, 1954–2004. Printing History (New Series), (6):??, July 2009. ISSN 0192-9275.

WP:2009:BRE

[WP09b] WP. Book review: Elizabeth Evenden, Patents, Pictures and Patronage: John Day and the Tudor Book Trade. Printing History (New Series), (6):??, July 2009. ISSN 0192-9275.

WP:2009:BRJb

[WP09c] WP. Book review: Jerry Kelly, The First Flowering: Bruce Rogers at the Riverside Press, 1896–1912. Printing History (New Series), (6):??, July 2009. ISSN 0192-9275.

WP:2009:BRJa

[WP09d] WP. Book review: John Buchanan-Brown, Early Victorian Illustrated Books: Britain, France, and Germany, 1820– 1860. Printing History (New Series), (5):??, January 2009. ISSN 0192-9275.

WP:2009:BRM

[WP09e] WP. Book review: Margaret Willes, Reading Matters: Five Centuries of Discovering Books. Printing History (New Series), (6):??, July 2009. ISSN 0192-9275.

WP:2009:BRN

[WP09f]  WP.  Book review: Neil

Macmillan, An A–Z of Type Designers. Printing History (New Series), (5):??, January 2009. ISSN 0192-9275.

WP:2009:BRRa

[WP09g] WP. Book review: Richard Landon, Humane Letters: Bruce Rogers, Designer of Books and Artist. Printing History (New Series), (5):??, January 2009. ISSN 0192-9275.

WP:2009:BRRd

[WP09h] WP. Book review: Rob Banham and Fiona Ross, eds., Non-Latin Typefaces. Printing History (New Series), (6):??, July 2009. ISSN 0192-9275.

WP:2009:BRRc

[WP09i] WP. Book review: Robert Oldham, A Field Guide to North American Hand Presses and Their Manufacturers. Printing History (New Series), (6): ??, July 2009. ISSN 0192-9275.

WP:2009:BRRb

[WP09j] WP. Book review: Robin Dodd, From Gutenberg to Opentype: An Illustrated History of Type from the Earliest Letterforms to the Latest Digital Fonts. Printing History (New Series), (5):??, January 2009. ISSN 0192-9275.

WP:2009:BRH

[WP09k] WP. Book review: T. H. Howard-Hill, The British Book Trade, 1475–1890: A Bibliography. Printing History (New Series), (6):??, July 2009. ISSN 0192-9275.

Wright:1992:DHS

[Wri92] Helena E. Wright. Dard Hunter at the Smithsonian. Printing History, 14(2):??, ???? 1992.

Wright:2007:FPM

[Wri07]  Helena Wright.  The fu-

ture of printing museums. Printing History, 25(2):??, ???? 2007. ISSN 0192-9275.

URL https://www.rit.edu/
press/printing-history-50-
0.

Wronker:1991:PPP

[Wro91] Erich Wronker. A picture portfolio of printing medals. Printing History, 14(1):??, ???? 1991.

Yeager:2012:SKB

[Yea12] Jonathan M. Yeager. Samuel Kneeland of Boston, colonial bookseller, printer and publisher of religion. Printing History (New Series), (11):??, January 2012. ISSN 0192-9275.

Young:2010:FTL

[You10a] Matthew McLennan Young.

Field and Tuer, the Leadenhall Press: a checklist with an appreciation of Andrew White Tuer. Oak Knoll Press, New Castle, DE, USA, 2010. ISBN 1-58456-280-3 (USA), 0-7123-5807-2 (UK). vii + 144 + 16 pp. LCCN Z232.L396 Y68 2010.

Young:2010:RFP

[You10b] Matthew McLennan Young. The rise and fall of the Printers’ International Specimen Exchange. Printing History (New Series), (8):??, July 2010. ISSN 0192-9275.

Yousif:2012:BGR

[You12] Keri Yousif. Balzac, Grandville, and the rise of book illustration. Ashgate, Farnham, Surrey, England, 2012. ISBN 1-4094-1808-1 (hardcover), 1-4094-1809-X (e-book). ix + 200 pp. LCCN NC980.5.G67 Y68 2012.

Zlatkes:2013:OHT

[Zla13] Gwido Zlatkes. One hundred and twenty thousand “duplicator hours”: Underground printing in Communist Poland, 1976–1989. Printing History (New Series), (13):??, January 2013. ISSN 0192-9275.